### v1

# Zeroth-order log-concave sampling

Abstract
We study the zeroth-order query complexity of log-concave sampling, specifically uniform sampling from convex bodies using membership oracles. We propose a simple variant of the proximal sampler that achieves the query complexity with matched Rényi orders between the initial warmness and output guarantee. Specifically, for any $\varepsilon>0$ and $q\geq 2$, the sampler, initialized at $\pi_{0}$, outputs a sample whose law is $\varepsilon$-close in $q$-Rényi divergence to $\pi$, the uniform distribution over a convex body in $\mathbb{R}^{d}$, using $\widetilde{O}(qM_{q}^{q/(q-1)}d^{2}\,\lVert\operatorname{cov}\pi\rVert\log\frac{1}{\varepsilon})$ membership queries, where $M_{q}=\lVert\text{d}\pi_{0}/\text{d}\pi\rVert_{L^{q}(\pi)}$.
We further introduce a simple annealing scheme that produces a warm start in $q$-Rényi divergence (i.e., $M_{q}=O(1)$) using $\widetilde{O}(qd^{2}R^{3/2}\,\lVert\operatorname{cov}\pi\rVert^{1/4})$ queries, where $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$. This interpolates between known complexities for warm-start generation in total variation and Rényi-infinity divergence. To relay a Rényi warmness across the annealing scheme, we establish hypercontractivity under simultaneous heat flow and translate it into an improved mixing guarantee for the proximal sampler under a logarithmic Sobolev inequality. These results extend naturally to general log-concave distributions accessible via evaluation oracles, incurring additional quadratic queries.
1 Introduction
We study the zeroth-order query complexity of sampling from an arbitrary log-concave distribution $\pi$, assuming access to an evaluation oracle for $-\log\pi$ (up to an additive constant). This is a very general setting for log-concave sampling—for example, it includes uniform sampling from convex bodies, which serves as a running example throughout this paper. Results in this general setting provide a theoretical baseline, highlighting the benefits of additional assumptions often made in other sampling work, such as log-smoothness (i.e., $-\nabla\log\pi$ is Lipschitz) or access to first-order oracles (i.e., $-\nabla\log\pi$). Indeed, much of our work is motivated by gaps between this general case and more favorable settings where first-order oracles are available for positive and log-smooth distributions. In practical terms, zeroth-order samplers are natural default algorithms, since evaluation oracles typically require only function value computations, making them the easiest to implement.
We now make the problem setup more explicit, focusing for now on uniform sampling. Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body (i.e., a convex and compact set with non-empty interior) containing the $d$-dimensional unit ball $B_{1}$. Given access to the membership oracle for $\mathcal{K}$ (which answers whether a queried point belongs to $\mathcal{K}$) [GLS93], we ask: how can we sample uniformly from $\mathcal{K}$, and how many oracle queries are required? More formally, given a target accuracy $\varepsilon>0$ and a probability divergence (or distance) $\mathsf{D}$—such as the total variation ($\mathsf{TV}$) distance or $\chi^{2}$-divergence—what is the query complexity of generating a sample whose law is $\varepsilon$-close to the uniform distribution $\pi$ over $\mathcal{K}$ in $\mathsf{D}$?
In this work, we propose a simple uniform sampler that essentially subsumes prior methods developed over the past three decades. Its simplicity allows us to establish the first balanced complexity guarantee in a general setting, meaning that the Rényi orders of the initial warmness and the output metric are the same. This result parallels guarantees found in recent sampling research in the log-smooth setting. Next, we provide a simple annealing scheme that, for the first time, generates a warm start in $q$-Rényi divergence. To this end, we prove hypercontractivity under simultaneous heat flow, given a logarithmic Sobolev inequality. Finally, we extend these results to log-concave distributions under evaluation oracles. Alongside, we present a streamlined analysis leveraging powerful tools from diverse mathematical fields.
1.1 Backgrounds and results
In this section, $\pi_{\mathcal{K}}$ denotes the uniform distribution over a convex body $\mathcal{K}$ with diameter $D$, and $\pi$ refers to log-concave distributions over $\mathbb{R}^{d}$. For clarity, we recall the $q$-Rényi divergence, defined as $\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{q}{q-1}\log\,\lVert\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}\rVert_{L^{q}(\nu)}$ for $q>1$, and its monotonicity in $q$, and $2\,\lVert\cdot\rVert_{\mathsf{TV}}^{2}\leq\mathsf{KL}=\lim_{q\downarrow 1}\mathcal{R}_{q}\leq\mathcal{R}_{2}=\log(1+\chi^{2})\leq\chi^{2}\leq\mathcal{R}_{\infty}=\log\operatorname{ess\,sup}_{\nu}\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}$. See §1.4 for more details. When a sampler is initialized from an initial distribution $\nu_{0}$ toward a target distribution $\nu$, we define the warmness parameter as $M_{q}=\lVert\mathrm{d}\nu_{0}/\mathrm{d}\nu\rVert_{L^{q}(\nu)}$.
1.1.1 Uniform sampling from a warm start
Classical algorithms: Ball walk and Hit-and-run.
The $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$ are well-studied random walks (i.e., Markov chain Monte Carlo methods) for uniform sampling. These algorithms are analyzed through the conductance framework, where a lower bound on conductance yields a mixing-time guarantee. Establishing such a lower bound involves two ingredients: one-step coupling (i.e., the closeness of transition kernels at two nearby points) and a Cheeger isoperimetric inequality for the target distribution. The Cheeger constant of a probability measure $\pi$ over $\mathbb{R}^{d}$ is defined as
$C_{\mathsf{Ch}}(\pi):=\inf_{S:\pi(S)\leq\frac{1}{2}}\frac{\pi(\partial S)}{\pi(S)}\,,$
where $S\subset\mathbb{R}^{d}$ is any open set with smooth boundary, $S_{\eta}=\{x:d(x,S)\leq\eta\}$, and $\pi(\partial S):=\liminf_{\eta\downarrow 0}\frac{\pi(S_{\eta})-\pi(S)}{\eta}$.
The $\mathsf{Ball\ walk}$ was introduced by Lovász [Lov90]; in each step, it samples a random point uniformly from a ball of fixed radius around the current point, accepting it if the point lies within $\mathcal{K}$. Later, Lovász and Simonovits established that the $\mathsf{Ball\ walk}$ satisfies $\mathcal{R}_{\infty}\to\mathsf{TV}$ with complexity $\operatorname{poly}(d,D,M_{\infty}/\varepsilon)$ [LS93, Remark on p. 398], meaning the initial warmness should be given in $\mathcal{R}_{\infty}$-divergence and the final output is $\varepsilon$-close to $\pi_{\mathcal{K}}$ in $\mathsf{TV}$-distance. They also analyzed an indirect approach that uses the $\mathsf{Ball\ walk}$ (with a Metropolis filter) to sample from a log-concave distribution and then post-processes the sample to bring it closer to $\pi_{\mathcal{K}}$. The resulting algorithm, also of type $\mathcal{R}_{\infty}\to\mathsf{TV}$, was shown to have query complexity $\operatorname{poly}(d,D,\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$.
This approach was further improved by Kannan, Lovász, and Simonovits [KLS97], who introduced the $\mathsf{Speedy\ walk}:\mathcal{R}_{\infty}\to\mathsf{TV}$. Similar to [LS93], it begins by sampling from a biased log-concave variant $\pi^{\prime}$ (of $\pi_{\mathcal{K}}$), which is invariant under the following per-step transition: the next point $x_{i+1}$ is sampled uniformly from $\mathcal{K}\cap B_{\delta}(x_{i})$ via rejection sampling, using the uniform distribution over $B_{\delta}(x_{i})$ as the proposal. Then, a post-processing step is used to reduce the bias. They showed that the mixing time of this chain is $O(d^{2}C_{\mathsf{Ch}}^{-2}(\pi^{\prime})\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$ and that each rejection sampling step requires $O(M_{\infty})$ queries, leading to an overall complexity of $\widetilde{O}(M_{\infty}d^{2}D^{2}\log\nicefrac{{1}}{{\varepsilon}})$ (due to $C_{\mathsf{Ch}}(\pi)\gtrsim D^{-1}$ [LS93]). Hereafter, by following convention, the $\mathsf{Ball\ walk}$ refers to the $\mathsf{Speedy\ walk}$ with the post-processing step. Later, progress on the Kannan-Lovász-Simonovits (KLS) conjecture [KLS95] improved this bound to $C_{\mathsf{Ch}}^{2}(\pi)\gtrsim(\lVert\operatorname{cov}\pi\rVert\log d)^{-1}$ [Kla23]. As a result, the complexity of the $\mathsf{Ball\ walk}$ is now $\widetilde{O}(M_{\infty}d^{2}\lVert\operatorname{cov}\pi\rVert\log\nicefrac{{1}}{{\varepsilon}})$.
The $\mathsf{Hit\text{-}and\text{-}Run}$ algorithm [Smi84] was first analyzed by Lovász [Lov99] via $s$-conductance [LS90], with subsequent refinements by Lovász and Vempala [LV06b]. Each step selects a random chord passing the current point and moves to a random point on the chord. Here, uniform sampling from the chord is implemented via rejection sampling with complexity $O(\log D)$, where the proposal distribution is uniform over a slightly longer chord (found by binary search). They established a novel Cheeger isoperimetric inequality in terms of the cross-cut ratio (rather than Euclidean distance) and showed $\mathsf{Hit\text{-}and\text{-}Run}:\mathcal{R}_{2}\to\mathcal{R}_{2}$ with complexity $\widetilde{O}(d^{2}D^{2}\log\nicefrac{{M_{2}}}{{\varepsilon}})$, making it the first random walk with a balanced guarantee. Thus, from an $O(1)$-warm start in $\mathcal{R}_{\infty}(\geq\mathcal{R}_{2})$, the $\mathsf{Ball\ walk}$ achieves faster mixing (and better query complexity) than $\mathsf{Hit\text{-}and\text{-}Run}$.
However, despite these advances, a complete and clean understanding of this fundamental problem is still lacking. First, the $\mathsf{Ball\ walk}$ takes an indirect approach to uniform sampling: as noted earlier, it samples from a biased distribution to obtain a mixing guarantee with $\log\nicefrac{{M_{\infty}}}{{\varepsilon}}$-dependence, followed by a post-processing step to reduce the bias. Second, conductance-based analyses require an algorithmic modification making the chains lazy (i.e., the chain remains at its current point with some probability). To achieve greater simplicity and elegance, it would be desirable to develop a direct sampling algorithm that avoids these additional processing steps. Finally, the highly unbalanced guarantee of the $\mathsf{Ball\ walk}$ (i.e., $\mathcal{R}_{\infty}\to\mathsf{TV}$) appears unnatural. As with $\mathsf{Hit\text{-}and\text{-}Run}$, we are led to ask whether we can obtain a balanced guarantee using the improved Cheeger constant $C_{\mathsf{Ch}}^{2}(\pi)\gtrsim(\lVert\operatorname{cov}\pi\rVert\log d)^{-1}$. This question becomes even more natural when considering recent developments in sampling under log-smoothness.
Sampling under smoothness.
Modern research in sampling has made substantial progress by analyzing discretizations of continuous-time processes. A central example is the Langevin diffusion $(X_{t})_{t\geq 0}$ in $\mathbb{R}^{d}$, described by the stochastic differential equation
$\mathrm{d}X_{t}=-\nabla V(X_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}B_{t}\,,\qquad X_{0}\sim\nu_{0}\,,$
where $V:\mathbb{R}^{d}\to\mathbb{R}$ is a potential function satisfying mild regularity conditions, and $B_{t}$ denotes Brownian motion. The law $\nu_{t}$ of $X_{t}$ converges to the stationary distribution $\nu\propto e^{-V}$ as $t\to\infty$. A spectral analysis of the infinitesimal generator for this process yields a connection between the rate of convergence and the Poincaré constant of $\nu$ (see [BGL14, §4]). A probability measure $\pi$ on $\mathbb{R}^{d}$ is said to satisfy a Poincaré inequality with constant $C$ if for any locally Lipschitz function $f\in L^{2}(\pi)$,
$\operatorname{Var}_{\pi}f:=\int\Bigl{(}f-\int f\,\mathrm{d}\pi\Bigr{)}^{2}\,\mathrm{d}\pi\leq C\int|\nabla f|^{2}\,\mathrm{d}\pi\,,$ ($\mathsf{PI}$)
and the smallest such $C$ is called the Poincaré constant $C_{\mathsf{PI}}(\pi)$. Furthermore, the seminal work of Jordan, Kinderlehrer, and Otto [JKO98] interpreted Langevin diffusion as the $2$-Wasserstein gradient flow of $\mathsf{KL}(\cdot\mathbin{\|}\nu)$, revealing a connection to the logarithmic Sobolev inequality, which is strictly stronger than the Poincaré inequality. A probability measure $\pi$ on $\mathbb{R}^{d}$ is said to satisfy a logarithmic Sobolev inequality with constant $C$ if for any locally Lipschitz function $f\in L^{2}(\pi)$,
$\operatorname{Ent}_{\pi}(f^{2}):=\int f^{2}\log f^{2}\,\mathrm{d}\pi-\int f^{2}\,\mathrm{d}\pi\cdot\log\int f^{2}\,\mathrm{d}\pi\leq 2C\int|\nabla f|^{2}\,\mathrm{d}\pi\,,$ ($\mathsf{LSI}$)
and the smallest such $C$ is referred to as the log-Sobolev constant $C_{\mathsf{LSI}}(\pi)$. Under ($\mathsf{PI}$) and ($\mathsf{LSI}$) respectively, the Langevin process satisfies the exponential convergences:
$\chi^{2}(\nu_{t}\mathbin{\|}\nu)\leq\exp\bigl{(}-\frac{2t}{C_{\mathsf{PI}}(\nu)}\bigr{)}\,\chi^{2}(\nu_{0}\mathbin{\|}\nu)\,,\quad\text{and}\quad\mathsf{KL}(\nu_{t}\mathbin{\|}\nu)\leq\exp\bigl{(}-\frac{2t}{C_{\mathsf{LSI}}(\nu)}\bigr{)}\,\mathsf{KL}(\nu_{0}\mathbin{\|}\nu)\,,$ (1.1)
which were later extended to $\mathcal{R}_{q}$-divergence in [CLL19, VW23].
With bounded smoothness (i.e., $\beta:=\lVert\nabla^{2}V\rVert<\infty$), a short-time discretization of the Langevin process yields an algorithm that preserves the balanced guarantees of the continuous process. Specifically, it has been shown to converge in $\mathsf{KL}$ (i.e., $\mathcal{R}_{1}\to\mathcal{R}_{1}$) with complexity $\operatorname{poly}(d/\varepsilon,\beta,C_{\mathsf{LSI}}(\nu),\log\mathsf{KL}(\nu_{0}\mathbin{\|}\nu))$ [VW23]. This result was further extended to $\mathcal{R}_{q}$-divergence [CEL+24] achieving a query complexity of $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{PI}}(\nu),\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$ or $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{LSI}}(\nu),\log\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$. We note that the dependence on $\nicefrac{{1}}{{\varepsilon}}$ can be made polylogarithmic by incorporating a Metropolis filter (e.g., $\mathsf{MALA}$ [CLA+21, WSC22, CG23a]) or by considering $\mathsf{MHMC}$ [CV22, CDWY20, CG23b]). For a comprehensive review, we refer readers to the monograph [Che25].
The complexity of the $\mathsf{Ball\ walk}$ can be partially described as $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{PI}}(\nu),\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$. In fact, for log-concave probability measures $\pi$, it holds that
$\frac{1}{4}\leq\frac{C_{\mathsf{Ch}}^{-2}(\pi)}{C_{\mathsf{PI}}(\pi)}\leq 9\,,$
where the first inequality is due to Cheeger [Che70], and the second follows from the Buser–Ledoux inequality [Bus82, Led04]. Thus, $C_{\mathsf{PI}}$ could be replaced with $C_{\mathsf{Ch}}^{-2}$, although it remains unclear how to interpret the smoothness parameter $\beta$ in the context of convex-body sampling.
Rise of the proximal sampler.
Following this line of work, the $\mathsf{Proximal\ sampler}$ ($\mathsf{PS}$) [LST21] has brought substantial progress to sampling research. For a target distribution $\pi^{X}\propto e^{-V}$ over $\mathbb{R}^{d}$, $\mathsf{PS}$ considers the augmented distribution $\pi(x,y)\propto\pi^{X}(x)\,\gamma_{h}(y-x)$ over $\mathbb{R}^{d}\times\mathbb{R}^{d}$, where $\gamma_{h}$ denotes the centered Gaussian with covariance $hI_{d}$. Each iteration consists of two steps: for step size $h>0$, (i) $y_{k+1}\sim\pi^{Y|X=x_{k}}=\mathcal{N}(x_{k},hI_{d})$, and (ii) $x_{k+1}\sim\pi^{X|Y=y_{k+1}}$. When $\pi^{X}$ is $\beta$-(log-)smooth, rejection sampling for the backward step $x_{i+1}\sim\pi^{X|Y=y_{i+1}}$ can be performed using $O(1)$ evaluation queries and one proximal query [^1] The proximal query for a function $V:\mathbb{R}^{d}\to\mathbb{R}$ and step size $h>0$ is the function defined as $x\in\mathbb{R}^{d}\mapsto\arg\min\{V(\cdot)+\frac{1}{2h}\,\lVert\cdot-x\rVert^{2}\}$. for any $y_{i+1}$ and $h\asymp(\beta d)^{-1}$.
Even more surprising results were obtained by [CCSW22], which showed that one iteration of $\mathsf{PS}$ can be interpreted as the composition of two continuous-time stochastic processes—Brownian motion and its time-reversal. Since Brownian motion naturally defines a heat semigroup $(P_{t})_{t\geq 0}$ from $L^{2}(\nu*\gamma_{t})$ to $L^{2}(\nu)$, defined by $P_{t}f:=f*\gamma_{t}$, and the time-reversal also admits a (non-homogeneous) semigroup $(Q_{t})_{t\geq 0}$ from $L^{2}(\nu)$ to $L^{2}(\nu*\gamma_{t})$, defined as $Q_{t}f:=P_{t}(f\nu)/P_{t}\nu$ [KP23], this allowed [CCSW22] to carry out a semigroup calculus, leading to contraction results similar to (1.1). For instance, this framework yields a query complexity of $O(qd\beta C_{\mathsf{PI}}(\nu)\,\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu)\log\nicefrac{{1}}{{\varepsilon}})$ for achieving $\varepsilon$-closeness in $\mathcal{R}_{q}$ (i.e., $\mathsf{PS}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ under ($\mathsf{PI}$)). See §3.1 for further details on these semigroups.
Proximal sampler for uniform sampling.
Kook, Vempala, and Zhang [KVZ24] adapted this idea to convex-body sampling by introducing the $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ algorithm. They extended the contraction result to constrained distributions and developed a new implementation for sampling from $\pi^{X|Y=y}\propto\gamma_{h}(\cdot-y)|_{\mathcal{K}}$. They introduced a threshold parameter $\tau$: if the number of retries in the rejection sampling with proposal $\mathcal{N}(y,hI_{d})$ exceeds $\tau$, then $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ halts and declares failure. To control this failure probability $\eta\in(0,1)$, they carried out involved analysis to carefully set the parameters $h,\eta,\tau$ so that with high probability, the backward step can be completed using $\widetilde{O}(M_{\infty}\operatorname{polylog}\nicefrac{{dD}}{{\eta\varepsilon}})$ queries. Consequently, they established $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ with query complexity $\widetilde{O}(qM_{\infty}d^{2}\Lambda\operatorname{polylog}\nicefrac{{D}}{{\eta\varepsilon}})$ for $\Lambda:=\lVert\operatorname{cov}\pi\rVert$, generalizing $\mathsf{Ball\ walk}$’s output guarantee. Later, Kook and Zhang [KZ25] established a balanced guarantee of $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ using ($\mathsf{LSI}$), with query complexity $\widetilde{O}(M_{\infty}d^{2}D^{2}\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$.
Nevertheless, algorithmic tricks (e.g., the failure probability) still remain in $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$, and the analysis becomes unwieldy due to intricate interdependencies among the parameters: $h,\eta,\tau$, and the number $N$ of iterations. Moreover, the required initial warmness remains stringent as $\mathcal{R}_{\infty}$. The warmness requirement was later relaxed to $\mathcal{R}_{c}$ with $c=2+\operatorname{polylog}\nicefrac{{dD}}{{\eta\varepsilon}}=\widetilde{O}(1)$ for achieving $\varepsilon$-closeness in $\mathcal{R}_{2}$ (i.e., $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{c}\to\mathcal{R}_{2}$ under ($\mathsf{PI}$)), with query complexity $\widetilde{O}(M_{c}d^{2}\Lambda\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$ [KV25a]. While this result gave hope for balancing the orders of warmness and final guarantee, the complex parameter dependencies continued to be a major obstacle.
Algorithm Type Complexity Reference
[…]
$\mathsf{Ball\ walk}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $\operatorname{poly}(d,D,\frac{M_{\infty}}{\varepsilon})$Rightarrow $\frac{M_{\infty}^{2}d^{2}\Lambda}{\varepsilon^{2}}$ [LS93]
$\mathsf{Speedy\ walk}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $M_{\infty}d^{2}D^{2}\log\frac{1}{\varepsilon}$Rightarrow $M_{\infty}d^{2}\Lambda\log\frac{1}{\varepsilon}$ [KLS97]
$\mathsf{Hit\text{-}and\text{-}Run}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $\frac{M_{\infty}^{2}d^{2}D^{2}}{\varepsilon^{2}}$ [Lov99]
$\mathcal{R}_{2}\to\mathcal{R}_{2}$ $d^{2}D^{2}\log\frac{M_{2}}{\varepsilon}$ [LV06b]
$\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ $\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ $qM_{\infty}d^{2}\Lambda\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KVZ24]
$\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ $M_{\infty}d^{2}D^{2}\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KZ25]#
$\mathcal{R}_{\widetilde{O}(1)}\to\mathcal{R}_{2}$ $M_{\widetilde{O}(1)}d^{2}\Lambda\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KV25a]
$\mathsf{PS}_{\textup{unif}}$ $\mathcal{R}_{q}\to\mathcal{R}_{q}$ $qM_{q}^{q/(q-1)}d^{2}\Lambda\log\frac{1}{\varepsilon}$ Theorem 1.1
Table 1: Query complexities of uniform samplers from a convex body. The right arrow ($\Rightarrow$) indicates subsequent improvements due to the improved Poincaré constant of log-concave probability measures. $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ requires the additional parameter $\eta$ referring to a failure probability of the algorithm. The $\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ result in # can be readily extended to $\mathsf{PS}_{\textup{unif}}$ as well, with guarantee $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ of complexity $M_{\infty}d^{2}D^{2}\log\frac{1}{\varepsilon}$ (i.e., removal of $\eta$).
Proximal sampler with restart for balanced guarantees (§2).
Our proposed sampler, $\mathsf{PS}_{\textup{unif}}$—essentially $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ with failure replaced by restart—resolves this issue: starting with $x_{0}\sim\pi_{0}^{X}$, step size $h$, and threshold $\tau$,
  • [Forward] $y_{k+1}\sim\pi^{Y|X=x_{k}}=\gamma_{h}(\cdot-x_{k})$.
[…]
One iteration refers to the successful completion of both steps. Note that each check $x_{k+1}\in\mathcal{K}$ incurs a single query to the membership oracle.
This algorithm allows $h$ and $\tau$ to depend solely on the dimension and initial warmness, leading to the same Rényi orders of initial warmness and final guarantee while leveraging the state-of-the-art bound on Poincaré constants of log-concave probability measures (i.e., $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ under ($\mathsf{PI}$)).
Theorem 1.1 (Uniform sampling from a warm start).
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ with $B_{1}\subset\mathcal{K}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$, given access to the membership oracle for $\mathcal{K}$. Let $\pi_{k}$ denote the law of the $k$-th iterate of $\mathsf{PS}_{\textup{unif}}$, initialized from $\pi_{0}$, with step size $h\asymp(d^{2}\,\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi))^{-1}$, and threshold $\tau\asymp\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)\,e^{\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}$ for $q\geq 2$. Then, for any $\varepsilon>0$, we can ensure $\mathcal{R}_{q}(\pi_{N}\mathbin{\|}\pi)\leq\varepsilon$ after $N\asymp qd^{2}\Lambda\,\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)\log d\log\frac{1}{\varepsilon}$ iterations, using
$2\tau N=\widetilde{O}\bigl{(}qd^{2}\Lambda\,e^{\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}\log\frac{1}{\varepsilon}\bigr{)}\qquad\text{membership queries in expectation}\,.$
Under the warmness notation, the query complexity reads $\widetilde{O}(qd^{2}\Lambda M_{q}^{\frac{q}{q-1}}\log\frac{1}{\varepsilon})$.
1.1.2 Order-preserving warm-start generation via annealing
$\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ exhibits polynomial dependence on $M_{\widetilde{O}(1)}$, and a “cold start” (e.g., the uniform distribution over a small ball) yields $M_{\widetilde{O}(1)}\lesssim\exp(d,D)$. Hence, warm-start generation is necessary to reduce the complexity overhead incurred by the cold start, so warm-start generating algorithms have been developed alongside sampling algorithms.
Warm-start generation in the well-conditioned setting.
The necessity of a warmness is not limited to this general setting; its importance also arises in the well-conditioned setting (i.e., $-\log\pi$ is $\alpha$-strongly convex and $\beta$-smooth). In fact, $\mathsf{MALA}$ from an $O(1)$-warm start in $\mathcal{R}_{2}$ can take a step size $h\asymp(\beta d^{1/2})^{-1}$, achieving the best first-order complexity of $\widetilde{O}(\kappa d^{1/2}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}}$) in the well-conditioned setting. In contrast, from a feasible start, it can only take a smaller step size $h\asymp(\beta d)^{-1}$, resulting in $\widetilde{O}(\kappa d\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}})$ complexity, where $\kappa:=\beta/\alpha$ denote the condition number of $-\log\pi$. Recently, [AC24] showed that combining the Underdamped Langevin Algorithm with $\mathsf{PS}$ can generate an $O(1)$-warm start in $\mathcal{R}_{2}$-divergence using $\widetilde{O}(\kappa d^{1/2})$ first-order queries, thus incurring no overhead for $\mathsf{MALA}$ as a result.
Simulated annealing.
In our general setup, a classical strategy for warm-start generation is an annealing scheme, which involves a sequence of probability distributions $\{\mu_{i}\}_{i\leq m}$ that interpolate between a tractable initial distribution $\mu_{0}$ and the target distribution $\pi_{\mathcal{K}}$. The key idea is to ensure that each pair of consecutive distributions $\mu_{i}$ and $\mu_{i+1}$ are $O(1)$-close in some appropriate metric (e.g., $\mathcal{R}_{\infty}$-divergence), thereby enabling efficient transitions by a sampler along the sequence.
In the context of convex-body sampling, several annealing schemes have been proposed: (i) uniform distributions over truncated convex bodies intersected with balls of increasing radii [DFK91, LS90, LS93, KLS97], (ii) exponential distributions of the form $e^{-cx}$ with decreasing $c$ over a convex set, derived from the so-called “pencil construction” [LV06c], and (iii) Gaussian distributions with increasing variance, truncated to the body [CV15, CV18, KZ25, KV25a].
Among these, all approaches except [KZ25] provide a warm start $\mu$ satisfying $\lVert\mu-\pi_{\mathcal{K}}\rVert_{\mathsf{TV}}\leq\varepsilon$, while [KZ25] guarantees $\mathcal{R}_{\infty}$-warmness with complexity $\widetilde{O}(d^{2}R^{2})$ for $R^{2}=\mathbb{E}_{\pi_{\mathcal{K}}}[|\cdot|^{2}]$. On the other hand, $\mathsf{TV}$-warmness can be achieved with better complexity $\widetilde{O}(d^{2}R^{3/2}\Lambda^{1/4}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}})$ [KV25a]. Unfortunately, no known result currently interpolates between warmness results in $\mathsf{TV}$ and $\mathcal{R}_{\infty}$ (i.e., either very weak or strong warmness, without known intermediate warmness).
Toward genuine warmness.
First of all, why do we even care about $\mathcal{R}_{q}$-warmness? Given a theoretical gap in known complexities (for generating $\mathcal{R}_{q}$-warmness) between the two settings discussed above, bridging this gap is theoretically intriguing in its own right. In our general setup, it is not even clear whether generating $\mathcal{R}_{q}$-warmness without significant overhead is feasible, let alone whether one can tighten the complexities of warm-start generation and sampling from a warm start.
Moreover, $\mathsf{TV}$-warmness serves only as a “fake” warm-start for downstream sampling in some sense, since subsequent sampling guarantees using this warmness are collapsed down to $\mathsf{TV}$-distance due to the necessity of applying triangle inequalities—an issue referred to as “$\mathsf{TV}$-collapse” by [KZ25]. In practice, this limitation leads to cumulative sampling errors, resulting in somewhat cumbersome post-processing steps, such as replacing the original target accuracy $\varepsilon$ by a reduced accuracy $\varepsilon/m$, where $m$ is the number of times the triangle inequality is applied. Lastly, when combined with $\mathsf{PS}_{\textup{unif}}$ that achieves balanced guarantees, a $\mathcal{R}_{q}$-warmness result would finally yield clean guarantees for uniform sampling in terms of any $\mathcal{R}_{q}$-divergences.
$\mathsf{TV}$-collapse.
A main difficulty in obtaining $\mathcal{R}_{q}$-warmness through annealing is that we do not yet know how to relay $\mathcal{R}_{q}$-warmness without compromise, leading nearly all previous work to encounter this $\mathsf{TV}$-collapse problem. To illustrate this, consider annealing distributions $\mu_{0},\mu_{1},\mu_{2}$, where a sampler transitions from $\mu_{0}$ toward $\mu_{1}$. MCMC samplers are inherently approximate, producing a measure $\bar{\mu}_{1}$ close to $\mu_{1}$, quantified via $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})\leq\varepsilon$ for some $q>1$. In the next phase, the sampler moves toward $\mu_{2}$ starting from the approximate measure $\bar{\mu}_{1}$ instead of the true $\mu_{1}$. Ideally, the complexity for this phase would directly depend on $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})$. Unfortunately, the lack of an exact triangle inequality for $\mathcal{R}_{q}$-divergence prevents us from ensuring $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})=O(1)$, even if $\mathcal{R}_{q}(\mu_{1}\mathbin{\|}\mu_{2})=O(1)$.
To circumvent this issue, prior studies have resorted to using the triangle inequality (or a coupling argument) for $\mathsf{TV}$-distance. Specifically, although $\bar{\mu}_{1}$ differs from $\mu_{1}$, annealing algorithms proceeds as if the initial distribution were $\mu_{1}$, not $\bar{\mu}_{1}$, when running a sampler toward $\mu_{2}$. Denoting the transition kernel of this sampler by $P$ and the output distribution of this sampler by $\bar{\mu}_{2}=\bar{\mu}_{1}P$, we can deduce that
$\lVert\bar{\mu}_{2}-\mu_{2}\rVert_{\mathsf{TV}}\leq\lVert\bar{\mu}_{2}-\mu_{1}P\rVert_{\mathsf{TV}}+\lVert\mu_{1}P-\mu_{2}\rVert_{\mathsf{TV}}\leq\lVert\bar{\mu}_{1}-\mu_{1}\rVert_{\mathsf{TV}}+\lVert\mu_{1}P-\mu_{2}\rVert_{\mathsf{TV}}\,,$
where the second inequality follows from the data-processing inequality (2.1). Since $\mathcal{R}_{q}$-divergence implies $\mathsf{TV}$-distance, and each term on the RHS is bounded by $\varepsilon$, this leads to the accumulation of error over $m$ phases: $\lVert\bar{\mu}_{m}-\mu_{m}\rVert_{\mathsf{TV}}\leq m\varepsilon$. This necessitates accuracy adjustments like replacing $\varepsilon$ with $\varepsilon/m$.
Boosting a Rényi-order.
The above discussion highlights the fundamental bottleneck arising from the absence of a triangle inequality for Rényi divergences. One promising direction to overcome this is to leverage a weak triangle inequality (1.3), namely, $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})\lesssim\mathcal{R}_{2q-1}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})+\mathcal{R}_{2q}(\mu_{1}\mathbin{\|}\mu_{2})$. As [KV25a] has already designed annealing schemes that can ensure $\mathcal{R}_{2q}(\mu_{1}\mathbin{\|}\mu_{2})=O(1)$, one might consider starting with a $\mathcal{R}_{2q-1}$-guarantee when sampling from $\mu_{1}$. However, this effectively amounts to requiring a $\mathcal{R}_{\Theta(2^{m}q)}$-guarantee for $\mu_{1}$-sampling when there are $m$ many phases. Since $\mathsf{PS}$ has linear dependence on $q$, and prior annealing approaches require $m\gtrsim d^{1/2}$, such a “backtracking” analysis would ultimately incur an $\Omega(2^{\sqrt{d}}q)$-multiplicative overhead in the final complexity.
Given the failure of this naïve approach, a more valid question is whether one can ensure $\mathcal{R}_{2q-1}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})=O(1)$ given $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})<\varepsilon$. Since $2q-1>q$ for $q>1$, reversing the monotonic property of $\mathcal{R}_{q}$-divergence at some overhead will be a main challenge. This reflection naturally introduces an algorithmic question: can we boost Rényi-orders algorithmically (i.e., can we transform $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)=O(1)$ into a stronger bound $\mathcal{R}_{2q-1}(\mu^{\prime}\mathbin{\|}\pi)=O(1)$)? Solving this question would enhance our theoretical toolkit, providing the final piece toward comprehensive convex-body sampling and potentially influencing applications in areas such as differential privacy.
Hypercontractivity of semigroups.
Perhaps the most relevant mathematical notion to tackle this question is Gross’ hypercontractivity of semigroups [Gro75] (see Proposition 3.1): for a semigroup $(P_{t})_{t\geq 0}$ reversible with respect to an invariant measure $\pi$, it is known that ($\mathsf{LSI}$) for $\pi$ is equivalent to the property that $\|P_{t}f\|_{L^{q(t)}(\pi)}\leq\|f\|_{L^{p}(\pi)}$ for any function $f\in L^{p}(\pi)$, where $q(t)$ is an increasing function with $q(0)=p$. For instance, the continuous-time Langevin diffusion satisfies this property, but it is not immediately clear whether such hypercontractivity is preserved upon discretization. Recent advancements, such as [CEL+24], showed that discretized Langevin processes retain hypercontractivity under LSI when the target $\pi$ is positive and log-smooth.
Nonetheless, extending hypercontractivity algorithmically beyond this favorable setting remains nontrivial. In particular, circumstances involving constraints—such as sampling from a convex body—or more general smoothness assumptions (e.g., $B_{1}\subseteq\mathcal{K}$) have resisted straightforward solutions. One might consider using reflected Langevin diffusions, but discretizations like the projected Langevin algorithm [BEL18] require access to a projection oracle for $\mathcal{K}$, and existing analytical guarantees similarly collapse to $\mathsf{TV}$-distance. Thus, achieving hypercontractivity in broader contexts remains as an open and intriguing direction.
Hypercontractivity under simultaneous heat flow (§3).
To address the boosting question, we explore an alternative strategy involving a different Markov semigroup that can be algorithmically realized via $\mathsf{PS}$. Specifically, we aim to establish hypercontractivity for the $L^{2}(\nu)$-adjoint $Q_{t}$ of the heat semigroups $P_{t}$, under ($\mathsf{LSI}$) for the target measure $\nu$. Formally, we state the following hypercontractivity result for $Q_{t}$:
Theorem 1.2 (Hypercontractivity under heat flow).
Let $\nu$ be a probability measure satisfying ($\mathsf{LSI}$) with $C_{\mathsf{LSI}}(\nu)<\infty$. We denote $\nu_{t}=\nu*\gamma_{t}$ for $t\geq 0$, and define $q:[0,\infty)\to(1,\infty)$ by
$\frac{q(t)-1}{p-1}=1+\frac{t}{C_{\mathsf{LSI}}(\nu)}\quad\text{for any }p\in(1,\infty)\,.$
Then, the $L^{2}(\nu)$-adjoint $Q_{t}$ of heat semigroups satisfies
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f\rVert_{L^{p}(\nu)}\quad\text{for all }p\in(1,\infty),t\geq 0,\text{ and }f\in L^{p}(\nu)\,.$
Its immediate consequence follows by selecting $f=\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$ for a probability measure $\mu\ll\nu$. Since $Q_{t}f=\frac{P_{t}(f\nu)}{P_{t}\nu}$, this yields $\lVert\mathrm{d}\mu_{t}/\mathrm{d}\nu_{t}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert\mathrm{d}\mu/\mathrm{d}\nu\rVert_{L^{p}(\nu)}$, which implies $\mathcal{R}_{q(t)}(\mu_{t}\mathbin{\|}\nu_{t})\leq\mathcal{R}_{p}(\mu\mathbin{\|}\nu).$ This result has algorithmic implications for $\mathsf{PS}$. Below, the constant $2$ can be replaced by any number larger than $1$.
Corollary 1.3 (Hypercontractivity of proximal sampler).
Let $\pi_{k}^{X}$ be the law of the $k$-th iterate returned by the $\mathsf{Proximal\ sampler}$ with step size $h>0$, initial $\pi_{0}^{X}$, and target $\pi^{X}$. Given $1<p<q$, we can ensure $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\mathcal{R}_{p}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$ after $N\lesssim h^{-1}C_{\mathsf{LSI}}(\pi^{X})\log_{2}\frac{q-1}{p-1}$ iterations. Moreover, given $\varepsilon>0$ and $q_{0},q>1$, we can attain $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
$N\lesssim h^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\pi_{0}^{X}\mathbin{\|}\pi^{X})}{(2\wedge q_{0}-1)\,\varepsilon}\quad\text{iterations\,.}$
This result translates hypercontractivity into practical algorithmic outcomes, enabling efficient implementation of warm-start generation. Moreover, it leads to $\widetilde{O}(\kappa d\log\frac{q}{\varepsilon})$-mixing of $\mathsf{PS}$ in the well-conditioned setting (see Proposition 3.2), improving the linear dependence of $\mathsf{PS}$ on $q$ [CCSW22].
Streamlined order-preserving warm-start generation (§4).
We streamline the design and analysis of a known Gaussian-annealing scheme for warm-start generation, incorporating the hypercontractivity established for $\mathsf{PS}$. This streamlined approach offers an elegant and clean picture for warm-start generation via annealing schemes.
As with $\mathsf{PS}_{\textup{unif}}$, we propose and analyze its Gaussian variant in §4.1, denoted by $\mathsf{PS}_{\textup{Gauss}}$, specifically tailored for truncated Gaussian distributions. Leveraging the hypercontractivity result for $\mathsf{PS}$ and the bound $C_{\mathsf{LSI}}(\gamma_{\sigma^{2}}|_{\mathcal{K}})\leq\sigma^{2}$, we derive an enhanced guarantee of $\mathsf{PS}_{\textup{Gauss}}$.
Theorem 1.4.
Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with $B_{1}(0)\subset\mathcal{K}$, given access to the membership oracle. For $q_{0}>1$, let $\mu_{k}$ denote the law of the $k$-th iterate of $\mathsf{PS}_{\textup{Gauss}}$, initialized from $\mu_{0}$ toward the target $\mu=\gamma_{\sigma^{2}}|_{\mathcal{K}}$, with step size $h\asymp(d^{2}\,\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu))^{-1}$ and threshold $\tau\asymp\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}$. Then, for any $\varepsilon>0$ and $q>1$, we can ensure $\mathcal{R}_{q}(\mu_{N}\mathbin{\|}\mu)\leq\varepsilon$ after $N\asymp d^{2}\sigma^{2}\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}{(2\wedge q_{0}-1)\,\varepsilon}$ iterations, using
$2\tau N=\widetilde{O}\bigl{(}d^{2}\sigma^{2}\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}\log\frac{q}{(2\wedge q_{0}-1)\,\varepsilon}\bigr{)}\qquad\text{membership queries in expectation}\,.$
Equipped with $\mathsf{PS}_{\textup{Gauss}}$, our warm-start generation process is very simple: (1) sample from $\mu_{1}\propto\pi_{\mathcal{K}}\gamma_{d^{-1}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$, and (2) while $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ toward $\mu_{i+1}\propto\pi_{\mathcal{K}}\gamma_{\sigma_{i+1}^{2}}$ using $\mathsf{PS}_{\textup{Gauss}}$, where $\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$ and $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,(1+\nicefrac{{\sigma_{i}}}{{R}})$.
This simpler algorithm eliminates unnecessary pre-processing steps, such as truncation of the convex body, and post-processing, such as adjusting target accuracy. Moreover, a key practical advantage of our streamlined method is the introduction of early-stopping at $\sigma^{2}\approx qR\Lambda^{1/2}\log^{1/2}d$. Consequently, faster-sampling phases beyond $\sigma^{2}\gtrsim R\Lambda^{1/2}\log^{1/2}d$ in [KV25a] become redundant.
On theoretical side, hypercontractivity enables the propagation of $\mathcal{R}_{q}$-warmness throughout the annealing scheme. Furthermore, we demonstrate in §4.2 that closeness between successive annealing distributions follows directly from Rényi versions of ($\mathsf{LSI}$) and ($\mathsf{PI}$), leading to the following comprehensive result for warm-start generation:
Theorem 1.5.
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ with $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$, given access to the membership oracle for $\mathcal{K}$. For any $q\geq 2$, there exists an annealing algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)=O(1)$, using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$ membership queries in total.
1.1.3 Log-concave sampling under an evaluation oracle
Algorithm Type Complexity Reference
[…]
$\mathsf{Ball\ walk}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $\frac{M_{\infty}^{4}d^{2}R^{2}}{\varepsilon^{4}}$ [LV07]
$\mathsf{Hit\text{-}and\text{-}Run}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $\frac{M_{\infty}^{4}d^{2}R^{2}}{\varepsilon^{4}}$ [LV07]
$\mathcal{R}_{2}\to\mathcal{R}_{2}$ $d^{2}R^{2}\operatorname{polylog}\frac{M_{2}}{\varepsilon}$ [LV06a]
$\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ $\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ $qM_{\infty}d^{2}(\Lambda\vee 1)\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KV25b]
$\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ $M_{\infty}d^{2}R^{2}\operatorname{polylog}\frac{1}{\eta\varepsilon}$
$\mathcal{R}_{\widetilde{O}(1)}\to\mathcal{R}_{2}$ $M_{\widetilde{O}(1)}d^{2}(\Lambda\vee 1)\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KV25a]
$\mathsf{PS}_{\textup{exp}}$ $\mathcal{R}_{q}\to\mathcal{R}_{q}$ $qM_{q}^{q/(q-1)}d^{2}(\Lambda\vee 1)\log\frac{1}{\varepsilon}$ Theorem 1.6
[…]
The uniform-sampling results extend naturally to general log-concave distributions $\pi\propto e^{-V}$, given access to an evaluation oracle for a convex function $V:\mathbb{R}^{d}\to\mathbb{R}\cup\{\infty\}$. In this setting, the regularity condition $B_{1}\subset\mathcal{K}$ translates to a condition on a level set of $V$, namely, that the ground set $\mathsf{L}_{\pi,g}:=\{x\in\mathbb{R}^{d}:V(x)-\min V\leq 10d\}$ contains a unit ball. This generalization is justified by the fact that $\mathsf{L}_{\pi,g}$ takes up a constant fraction of the $\pi$-measure (see [KV25b, Lemma 2.6]).
[…]
Recent work [KV25b, KV25a] revisited this problem by adapting $\mathsf{PS}$ to the general log-concave setting. Although $\mathsf{PS}$ can in principle be extended to any distribution where $\pi^{X|Y=y}\propto\pi^{X}(\cdot)\,\gamma_{h}(\cdot-y)$ is readily samplable (e.g., when given access to a proximal oracle in the log-smooth setting), it is unclear how to implement this step for arbitrary $\pi^{X}$. To address this, [KV25b] introduced the exponential lifting technique: for $\pi^{X}\propto e^{-V}$, consider the lifted distribution
$\mathrm{d}\pi^{X,T}(x,t)\propto e^{-dt}\,\mathds{1}_{\mathcal{K}}(x,t)\,\mathrm{d}x\mathrm{d}t\,,\quad\text{where }\mathcal{K}:=\{(x,t)\in\mathbb{R}^{d}\times\mathbb{R}:V(x)\leq dt\}\,.$ ($\mathsf{exp\text{-}lifting}$)
Here, the $X$-marginal of $\pi^{X,T}$ is exactly $\pi^{X}$, and $\mathcal{K}$ is convex whenever $\pi^{X}$ is log-concave. Thus, log-concave sampling reduces to the simpler problem of sampling from a log-concave distribution with a linear potential over a convex set. We then apply $\mathsf{PS}$ to this exponential distribution, denoted by $\mathsf{PS}_{\textup{exp}}$: for $z=(x,t)\in\mathbb{R}^{d+1}$ and $y\in\mathbb{R}^{d+1}$, it alternates $y\sim\pi^{Y|Z=z}=\mathcal{N}(z,hI_{d+1})$ and $z\sim\pi^{Z|Y=y}\propto\mathcal{N}(y-h\alpha,hI_{d+1})|_{\mathcal{K}}$ for $\alpha=de_{d+1}$, where the second step is implemented via rejection sampling, using the proposal $\mathcal{N}(y-h\alpha,hI_{d+1})$ with threshold $\tau$.
In [KV25b], it was shown that $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ with query complexity $\widetilde{O}(qM_{\infty}d^{2}(\Lambda\vee 1)\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$ and $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ with complexity $\widetilde{O}(M_{\infty}d^{2}(d\vee R^{2})\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$. As in the uniform sampling case, the order of initial warmness was relaxed to $\mathcal{R}_{c}$ with $c=\widetilde{O}(1)$ in [KV25a], where $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{c}\to\mathcal{R}_{2}$ achieves query complexity $\widetilde{O}(M_{c}d^{2}(\Lambda\vee 1)\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$. However, they were unable to balance the orders of initial warmness and final guarantee due to intricate interdependencies among algorithmic parameters.
Extension to log-concave sampling (§5).
Our redesigned $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ resolves this issue, aligning the order of initial warmness with that of the final guarantee. At an additive cost of $qd^{2}$ queries, our result effectively tames a long line of work on sampling in this classical setting. See §5.1 for further details.
Theorem 1.6 (Zeroth-order log-concave sampling from a warm start).
Let $\pi\propto e^{-V}$ be a log-concave distribution over $\mathbb{R}^{d}$ with $B_{1}\subset\mathsf{L}_{\pi,g}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and an evaluation oracle for $V$. Let $\pi_{k}$ be the $X$-marginal of the $k$-th iterate of $\mathsf{PS}_{\textup{exp}}$ with initial distribution $\pi_{0}$, step size $h\asymp(d^{2}\,\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi))^{-1}$, and threshold $\tau\asymp\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)\,e^{\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}$ for $q\geq 2$. Given $\varepsilon>0$, we can ensure $\mathcal{R}_{q}(\pi_{N}\mathbin{\|}\pi)\leq\varepsilon$ with $N\asymp qd^{2}(\Lambda\vee 1)\,\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)\log d\log\frac{1}{\varepsilon}$, using
$2\tau N=\widetilde{O}\bigl{(}qd^{2}(\Lambda\vee 1)\,e^{\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}\log\frac{1}{\varepsilon}\bigr{)}\qquad\text{evaluation queries in expectation}\,.$
We remark on a connection to the well-conditioned setting. In this case, the ground set contains a ball of radius $\Theta((d/\beta)^{1/2})$ centered at $\arg\min V$. Since $\Lambda\leq\alpha^{-1}$ by the Brascamp–Lieb or Lichnerowicz inequality, $\mathsf{PS}_{\textup{exp}}$ uses $\widetilde{O}(q\,(\kappa d+d^{2})\log\nicefrac{{1}}{{\varepsilon}})$ queries from an $O(1)$-warm start in $\mathcal{R}_{q}$. With the additional $qd^{2}$ queries, this matches the known $\widetilde{O}(q\kappa d\log\nicefrac{{1}}{{\varepsilon}})$-complexity results for zeroth-order samplers in this setting when additionally given a proximal oracle. See §1.3 for further details.
Annealing approaches for uniform sampling have also been extended to arbitrary log-concave distributions [LV06a, KV25b, KV25a]. In the general case, the best-known complexity for warm-start generation remains comparable to that of the convex-body case for $\mathsf{TV}$ and $\mathcal{R}_{\infty}$-warmness, with algorithms requiring additional $d^{2}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}}$ queries. In §5.2, we extend our order-preserving warm-start generation to general log-concave distributions in a similar fashion:
Theorem 1.7.
Let $\pi\propto e^{-V}$ be a log-concave distribution over $\mathbb{R}^{d}$ with $B_{1}(0)\subset\mathsf{L}_{\pi,g}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$, and access to an evaluation oracle for $V$. For $q\geq 2$, there exists an annealing algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)=O(1)$, using $\widetilde{O}(d^{2.5}+qd^{2}(R^{3/2}\vee 1)(\Lambda^{1/4}\vee 1))$ evaluation queries in total.
1.2 Technical overview
Proximal sampler with restart.
Compared to the standard $\mathsf{PS}$, our approach has a key modification—restarting the current iteration if the number of attempts in rejection sampling for the backward step exceeds a fixed threshold. The analysis of $\mathsf{PS}$ is split to two parts: the number of iterations required to achieve desired sampling accuracy, and the per-iteration query complexity arising from rejection sampling. Both components have been well-studied in prior works. The convergence rate of $\mathsf{PS}$ under ($\mathsf{PI}$) was established in [CCSW22], along with a rigorous justification for the general de Bruijn identity [KO25]. Moreover, [KVZ24] proposed carefully tuned parameter choices for $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ that keep the per-step complexity under control.
Given these foundations, the main technical questions concern the effect of the restarting mechanism. We must ensure that restarting neither alters the transition kernel of $\mathsf{PS}$ nor introduces excessive overhead in query complexity. The first point is intuitively straightforward: each iteration of $\mathsf{PS}$ can be viewed as sampling from the true transition kernel followed by an acceptance step. The distribution of the first successful sample thus corresponds to the intended transition kernel (Lemma 2.7). The second point follows from the same perspective. Suppose that each iteration succeeds with probability $p$, and the threshold for rejection retries is set to $\tau$. Then, the expected number of retries until a successful iteration is bounded by $\tau/p$. This bound remains moderate so long as the success probability is bounded away from zero, allowing us to employ simpler choices of step size $h$ and threshold $\tau$ than those used in $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$.
Hypercontractivity under heat flow.
Overall, we follow a proof sketch in [BGL14, Theorem 5.2.3] along with technical tweaks. To begin, for any compactly supported, smooth, positive function $f$, we define the probability measure $\mathrm{d}\mu\propto f\,\mathrm{d}\pi$. Denoting the convolved measures $\mu_{t}:=\mu*\gamma_{t}$ and $\pi_{t}:=\pi*\gamma_{t}$, we introduce the functionals $\mathcal{F}(t,q):=\int(\nicefrac{{\mathrm{d}\mu_{t}}}{{\mathrm{d}\pi_{t}}})^{q}\,\mathrm{d}\pi_{t}$ and $\mathcal{G}(t):=\frac{1}{q(t)}\,\log\mathcal{F}(t,q(t))$. Our goal is to construct a suitable function $q(t)$ such that $\partial_{t}\mathcal{G}\leq 0$. Our proof proceeds by differentiating $\mathcal{G}$ in $t$, which involves computing $\partial_{t}\mathcal{F}$ and $\partial_{q}\mathcal{F}$. The time derivative $\partial_{t}\mathcal{F}$ follows from the general de Bruijn identity [KO25], while the exponent derivative $\partial_{q}\mathcal{F}$ gives rise to $\operatorname{Ent}_{\pi_{t}}$, where the log-Sobolev constant $C_{\mathsf{LSI}}(\pi_{t})$ comes into picture.
Combining these ingredients, we derive a differential inequality for $\mathcal{G}(t)$, and $\partial_{t}\mathcal{G}\leq 0$ holds provided that $C_{\mathsf{LSI}}(\pi_{t})\leq(q(t)-1)/(q(t)-1)^{\prime}$. Leveraging $C_{\mathsf{LSI}}(\pi_{t})\leq C_{\mathsf{LSI}}(\pi)+t$ [Cha04], we solve $C_{\mathsf{LSI}}(\pi)+t\leq(q(t)-1)/(q(t)-1)^{\prime}$, which holds when $q(t)-1=(1+\nicefrac{{t}}{{C_{\mathsf{LSI}}(\pi)}})\,(p-1)$. Altogether, these arguments establish the hypercontractivity of the heat adjoint $Q_{t}$. Lastly, we extend this result beyond the space of compactly supported, positive, smooth functions by employing standard approximation techniques in analysis, such as the dominated convergence theorem and Fatou’s lemma, thereby generalizing the statement to $L^{p}(\pi)$-space.
Streamlined analysis of Gaussian annealing.
In our analysis of Gaussian annealing (§4.2), we present a more principled and streamlined approach to bounding the closeness between successive annealing distributions. Whereas prior works relied on ad-hoc arguments to establish such closeness, we show that this property follows more cleanly from LSI, combined with a mollification argument that allows for a limiting process.
To formalize this idea, we recall a Rényi version of LSI established in [VW23], which states that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\nicefrac{{qC_{\mathsf{LSI}}(\pi)}}{{2}}\,\mathsf{RFI}_{q}(\mu\mathbin{\|}\pi)$ for any $q>1$ and smooth $\mu/\pi$, where $\mathsf{RFI}_{q}(\mu\mathbin{\|}\pi)=q\,\mathbb{E}_{\pi}[(\frac{\mathrm{d}\mu}{\mathrm{d}\pi})^{q}|\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\pi}|^{2}]/\mathbb{E}_{\pi}[(\frac{\mathrm{d}\mu}{\mathrm{d}\pi})^{q}]$. When bounding $\mathcal{R}_{q}(\gamma_{\sigma^{2}}|_{\mathcal{K}}\mathbin{\|}\gamma_{\sigma^{2}(1+\alpha)}|_{\mathcal{K}})$ for a convex body $\mathcal{K}$ and some $\alpha>0$, a naïve calculation—ignoring the truncation to $\mathcal{K}$—suggests that one can take $\alpha\asymp\nicefrac{{\sigma}}{{qR}}$ to make it $O(1)$. This heuristic choice recovers the scaling behavior established in earlier results.
To rigorously account for the effects of truncation to $\mathcal{K}$, we employ a mollification argument using a standard mollifier $\eta$ supported on the unit ball (see (4.2)). For $\varepsilon>0$, we define $\eta_{\varepsilon}(x):=\varepsilon^{-d}\,\eta(x/\varepsilon)$, and convolve both truncated Gaussians with $\eta_{\varepsilon}$. Since convolution with $\eta_{\varepsilon}$ yields a smooth approximation $f_{\varepsilon}:=f*\eta_{\varepsilon}$ that converges to $f$ almost everywhere as $\varepsilon\to 0$ for any locally integrable function $f$, this mollification step allows us to apply the Rényi-LSI to the regularized densities and pass to the limit, thereby recovering the desired closeness bound.
An additional benefit of this approach is that it enables a clean justification for early-stopping in the annealing schedule. Unlike known annealing algorithms, we can terminate annealing when the variance $\sigma^{2}$ reaches approximately $qR\Lambda^{1/2}$, at which point $\mathcal{R}_{q}(\gamma_{\sigma^{2}}|_{\mathcal{K}}\mathbin{\|}\pi_{\mathcal{K}})=O(1)$. This follows directly from a Rényi version of PI (see ($\mathsf{R\text{-}PI}$)), combined with the mollification argument described above. Overall, this approach simplifies the analysis and unifies the reasoning behind closeness and early-stopping in Gaussian annealing.
1.3 Related work
We survey the literature on applications of zeroth-order log-concave sampling and log-concave sampling in other settings (e.g., given access to a different set of oracles).
More on uniform sampling.
[…]
As mentioned earlier, Lovász and Simonovits [LS93] used the $\mathsf{Ball\ walk}$ to develop an indirect approach with improved complexity $\operatorname{poly}(d,D,\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$. It was achieved by refining their earlier conductance framework and localization machinery; they proved that any log-concave distribution $\pi$ supported on a convex body of diameter $D$ also satisfies $C_{\mathsf{Ch}}(\pi)\gtrsim D^{-1}$.
Recently, $\mathsf{Coordinate\ Hit\text{-}and\text{-}Run}$ ($\mathsf{CHR}$), a special variant of $\mathsf{Hit\text{-}and\text{-}Run}$, has been shown to mix in polynomial time with respect to the dimension and ($\ell_{2}$ or $\ell_{\infty}$) diameter [NS22, LV23, FV23], and with a logarithmic dependence on the initial warmness [NRS25]. This thus results in $\mathsf{CHR}:\mathcal{R}_{2}\to\mathcal{R}_{2}$.
KLS conjecture.
The Kannan-Lovász-Simonovits (KLS) conjecture [KLS95] posits that the Cheeger constant of any log-concave probability measure is bounded below by $\Omega(\lVert\operatorname{cov}\pi\rVert^{-1/2})$. Starting from $C_{\mathsf{Ch}}^{2}\gtrsim\operatorname{tr}^{-1}(\operatorname{cov})$, subsequent work [Eld13, LV24, Che21, KL22, Kla23]—notably through Eldan’s stochastic localization—has improved it to $C_{\mathsf{Ch}}^{2}\gtrsim(\lVert\operatorname{cov}\pi\rVert\log d)^{-1}$ [Kla23]. However, the improved Cheeger constant in this setting does not translate to a better Cheeger constant with respect to the cross-cut ratio.
Rounding.
A notable application is the rounding of convex bodies, or more generally, log-concave functions. It aims to make the geometry of the input object more regular—typically by applying an affine transformation that brings it into a nearly-isotropic position. This step is often a necessary preprocessing phase for achieving a fully polynomial-time algorithm, since worst-case shapes (e.g., highly skewed bodies) can cause sampling algorithms to suffer from poor mixing.
Earlier approaches to rounding relied on deterministic methods based on the ellipsoid algorithm [DFK91, LS90], which guarantees $B_{1}(0)\subset\mathcal{K}\subset B_{d^{3/2}}(0)$ after a suitable affine transformation. To improve the sandwiching ratio, randomized rounding methods were pioneered in [LS93] and further developed in [KLS97, LV06c, JLLV21, JLLV24]. These methods use sampling to estimate the covariance matrix and then apply affine transformations to round the body. The best known complexity for isotropic rounding of a convex body is $\widetilde{O}(d^{3.5}\operatorname{polylog}R)$ [JLLV24], with the analysis streamlined in [KZ24]. These rounding techniques have also been extended to log-concave distributions [LV06a, KV25b], while preserving the query complexity.
Volume computation and integration.
One of the most well-known applications of uniform sampling is volume computation, a problem with a long and rich history in theoretical computer science and convex geometry; given $\varepsilon>0$, the goal is to estimate the volume $V$ of a convex body within a relative $\varepsilon$-error, $(1-\varepsilon)\operatorname{vol}\mathcal{K}\leq V\leq(1+\varepsilon)\operatorname{vol}\mathcal{K}$.
Computing the volume of a high-dimensional convex body is known to be intractable via deterministic algorithms [Ele86, BF87]. The breakthrough result by Dyer, Frieze, and Kannan [DFK91] showed that this task admits a randomized polynomial-time algorithm using random walks. The standard technique introduced therein, referred to as the multi-phase Monte Carlo method, constructs a sequence of convex bodies such that each successive pair is close in some sense. Rewriting the volume of $\mathcal{K}$ as a telescoping product of volume ratios between successive bodies, one can efficiently estimate each ratio using samples obtained via uniform sampling over convex bodies.
This framework initiated a long line of research focused on improving the mixing time, oracle complexity, and practical implementability of such algorithms [Lov90, AK91, LS90, LS93, KLS95, KLS97, LV06b, CV15, CV18, JLLV21]. The current fastest algorithm for isotropic convex bodies is due to [CV18] with complexity $\widetilde{O}(d^{3}/\varepsilon^{2})$. A natural generalization of volume computation is the integration of a log-concave function. This general problem was studied in [LV06a, KV25b], with the same complexity $\widetilde{O}(d^{3}/\varepsilon^{2})$. See [KV25b, §5.2] for a streamlined analysis.
Sampling with additional assumptions.
For constrained sampling problems—such as the convex-body sampling—there exist approaches beyond general-purpose MCMC algorithms like the $\mathsf{Ball\ walk}$, $\mathsf{Hit\text{-}and\text{-}Run}$, and $\mathsf{PS}_{\textup{unif}}$. These alternatives often leverage additional geometric structure of the constraint set $\mathcal{K}$, assume additional regularities of target distributions (e.g., smoothness), or assume access to stronger oracles, such as projection or first-order oracles.
(1) Sampling with local geometry. This approach assumes the existence of a tractable self-concordant barrier $\phi$ for the constraint $\mathcal{K}$. At each step, a sampler draws from an anisotropic Gaussian distribution whose covariance is pertinent to the local geometry encoded by $\nabla^{2}\phi$. This leads to an affine-invariant sampler that naturally adapts to the shape of $\mathcal{K}$, without requiring a separate rounding procedure. The most notable example is the $\mathsf{Dikin\ walk}:\mathcal{R}_{2}\to\mathcal{R}_{2}$ [KN12], which, when used with a log-barrier for a polytope with $m$ linear constraints, achieves a mixing time of $O(md\log\nicefrac{{M_{2}}}{{\varepsilon}})$, and its analysis was streamlined in [SV16].
This barrier-based sampler has been generalized and improved through more efficient barrier functions [Nar16, CDWY18, GN23]. A unifying framework incorporating these developments—as well as procedures for warm-start generation—was presented in [KV24]. Moreover, complexity analysis beyond worst case was explored in [JC24], and alternative non-Euclidean schemes using log-Laplace transformations was proposed in [GLL+23].
(2) Sampling with first-order oracle. When a first-order oracle is available, faster non-Euclidean samplers become feasible. These include Riemannian Hamiltonian Monte Carlo  [GC11, LV18, GKV24], which uses Hamiltonian dynamics on the Hessian manifold with metric $\nabla^{2}\phi$, and Riemannian Langevin Algorithm [GC11, LE23], which extends Langevin dynamics to a Riemannian manifold. Other approaches include the Mirror Langevin Algorithm [ZPFP20, Jia21, AC21, LTVW22, SWW24], which use mirror maps to simulate Langevin-type updates in dual space, and the Preconditioned Langevin Algorithm [SWW25].
(3) Sampling with projection oracle. In this setting, access to a projection oracle—computing the projection of a point onto $\mathcal{K}$—enables other sampling methods: unconstrained steps (e.g., Euler–Maruyama discretization of the Langevin diffusion) are followed by projections back into the feasible set [BEL18, Leh23]. Alternatively, soft-constraint approaches have also been proposed, in which the target distribution is modified with penalty terms that discourage constraint violations while remaining close to the original distribution. Such penalized schemes have been studied in [BDMP17, GHZ24, Yu24].
(4) Zeroth-order sampling for well-conditioned distributions. Those three general-purpose samplers—the $\mathsf{Ball\ walk}$, $\mathsf{Hit\text{-}and\text{-}Run}$, and $\mathsf{Proximal\ sampler}$—have also been studied by [ALPW24], [ALZ24], and [LST21, CCSW22] in the unconstrained well-conditioned setting with access to an evaluation oracle. These samplers achieve the mixing times of $\widetilde{O}(\kappa d\log\nicefrac{{1}}{{\varepsilon}})$ under various metrics (e.g., $\mathcal{W}_{2}$, $\mathsf{KL}$, $\chi^{2}$).
Real-world applications and impacts.
Sampling research is not only of theoretical interest but also has far-reaching implications for real-world applications. A prominent application arises in systems biology, where uniform sampling from a polytope can be used to analyze the efficiency of human metabolic networks [T+13, HCT+17]. Advances in sampling techniques have enabled more scalable biological inferences. For instance, the development of non-Euclidean samplers has led to improved mixing times and computational efficiency [YK17, CF21, CK21, KLSV22, KLSV23, SC24]. Moreover, practical implementations of volume computation for convex bodies have been developed in [CV16], bridging the gap between theoretical developments and real-world usability. Collectively, these advancements demonstrate how sampling research continues to yield practical tools with deep and lasting impact across science and engineering disciplines.
1.4 Preliminaries
Notation.
For $t>0$, we reserve $\gamma_{t}$ for the centered Gaussian distribution with covariance matrix $tI_{d}$. The indicator function of a set $S\subseteq\mathbb{R}^{d}$ is denoted by $\mathds{1}_{S}(x):=[x\in S]$, and $\mu|_{S}$ refers to a distribution $\mu$ truncated to $S$ (i.e., $\mu|_{S}\propto\mu\cdot\mathds{1}_{S}$). For two probability measures $\mu,\pi$, we use $\mu\pi$ to denote the new distribution with density proportional to $\mu\pi$.
Both $a\lesssim b$ and $a=O(b)$ mean $a\leq cb$ for a universal constant $c>0$. $a=\Omega(b)$ means $a\gtrsim b$, and $a\asymp b$ means $a=O(b)$ and $a=\Omega(b)$. Lastly, $a=\widetilde{O}(b)$ means $a=O(b\operatorname{polylog}b)$. For a positive semi-definite matrix $\Sigma$, $\lVert\Sigma\rVert$ denote the operator norm of $\Sigma$.
Log-concavity.
We call a function $f:\mathbb{R}^{d}\to[0,\infty)$ log-concave if $-\log f$ is convex in $\mathbb{R}^{d}$, and a probability measure $\pi$ (or distribution) log-concave if it has a log-concave density function with respect to the Lebesgue measure. We abuse notation by using the same symbol for a distribution and density. We assume that any log-concave distributions considered in this work are non-degenerate; otherwise, we could simply work on an affine subspace on which degenerate distributions are supported. For $t\geq 0$, $\pi$ is called $t$-strongly log-concave if $-\log\pi$ is $t$-strongly convex (i.e., $-\log\pi-\frac{t}{2}\,|\cdot|^{2}$ is convex). Clearly, log-concavity is preserved under multiplication, and a classical result by Prékopa and Leindler ensures that the convolution also preserves log-concavity. A distribution is called isotropic if its barycenter is at the origin and has the identity covariance matrix. Note that as the log-concave distributions decay exponentially fast at infinity, they have finite moments of all orders.
Probability divergences and distances.
Let $\mu$ and $\nu$ denote probability measures over $\mathbb{R}^{d}$, and $\varphi:\mathbb{R}_{+}\to\mathbb{R}$ be a convex function with $\varphi(1)=0$. The $\varphi$-divergence of $\mu$ toward $\nu$ with $\mu\ll\nu$ is defined as
$D_{\varphi}(\mu\mathbin{\|}\nu):=\int\varphi\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}\,\mathrm{d}\nu\,.$
The total variation ($\mathsf{TV}$) distance, the $\mathsf{KL}$-divergence, and $\chi^{q}$-divergence for $q\in(1,\infty)$ can be recovered through $\varphi(x)=\frac{1}{2}\,|x-1|$, $x\log x$, and $x^{q}-1$, respectively. The $q$-Rényi divergence is defined as
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{1}{q-1}\,\log\bigl{(}1+\chi^{q}(\mu\mathbin{\|}\nu)\bigr{)}=\frac{1}{q-1}\,\log\,\Bigl{\|}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{\|}_{L^{q}(\nu)}^{q}\,.$ (1.2)
The Rényi-infinity divergence is defined as $\mathcal{R}_{\infty}(\mu\mathbin{\|}\nu):=\log\operatorname{ess\,sup}_{\nu}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$. A weak triangle inequality holds for Rényi divergence: for any $q>1,\lambda\in(0,1)$, and probability measures $\mu,\nu,\pi$, it holds that
$\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\frac{q-\lambda}{q-1}\,\mathcal{R}_{\frac{q}{\lambda}}(\mu\mathbin{\|}\nu)+\mathcal{R}_{\frac{q-\lambda}{1-\lambda}}(\nu\mathbin{\|}\pi)\,.$ (1.3)
For $\lambda=q/(2q-1)$, we have $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\frac{2q}{2q-1}\,\mathcal{R}_{2q-1}(\mu\mathbin{\|}\nu)+\mathcal{R}_{2q}(\nu\mathbin{\|}\pi)$. We refer readers to [vEH14, Mir17] for more properties of the Rényi divergence.
2 Uniform sampling from convex bodies
We begin by describing our approach for the general proximal sampler ($\mathsf{PS}$) and then focus specifically on $\mathsf{PS}_{\textup{unif}}$ in §2.2. $\mathsf{PS}$ with step size $h$ alternates between two steps: starting with $k=0$ and $x_{0}\sim\pi_{0}^{X}$, (i) sample $y_{k+1}\sim\pi^{Y|X=x_{k}}$ and (ii) sample $x_{k+1}\sim\pi^{X|Y=y_{k+1}}$. Precisely, let $P_{h}$ and $Q_{h}$ be the transition kernel of the forward and backward step respectively, defined as
$P_{h}(x,\mathrm{d}y)=\pi^{Y|X=x}=\mathcal{N}(x,hI_{d})\,,\qquad Q_{h}(y,\mathrm{d}x)=\pi^{X|Y=y}\propto\pi^{X}(\cdot)\,\gamma_{h}(\cdot-y)\,.$
Assume that there exists a suitable proposal distribution $\mu_{y}$ such that $(\mu_{y}\,|\,\text{accepted})=Q_{h}(y,\cdot)$. In case of uniform sampling, we use $\mu_{y}=\mathcal{N}(y,hI_{d})$ and accept the proposal if it lies inside of $\mathcal{K}$.
Given this setup, we implement the backward step using rejection sampling with proposal $\mu_{y_{k+1}}$. If this rejection loop exceeds $\tau$ trials, we discard $y_{k+1}$ and return to the forward step using the same $X=x_{k}$. We group each forward–backward pair (with rejection sampling up to $\tau$ retries) into a stage. For instance, if all $\tau$ retries fail in stage $1$, then we begin stage $2$ by drawing a new $\bar{y}_{k+1}\sim\pi^{Y|X=x_{k}}$ and sampling from $x_{k+1}\sim\pi^{X|Y=\bar{y}_{k+1}}$. Stages are repeated until one iteration completes successfully.
We now analyze the query complexity of $\mathsf{PS}$ for obtaining a sample whose law is $\varepsilon$-close to the target $\pi^{X}$ in a chosen metric (Theorem 1.1). To this end, we decouple the analysis into two components: (1) a mixing analysis and (2) a query complexity of the backward step. The first part determines how many iterations $N$ are required for the sample $X_{N}$ to be $\varepsilon$-close to $\pi^{X}$. The second one determines how many queries $\mathsf{PS}$ uses in each iteration. Therefore, the total query complexity of the algorithm is the product of these two components.
2.1 Mixing analysis
The data-processing inequality (DPI) is a handy tool in information theory; for any probability measures $\mu,\nu$, Markov kernel $P$, $\varphi$-divergence $D_{\varphi}$, and $q\in(1,\infty]$, it holds that
$D_{\varphi}(\mu P\mathbin{\|}\nu P)\leq D_{\varphi}(\mu\mathbin{\|}\nu)\,,\qquad\quad\text{and}\qquad\quad\mathcal{R}_{q}(\mu P\mathbin{\|}\nu P)\leq\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\,,$ (2.1)
where $\mu P:=\int P(x,\cdot)\,\mu(\mathrm{d}x)$ denotes the distribution obtained by applying one step of $P$ starting from $\mu$. This implies that for $D\in\{D_{\varphi},\mathcal{R}_{q}\}$ and any $k\geq 0$,
$D(\pi_{k+1}^{X}\mathbin{\|}\pi^{X})=D(\pi_{k+1}^{Y}Q_{h}\mathbin{\|}\pi^{Y}Q_{h})\leq D(\pi_{k+1}^{Y}\mathbin{\|}\pi^{Y})=D(\pi_{k}^{X}P_{h}\mathbin{\|}\pi^{X}P_{h})\leq D(\pi_{k}^{X}\mathbin{\|}\pi^{X})\,.$
However, this is not sufficient to establish a mixing rate. It turns out that adding an independent Gaussian random variable to $X_{k}$ (i.e., convolving $\operatorname{law}X_{k}$ with $\gamma_{h}$) makes $\operatorname{law}X_{k}$ contract directly toward $\pi^{X}*\gamma_{h}=\pi^{Y}$, with the contraction rate dependent on $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$.
This phenomenon, known as the strong data-processing inequality (SDPI) for the Gaussian channel [AG76], has been extensively studied in information theory [PW16, CPW18]. Given a $\varphi$-divergence $D$, the contraction coefficient is defined as
$\eta_{D}(\mu,t):=\sup_{\nu:\,D(\nu\mathbin{\|}\mu)\in\mathbb{R}_{>0}}\frac{D(\nu_{t}\mathbin{\|}\mu_{t})}{D(\nu\mathbin{\|}\mu)}\,,$
where $\nu_{t}:=\nu*\gamma_{t}$ denotes the probability measure obtained by convolving $\nu$ with $\gamma_{t}$. Clearly, $0\leq\eta_{D}(\mu,t)\leq 1$ by the DPI.
Bounding $\eta_{D}$ relies on the well-known de Bruijn identity, $\partial_{t}\mathsf{KL}(\nu_{t}\mathbin{\|}\mu_{t})=-\frac{1}{2}\,\mathsf{FI}(\nu_{t}\mathbin{\|}\mu_{t})$, where the relative Fisher information is defined as $\mathsf{FI}(\nu\mathbin{\|}\mu):=\mathbb{E}_{\nu}[|\nabla\log\frac{\mathrm{d}\nu}{\mathrm{d}\mu}|^{2}]$; see rigorous proofs by Barron [Bar84] and Klartag and Ordentlich [KO25] (for its generalizations). Other derivations can be found in [Guo09, VW23, CCSW22], assuming that probability measures involved in proofs are sufficiently smooth to justify differentiation under the integral sign and vanishment of boundary terms in integration by parts.
Combining this identity with, for example, ($\mathsf{LSI}$) for $\mu$ (i.e., $\mathsf{KL}(\nu_{t}\mathbin{\|}\mu_{t})\leq\nicefrac{{C_{\mathsf{LSI}}(\mu_{t})}}{{2}}\,\mathsf{FI}(\nu_{t}\mathbin{\|}\mu_{t})$) and $C_{\mathsf{LSI}}(\mu_{t})\leq C_{\mathsf{LSI}}(\mu)+t$ [Cha04], one can obtain a differential inequality in terms of $\mathsf{KL}$, and solving this leads to
$\eta_{\mathsf{KL}}(\mu,t)\leq\bigl{(}1+\frac{t}{C_{\mathsf{LSI}}(\mu)}\bigr{)}^{-1}\,.$
Under similar regularity assumptions, the SDPI for Rényi divergence under ($\mathsf{PI}$), ($\mathsf{LSI}$), and Latała–Oleszkiewicz inequality was studied in [CCSW22] and further extended to constrained probability measures in [KVZ24]. These results were rigorously unified and investigated in [KO25] under a $\varphi$-Sobolev inequality [Cha04] (i.e., a generalization of ($\mathsf{PI}$) and ($\mathsf{LSI}$)), with regularity assumptions explicitly identified. We refer readers to [KO25, Regularity conditions] for a detailed account.
Specifically, [CCSW22] showed that under the regularity assumptions, for any $q\geq 2$ and $\nu\ll\mu$, the following bounds hold:
$\mathcal{R}_{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq\begin{cases}\mathcal{R}_{q}(\nu\mathbin{\|}\mu)-\frac{\log(1+t/C_{\mathsf{PI}}(\mu))}{q}&\text{if }\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\nu\mathbin{\|}\mu)}{(1+t/C_{\mathsf{PI}}(\mu))^{1/q}}&\text{if }\mathcal{R}_{q}(\nu\mathbin{\|}\mu)<1\,,\end{cases}$ (2.2)
and $\chi^{2}(\nu_{t}\mathbin{\|}\mu_{t})\leq(1+t/C_{\mathsf{PI}}(\mu))^{-1}\,\chi^{2}(\nu\mathbin{\|}\mu)$. Applying this result to the forward step, along with the DPI for the backward step, implies that $\mathsf{PS}$ achieves $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ with
$N\lesssim qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{M_{q}}{\varepsilon}\,,\quad\text{where }M_{q}:=\Bigl{\|}\frac{\mathrm{d}\pi_{0}^{X}}{\mathrm{d}\pi^{X}}\Bigr{\|}_{L^{q}(\pi^{X})}\,.$
We provide a version of this result in terms of $\chi^{q}$-divergence (which generalizes their $\chi^{2}$-result), as working with $\chi^{q}$-divergence leads to cleaner expressions and a bound on the contraction coefficient $\eta_{\chi^{q}}$ under ($\mathsf{PI}$). Our proof follows the argument sketched above, combining the general de Bruijn identity in terms of $\chi^{q}$-divergence under ($\mathsf{PI}$). Below, note that any log-concave probability measure satisfies ($\mathsf{PI}$) with $C_{\mathsf{PI}}<\infty$, as established in [KLS95, Bob99].
Lemma 2.1.
Let $\mu$ be a log-concave probability measure over $\mathbb{R}^{d}$. For any $q\geq 2$, it holds that
$\eta_{\chi^{q}}(\mu,t)\leq\bigl{(}1+\frac{t}{C_{\mathsf{PI}}(\mu)}\bigr{)}^{-2/q}\quad\text{for }t>0\,.$
To prove this result, we recall the general version of de Bruijn’s identity from [KO25, Proposition 1.12].
Proposition 2.2 (General de Bruijn’s identity).
Let $\mu$ and $\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\chi^{q}(\nu\mathbin{\|}\mu)<\infty$ and $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$. Then, for any $t>0$ and $q>1$,
$\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})=-\frac{q\,(q-1)}{2}\,\mathbb{E}_{\mu_{t}}\Bigl{[}\bigl{(}\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\bigr{)}^{q}\,\bigl{|}\nabla\log\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\bigr{|}^{2}\Bigr{]}\underset{\text{if }q\geq 2}{=}-\frac{q-1}{2}\,\Bigl{\|}\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\Bigr{\|}_{L^{q}(\mu_{t})}^{q}\,\mathsf{RFI}_{q}(\nu_{t}\mathbin{\|}\mu_{t})\,,$
where $\mathsf{RFI}_{q}(\nu\mathbin{\|}\mu)=q\,\mathbb{E}_{\mu}[(\frac{\mathrm{d}\nu}{\mathrm{d}\mu})^{q}|\nabla\log\frac{\mathrm{d}\nu}{\mathrm{d}\mu}|^{2}]/\mathbb{E}_{\mu}[(\frac{\mathrm{d}\nu}{\mathrm{d}\mu})^{q}]$ denotes the $q$-Rényi Fisher information.
We also recall a Rényi version of ($\mathsf{PI}$) from [VW23, Lemma 9] that for $q\geq 2$ and probability measure $\nu$ and $\mu$ with $\nu\ll\mu$ and smooth $\nu/\mu$,
$1-\exp\bigl{(}-\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\bigr{)}\leq\frac{qC_{\mathsf{PI}}(\mu)}{4}\,\mathsf{RFI}_{q}(\nu\mathbin{\|}\mu)\,.$ ($\mathsf{R\text{-}PI}$)
Proof of Lemma 2.1.
We combine general de Bruijn’s identity and the Rényi PI. For the former, we note that $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$ since ($\mathsf{PI}$) ensures exponential integrability [BU83, GM83] (i.e., $\int\exp(s\,|\cdot|)\,\mathrm{d}\mu<\infty$ if $s^{2}<4/C_{\mathsf{PI}}(\mu)$; see [BGL14, Proposition 4.4.2]). For the latter, note that $\nu_{t}$ and $\mu_{t}$ are smooth due to an elementary property of the heat kernel.
Using those two with $\chi^{q}+1=\lVert\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\rVert_{L^{q}(\mu_{t})}^{q}:=L^{q}\geq 1$, we deduce that
$\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq-\frac{2}{qC_{\mathsf{PI}}(\mu_{t})}\,(q-1)\,L^{q}(1-L^{-\frac{q}{q-1}})\,.$
Since $(q-1)\cdot x\,(1-x^{-1/(q-1)})\geq x-1$ for $x\geq 1$, we have
$\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq-\frac{2}{qC_{\mathsf{PI}}(\mu_{t})}\,(L^{q}-1)=-\frac{2}{qC_{\mathsf{PI}}(\mu_{t})}\,\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\,.$
Using $C_{\mathsf{PI}}(\mu_{t})\leq C_{\mathsf{PI}}(\mu)+t$ and solving the differential inequality, we obtain that
$\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq\bigl{(}1+\frac{t}{C_{\mathsf{PI}}(\pi)}\bigr{)}^{-2/q}\,\chi^{q}(\nu\mathbin{\|}\mu)\,,$
which completes the proof. ∎
Remark 2.3 (Rényi version).
Under the same condition $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$, one can rederive (2.2) by following the argument in [CCSW22]. Using the identity $\mathcal{R}_{q}=\frac{1}{q-1}\,\log(1+\chi^{q})$ together with the chain rule and and de Bruijn’s identity, we obtain:
$\partial_{t}\mathcal{R}_{q}(\nu_{t}\mathbin{\|}\mu_{t})=\frac{1}{q-1}\,\frac{\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})}{L^{q}}=-\frac{1}{2}\,\mathsf{RFI}_{q}(\nu_{t}\mathbin{\|}\mu_{t})\,.$
When applying the Rényi PI, they used the inequality $21-e^{-x}\geq 1/$ 2 for $x\geq 1$ when $\mathcal{R}_{q}\geq 1$, and the inequality $21-e^{-x}\geq x/$ 2 for $x\leq 1$ when $\mathcal{R}_{q}\leq 1$. These inequalities are used to control the dissipation of the Rényi divergence in different regimes.
The mixing time of $\mathsf{PS}$ follows as a direct consequence of the contraction result above. Here, we ignore via the DPI additional contraction due to the backward step, although the backward step is also known to incur similar contraction under suitable regularity assumptions [CCSW22].
Proposition 2.4.
Let $q\geq 2$ and $\mathbb{E}_{\pi^{X}}[|\cdot|^{4}]<\infty$. Then, the law $\pi_{k}^{X}$ of $X_{k}$ returned by the $\mathsf{Proximal\ sampler}$ satisfies $\chi^{q}(\pi_{k+1}^{X}\mathbin{\|}\pi^{X})\leq\eta_{\chi^{q}}(\pi^{X},h)\,\chi^{q}(\pi_{k}^{X}\mathbin{\|}\pi^{X})$. Given $\varepsilon\in(0,1)$, we can ensure that $\chi^{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
$N\lesssim qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{\chi^{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})}{\varepsilon}\quad\text{iterations\,.}$
Moreover, $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
$N\lesssim q\,\bigl{(}1\vee h^{-1}C_{\mathsf{PI}}(\pi^{X})\bigr{)}\log\frac{M_{q}}{\varepsilon}\quad\text{iterations\,.}$
We adapt this result to the setting of log-concave probability measures.
Lemma 2.5.
For any $\varepsilon\in(0,1)$, $q\geq 2$, and log-concave probability measure $\pi^{X}$ over $\mathbb{R}^{d}$, we can ensure $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
$N\lesssim q\,\bigl{(}1\vee h^{-1}\lVert\operatorname{cov}\pi^{X}\rVert\log d\bigr{)}\log\frac{M_{q}}{\varepsilon}\quad\text{iterations\,.}$
Proof.
Any log-concave probability measures satisfy ($\mathsf{PI}$) (and thus $\mathbb{E}_{\pi^{X}}[|\cdot|^{4}]<\infty$). The claim then follows from the proposition above with $C_{\mathsf{PI}}(\pi^{X})\lesssim\lVert\operatorname{cov}\pi^{X}\rVert\log d$ for log-concave probability measures [Kla23]. ∎
Correctness under algorithmic modifications.
It may not be immediately clear whether the transition kernel of $\mathsf{PS}$ remains unchanged under two algorithmic modifications: (1) setting a retry threshold $\tau$, and (2) restarting an iteration if all $\tau$ retries fail.
For completeness, we prove that the transition kernel is indeed preserved under these changes. We first show that imposing the threshold does not alter the kernel of the backward step.
Lemma 2.6.
Upon success of the rejection sampling in the backward step, the accepted sample is distributed according to $\pi^{X|Y=y}$.
Proof.
Consider the proposal $\mu_{y}$ with success probability $p_{y}>0$ such that $X_{i}\sim\mu_{y}$ for each retry $i\in[\tau]$ satisfies $\operatorname{law}(X_{i}\,|\,\text{accepted})=\pi^{X|Y=y}$. Let $X$ be the random sample given by the backward step. For any event $A$ in $\mathbb{R}^{d}$, we have
$\mathbb{P}(X\in A\,|\,\text{accepted})$ $=\frac{\mathbb{P}(X\in A,\,\text{success})}{\mathbb{P}(\text{success})}$
$=\frac{\sum_{i=1}^{\tau}\mathbb{P}(\text{fail in the first }(i-1)\text{ trials},\,X_{i}\ \text{accepted},\,X_{i}\in A)}{\sum_{i=1}^{\tau}\mathbb{P}(\text{success in the }i\text{-th trial})}$
$=\frac{\sum_{i=1}^{\tau}(1-p_{y})^{i-1}\,\mathbb{P}(X_{i}\in A,\,X_{i}\ \text{accepted})}{\sum_{i=1}^{\tau}(1-p_{y})^{i-1}p_{y}}=\frac{\mathbb{P}(X_{1}\in A,\,X_{1}\ \text{accepted})}{p_{y}}$
$=\mathbb{P}(X_{1}\in A\,|\,\text{accepted})=\pi^{X|Y}(A)\,,$
which completes the proof. ∎
Restarting the forward step when $\tau$ retries fail also does not alter the transition kernel of $\mathsf{PS}$.
Lemma 2.7.
The transition kernel of $\mathsf{PS}$ is $P_{h}Q_{h}$.
Proof.
Let $X_{i}=x$ be a current point, and let $f_{x}$ be the failure probability of a stage. For any event $A$,
[…]
=$\underset{(i)}{=}\sum_{j=1}^{\infty}\int\pi^{X|Y=y}(A)\,\gamma_{h}(y-x)\,\mathrm{d}y\times f_{x}^{j-1}(1-f_{x})$
$=\int\pi^{X|Y=y}(A)\,\gamma_{h}(y-x)\,\mathrm{d}y=\int\pi^{X|Y=y}(A)\,(\delta_{x}P_{h})(y)\,\mathrm{d}y$
$=(\delta_{x}P_{h}Q_{h})(A)\,,$
where $(i)$ follows from Lemma 2.6. Thus, $X_{i+1}$ is distributed as $\pi_{i}^{X}P_{h}Q_{h}$ when $X_{i}\sim\pi_{i}^{X}$. ∎
2.2 Query complexity per iteration
We now focus on the case of $\pi^{X}\propto\mathds{1}_{\mathcal{K}}$, assuming $B_{1}(0)\subset\mathcal{K}$ by translation. Hereafter, we assume that the starting distribution $\pi_{0}^{X}$ is $M_{q}$-warm with respect to $\pi^{X}$. By the DPI, the same holds between $\pi_{1}^{Y}=\pi_{0}^{X}*\gamma_{h}$ and $\pi^{Y}=\pi^{X}*\gamma_{h}$ (i.e., $\lVert\mathrm{d}\pi_{1}^{Y}/\mathrm{d}\pi^{Y}\rVert_{L^{q}(\pi^{Y})}<\infty$), and this property is preserved throughout subsequent iterations (i.e., $\mathcal{R}_{q}(\pi_{k}^{X}\mathbin{\|}\pi^{X})\leq\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$). For brevity, we will use the $L^{q}$-warmness, $\chi^{q}$-warmness, or $\mathcal{R}_{q}$-warmness interchangeably, as justified by (1.2).
Recall from [KVZ24] a concentration phenomenon for $\pi^{Y}=\pi^{X}*\gamma_{h}$.
Lemma 2.8.
For a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ containing a unit ball,
$\pi^{Y}(\mathbb{R}^{d}\backslash\mathcal{K}_{\delta})\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+\delta d\bigr{)}\,,$
where $\mathcal{K}_{\delta}=\{x\in\mathbb{R}^{d}:d(x,\mathcal{K})\leq\delta\}$.
Taking $h=(2cd^{2})^{-1}$ and $\delta=1/d$ for some $c>0$ ensures $\pi^{Y}(\mathbb{R}^{d}\backslash\mathcal{K}_{\delta})\leq e^{-c+1}$. Note that the success probability of the rejection sampling at $Y_{k}=y$ is $\ell(y):=\mathbb{P}_{\mathcal{N}(y,hI_{d})}(\cdot\in\mathcal{K})$, so the failure probability of a stage in the $k$-th iteration is
$\int\bigl{(}1-\ell(y)\bigr{)}^{\tau}\,\mathrm{d}\pi_{k}^{Y}(y)=\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]=:f_{k}\,.$
Let $n_{k}$ denote the number of trials required to complete the $k$-th iteration. Then, the total expected number of queries over $N$ iterations is $\sum_{k=1}^{N}\mathbb{E}n_{k}$. By the law of total expectation,
$\mathbb{E}n_{k}=\sum_{i=1}^{\infty}\mathbb{E}[n_{k}\,|\,\text{success in }i\text{-th stage}]\,\mathbb{P}(\text{success in }i\text{-th stage})\leq\sum_{i=1}^{\infty}i\tau\times f_{k}^{i-1}(1-f_{k})=\frac{\tau}{1-f_{k}}\,.$
Therefore, bounding the per-iteration complexity reduces to bounding $f_{k}$ for each $k\in[N]$.
Let $p$ be the conjugate of $q$ (i.e., $p^{-1}+q^{-1}=1$ for $q\geq 2$). Then,
$\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq M_{q}\,\bigl{(}\mathbb{E}_{\pi^{Y}}[(1-\ell)^{\tau p}]\bigr{)}^{1/p}\,,$
and
$\int_{\mathbb{R}^{d}}(1-\ell)^{\tau p}\,\mathrm{d}\pi^{Y}=\int_{\mathbb{R}^{d}\backslash\mathcal{K}_{\delta}}\cdot+\int_{\mathcal{K}_{\delta}\cap[\ell\geq\tau^{-1}\log(3M_{q})]}\cdot+\int_{\mathcal{K}_{\delta}\cap[\ell<\tau^{-1}\log(3M_{q})]}\cdot=:\mathsf{A}+\mathsf{B}+\mathsf{C}\,.$
Using $\pi^{Y}=\ell/\operatorname{vol}\mathcal{K}$ when bounding $\mathsf{C}$ below,
$\mathsf{A}$ $\leq\pi^{Y}(\mathbb{R}^{d}\backslash\mathcal{K}_{\delta})\leq e^{-c+1}\,,$
$\mathsf{B}$ $\leq\int_{[\ell\geq\tau^{-1}\log(3M_{q})]}\exp(-\ell\tau p)\,\mathrm{d}\pi^{Y}\leq\bigl{(}\frac{1}{3M_{q}}\bigr{)}^{p},$
$\mathsf{C}$ $\leq\int_{\mathcal{K}_{\delta}\cap[\ell<\tau^{-1}\log(3M_{q})]}\frac{\ell(y)}{\operatorname{vol}\mathcal{K}}\,\mathrm{d}y\leq\frac{\log(3M_{q})}{\tau}\,\frac{\operatorname{vol}\mathcal{K}_{\delta}}{\operatorname{vol}\mathcal{K}}\leq\tau^{-1}e\log(3M_{q})\,,$
where we used $\operatorname{vol}\mathcal{K}_{\delta}\subset\operatorname{vol}((1+\delta)\,\mathcal{K})=(1+\delta)^{d}\,\operatorname{vol}\mathcal{K}\leq e\operatorname{vol}\mathcal{K}$ (due to $B_{1}(0)\subset\mathcal{K}$). Taking $c\asymp\log(3M_{q})^{p}$ and $\tau\asymp M_{q}^{p}\log(3M_{q})$, we can ensure that $\mathsf{A}+\mathsf{B}+\mathsf{C}\leq(\nicefrac{{1}}{{2M_{q}}})^{p}$ and thus $f_{k}\leq\frac{1}{2}$.
Putting these together, we bound the total query complexity during $N$ iterations as
$\sum_{k=1}^{N}\mathbb{E}n_{k}\leq\sum_{k=1}^{N}\frac{\tau}{1-f_{k}}\leq 2\tau N\,.$
Using this, we derive the complexity of $\mathsf{PS}_{\textup{unif}}$ with balanced guarantees.
Proof of Theorem 1.1.
For $M:=\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$, set $h^{-1}\asymp d^{2}M$ and $\tau\asymp Me^{M}$. To achieve $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$, by Lemma 2.5, it suffices to iterate at most
$N\lesssim q\,(1\vee h^{-1}\Lambda\log d)\log\frac{M}{\varepsilon}\lesssim q\Lambda d^{2}M\log d\log\frac{M}{\varepsilon}\quad\text{times}\,.$
Therefore, the expected number of total queries during $N$ iterations is
$2\tau N\lesssim q\Lambda d^{2}M^{2}e^{M}\log d\log\frac{M}{\varepsilon}=\widetilde{O}\bigl{(}q\Lambda d^{2}e^{\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})}\log\frac{1}{\varepsilon}\bigr{)}\,,$
which completes the proof. ∎
3 Hypercontractivity under the simultaneous heat flow
In this section, we establish hypercontractivity under simultaneous heat flow, which will serve as a stepping stone for relaying a Rényi warmness in annealing in the next section.
3.1 Adjoint of the heat semigroups and the proximal sampler
We first recap the Markov semigroup theory for a gentle introduction to heat semigroup and its adjoint. Interested readers may want to refer to [VH14, BGL14, Che25] for further details on the semigroup theory. Meanwhile, we take detour to $\mathsf{PS}$, since it sets up a context in which the introduction of heat semigroups and its adjoint is natural.
Markov semigroup theory.
Given a (time-homogeneous) Markov process $(X_{t})_{t\geq 0}$ on a domain $E$, its associated Markov semigroup $(P_{t})_{t\geq 0}$ is defined as a family of functional operators such that
$(P_{t}f)(x):=\mathbb{E}[f(X_{t})|X_{0}=x]\quad\text{for every bounded measurable function }f:E\to\mathbb{R}\,.$
One could readily check that $P_{t}$ is a linear operator with $P_{0}=\mathrm{id}$, $P_{t+s}=P_{t}\circ P_{s}$ for any $t,s\geq 0$ (semigroup property), $P_{t}\mathds{1}=\mathds{1}$ (mass conservation), and $P_{t}f\geq 0$ for any $f\geq 0$ (positivity preserving). Below, we omit $E$ in the integral sign.
A measure $\mu$ is said to be invariant (or stationary) for $P_{t}$ if $\int P_{t}f\,\mathrm{d}\mu=\int f\,\mathrm{d}\mu$ for every $t\geq 0$. In addition, it is said to be reversible for $P_{t}$ if $\int P_{t}f\,g\,\mathrm{d}\mu=\int f\,P_{t}g\,\mathrm{d}\mu$ for any $f,g\in L^{2}(\mu)$ and $t\geq 0$. For any $p\geq 1$, it follows from Jensen’s inequality that $P_{t}$ is a contraction operator on the bounded functions in $L^{p}(\mu)$ (i.e., $\lVert P_{t}f\rVert_{L^{p}(\mu)}\leq\lVert f\rVert_{L^{p}(\mu)}$). This furnishes the extension of $P_{t}$ to $L^{p}(\mu)$ by a density argument.
The proximal sampler.
As examined in [CCSW22] (with further details in [Che25]), the forward step (i.e., $y_{k+1}\sim\pi^{Y|X=x_{k}}=\mathcal{N}(x_{k},hI_{d})$) can be viewed as simulating Brownian motion for time $h$ started at $x_{k}$; when $\mathrm{d}X_{t}=\mathrm{d}B_{t}$ with $X_{0}=x_{k}$, we have $X_{h}\sim\operatorname{law}y_{k+1}$. The associated Markov semigroup of Brownian motion is often called the heat semigroup, which admits the formula of $P_{t}f=f*\gamma_{t}=:f_{t}$. Based on the classical heat equation $\partial_{t}P_{t}f=\frac{1}{2}\,\Delta P_{t}f$, rich theory for the heat semigroup facilitates one to understand the forward step in a better way (e.g., contraction under simultaneous heat flow).
This prompts a natural question whether one could also introduce a semigroup structure to the backward step. It turns out that the backward step can be interpreted as an SDE $(Z_{t}^{\leftarrow})_{t\geq 0}$, given by
$\mathrm{d}Z_{t}^{\leftarrow}=\nabla\log(\pi^{X}*\gamma_{h-t})\,(Z_{t}^{\leftarrow})\,\mathrm{d}t+\mathrm{d}\bar{B}_{t}\qquad\text{for another Brownian motion }\bar{B}_{t}\,,$
so that if $Z_{0}^{\leftarrow}\sim\pi^{Y}$, then $Z_{h}^{\leftarrow}\sim\pi^{X}$. Though it is non-trivial to justify, one could guess that $Z_{h}^{\leftarrow}\sim\pi^{X|Y=y}$ if $Z_{0}^{\leftarrow}=y$. Hence, one may be able to introduce a semigroup structure to the backward step via, for any suitable test function $f$,
$\overline{Q}_{h}f(y):=\mathbb{E}[f(Z_{h}^{\leftarrow})|Z_{0}^{\leftarrow}=y]=\mathbb{E}_{\pi^{X|Y=y}}f\,.$
As seen shortly, Klartag and Putterman [KP23] defined the same object as the $L^{2}(\pi)$-adjoint of the heat semigroups, providing solid ground where similar rigorous theory can be developed for the backward step.
$L^{2}$-adjoint of heat semigroups.
Given a probability measure $\pi$, [KP23] viewed the heat semigroup $P_{t}$ as a contraction operator from $L^{2}(\pi_{t})$ to $L^{2}(\pi)$:
$\lVert P_{t}f\rVert_{L^{2}(\pi)}^{2}=\int(P_{t}f)^{2}\,\pi\leq\int P_{t}(f^{2})\,\pi=\int f^{2}\,P_{t}\pi=\lVert f\rVert_{L^{2}(\pi_{t})}^{2}\,,$
where the equality in the middle follows simply from reversibility of the heat semigroups. This perspective led them to define the adjoint operator $Q_{t}:L^{2}(\pi)\to L^{2}(\pi_{t})$ as $Q_{0}=\mathrm{id}$ and
$Q_{t}f=\frac{P_{t}(f\pi)}{P_{t}\pi}=\frac{P_{t}(f\pi)}{\pi_{t}}\quad\text{for }t>0\,.$
This heat adjoint could be viewed as the backward step of $\mathsf{PS}$ with $\pi^{X}=\pi$ and $t=h$, since
$\overline{Q}_{t}f(y)=\int f\,\mathrm{d}\pi^{X|Y=y}=\frac{\int f(x)\,\pi^{X}(x)\,\gamma_{t}(y-x)\,\mathrm{d}x}{\int\pi^{X}(x)\,\gamma_{t}(y-x)\,\mathrm{d}x}=\frac{(f\pi)*\gamma_{t}}{\pi*\gamma_{t}}=\frac{P_{t}(f\pi)}{\pi_{t}}=(Q_{t}f)(y)\,.$
It follows from this formula that the adjoint operator satisfies mass preservation ($Q_{t}\mathds{1}=\mathds{1}$) and positivity preserving ($Q_{t}f\geq 0$ if $f\geq 0$). Moreover, it is a contraction operator; for $f\in L^{2}(\pi)\backslash\{0\}$, by Cauchy-Schwarz
$(Q_{t}f)^{2}\,\pi_{t}=\frac{\bigl{(}P_{t}(f\pi)\bigr{)}^{2}}{\pi_{t}}\leq\frac{P_{t}(f^{2}\pi)\,P_{t}\pi}{\pi_{t}}=P_{t}(f^{2}\pi)\,.$
and integration of both sides yields $\lVert Q_{t}f\rVert_{L^{2}(\pi_{t})}^{2}\leq\lVert f\rVert_{L^{2}(\pi)}^{2}$.
3.2 Hypercontractivity of an adjoint of the heat semigroups
We now show something stronger than mere contraction of $Q_{t}$ from $L^{2}(\pi)$ to $L^{2}(\pi_{t})$, called hypercontractivity. To begin with, we recall hypercontractivity of Markov semigroups under ($\mathsf{LSI}$):
Proposition 3.1 (Gross).
For $C>0$ and $p\in(1,\infty)$, let $q(t)=1+(p-1)\,e^{2t/C}$ for $t\geq 0$, and $P_{t}$ be a reversible Markov semigroup with stationary measure $\nu$. Then, $\nu$ satisfies ($\mathsf{LSI}$) with parameter $C$ if and only if
$\lVert P_{t}f\rVert_{L^{q(t)}(\nu)}\leq\lVert f\rVert_{L^{p}(\nu)}\quad\text{for all }p\in(1,\infty),t\geq 0,\text{ and }f\in L^{p}(\nu)\,.$
It means that when a stationary measure satisfies ($\mathsf{LSI}$), the semigroup $P_{t}$ not only contracts in any $L^{p}$-spaces but also improves integrability (recall $L^{q}\subset L^{p}$ for $p<q$ due to monotonicity of $\lVert\cdot\rVert_{L^{p}}$ in $p$). We now prove Theorem 1.2, hypercontractivity under the simultaneous heat flow with different $q(t)=1+(p-1)\,(1+\nicefrac{{t}}{{C_{\mathsf{LSI}}(\nu)}})$;
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f\rVert_{L^{p}(\nu)}\quad\text{for all }p\in(1,\infty),t\geq 0,\text{ and }f\in L^{p}(\nu)\,.$
Note that the reference measure $\nu_{t}$ also evolves over time $t$.
Proof of Theorem 1.2.
We first prove the claim for $f\in C_{c,+}^{\infty}(\mathbb{R}^{d}):=\{f+c:f\in C_{c}^{\infty}(\mathbb{R}^{d}),f\geq 0,c>0\}$. Let $\bar{f}$ be the normalization of $f$ with respect to $\nu$ (i.e., $\bar{f}=f/\int f\,\mathrm{d}\nu$).
Denoting a new probability measure $\mu$ by $\mathrm{d}\mu=\bar{f}\,\mathrm{d}\nu$, we define $\bar{f}_{t}:=Q_{t}\bar{f}=\mathrm{d}\mu_{t}/\mathrm{d}\nu_{t}$ and
$\mathcal{F}(t,q):=\int\bigl{(}\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr{)}^{q}\,\mathrm{d}\nu_{t}=\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\,,\qquad\mathcal{G}(t):=\frac{1}{q(t)}\,\log\mathcal{F}\bigl{(}t,q(t)\bigr{)}\,.$
It then suffices to show that $\partial_{t}\mathcal{G}\leq 0$ for a suitable choice of $q(t)$. Note that
$\partial_{t}\mathcal{G}=-\frac{q^{\prime}(t)}{q^{2}(t)}\,\log\mathcal{F}+\frac{\partial_{t}\mathcal{F}+q^{\prime}(t)\,\partial_{q}\mathcal{F}}{q(t)\,\mathcal{F}}\,.$
With assumptions of $q(t),q^{\prime}(t)>0$ for a moment, $\partial_{t}\mathcal{G}\leq 0$ is equivalent to
$\frac{1}{q^{\prime}(t)}\,\partial_{t}\mathcal{F}+\partial_{q}\mathcal{F}-\frac{1}{q(t)}\,\mathcal{F}\log\mathcal{F}\leq 0\,.$
We now compute $\partial_{t}\mathcal{F}$ and $\partial_{q}\mathcal{F}$. By general de Bruijn’s identity in Proposition 2.2 (note that $\mathbb{E}_{\nu}[|\cdot|^{4}]<\infty$ since ($\mathsf{LSI}$) implies ($\mathsf{PI}$)),
$\partial_{t}\mathcal{F}=-\frac{q\,(q-1)}{2}\,\mathbb{E}_{\nu_{t}}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr{)}^{q-2}\,\bigl{|}\nabla\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr{|}^{2}\bigr{]}=-\frac{q\,(q-1)}{2}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]\,.$
As for $\partial_{q}\mathcal{F}$, we have
$\partial_{q}\mathcal{F}$ $=\partial_{q}\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\underset{(i)}{=}\int\bar{f}_{t}^{q}\log\bar{f}_{t}\,\mathrm{d}\nu_{t}=\frac{1}{q}\,\Bigl{(}\operatorname{Ent}_{\nu_{t}}(\bar{f}_{t}^{q})+\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\,\log\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\Bigr{)}$
$\leq\frac{1}{q}\,\Bigl{(}2C_{\mathsf{LSI}}(\nu_{t})\,\mathbb{E}_{\nu_{t}}[|\nabla(\bar{f}_{t}^{q/2})|^{2}]+\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\,\log\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\Bigr{)}$
$=\frac{1}{q}\,\bigl{(}\frac{q^{2}C_{\mathsf{LSI}}(\nu_{t})}{2}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]+\mathcal{F}\log\mathcal{F}\bigr{)}\,.$
Differentiation under the integral sign in $(i)$ can be justified as follows: $\partial_{q}\bar{f}_{t}^{q}=(Q_{t}\bar{f})^{q}\log Q_{t}\bar{f}$ is continuous in $q$, and we can take a small $\delta>0$ such that for any $q^{\prime}\in(q-\delta,q+\delta)$ with $q-\delta>1$,
[…]
Hence, it amounts to showing integrability of the RHS. To this end, we first show that $\bar{f}_{t}\in L^{r}(\nu_{t})$ for any $r>1$. By Jensen’s inequality,
$\lVert\bar{f}_{t}\rVert_{L^{r}(\nu_{t})}^{r}=\mathbb{E}_{\nu_{t}}[(Q_{t}\bar{f})^{r}]\leq\mathbb{E}_{\nu_{t}}[Q_{t}(\bar{f}^{r})]=\mathbb{E}_{\nu}[\bar{f}^{r}]=\lVert\bar{f}\rVert_{L^{r}(\nu)}^{r}\,,$
and $\bar{f}\in C_{c,+}^{\infty}(\mathbb{R}^{d})$ ensures that $\lVert\bar{f}\rVert_{L^{r}(\nu)}<\infty$. Thus, integrability of the RHS follows from $L^{s}\subset L^{r}\log L$ for any $r<s$.
Combining the bounds on $\partial_{t}\mathcal{F}$ and $\partial_{q}\mathcal{F}$,
$\frac{1}{q^{\prime}(t)}\,\partial_{t}\mathcal{F}+\partial_{q}\mathcal{F}-\frac{1}{q(t)}\,\mathcal{F}\log\mathcal{F}$
$\leq-\frac{q\,(q-1)}{2\,q^{\prime}(t)}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]+\frac{1}{q}\,\bigl{(}\frac{q^{2}C_{\mathsf{LSI}}(\nu_{t})}{2}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]+\mathcal{F}\log\mathcal{F}\bigr{)}-\frac{1}{q(t)}\,\mathcal{F}\log\mathcal{F}$
$=\bigl{(}C_{\mathsf{LSI}}(\nu_{t})-\frac{q-1}{q^{\prime}(t)}\bigr{)}\times\frac{q}{2}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]\,.$
Hence, it suffices to enforce
$\frac{(q-1)^{\prime}}{q-1}\leq\frac{1}{C_{\mathsf{LSI}}(\nu_{t})}\,.$
As $C_{\mathsf{LSI}}(\nu_{t})\leq C_{\mathsf{LSI}}(\nu)+t$, we can guarantee this by solving
$\frac{(q-1)^{\prime}}{q-1}\leq\frac{1}{C_{\mathsf{LSI}}(\nu)+t}\,.$
Integrating both sides over $[0,t]$, and setting $q(0)=p$,
$\log\frac{q(t)-1}{p-1}\leq\log\frac{C_{\mathsf{LSI}}(\nu)+t}{C_{\mathsf{LSI}}(\nu)}\,,$
and thus we can simply take
$\frac{q(t)-1}{p-1}=1+\frac{t}{C_{\mathsf{LSI}}(\nu)}\,,$
which clearly satisfies $q,q^{\prime}\geq 0$. Thus, $\lVert\bar{f}_{t}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert\bar{f}\rVert_{L^{p}(\nu)}$ (so $\lVert f_{t}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f\rVert_{L^{p}(\nu)}$), and hypercontractivity holds on $C_{c,+}^{\infty}(\mathbb{R}^{d})$.
Every $f\in C_{c,+}^{\infty}(\mathbb{R}^{d})$ can be written as $f=f_{\delta}=g+\delta$ for some non-negative $g\in C_{c}^{\infty}(\mathbb{R}^{d})$ and $\delta>0$. We then apply the dominated convergence theorem to $\lVert Q_{t}f_{\delta}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f_{\delta}\rVert_{L^{p}(\nu)}$ as $\delta\to 0$, extending the main claim from $C_{c,+}^{\infty}(\mathbb{R}^{d})$ to $C_{c,\geq 0}^{\infty}(\mathbb{R}^{d}):=C_{c}^{\infty}(\mathbb{R}^{d})\cap\{f\geq 0\}$.
Now, for any non-negative $f\in L^{p}(\nu)$, we can take a sequence $\{f_{n}\}$ of functions in $C_{c,\geq 0}^{\infty}(\mathbb{R}^{d})$ such that $f_{n}\to f$ in $L^{p}$. Moreover, we can take a subsequence $\{n_{k}\}_{k}$ such that $f_{n_{k}}\to f$ almost everywhere. Since $Q_{t}f_{n}\geq 0$ due to positivity-preserving, Fatou’s lemma implies that
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}^{q(t)}=\int\liminf_{k\to\infty}(Q_{t}f_{n_{k}})^{q(t)}\,\mathrm{d}\nu_{t}\leq\liminf_{k\to\infty}\int(Q_{t}f_{n_{k}})^{q(t)}\,\mathrm{d}\nu_{t}\,,$
and thus,
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}\leq\liminf_{k\to\infty}\,\lVert Q_{t}f_{n_{k}}\rVert_{L^{q(t)}(\nu_{t})}\leq\liminf_{k\to\infty}\,\lVert f_{n_{k}}\rVert_{L^{p}(\nu)}=\lVert f\rVert_{L^{p}(\nu)}\,.$
Lastly, for any $f\in L^{p}(\nu)$, note that $|f|=f_{+}+f_{-}$ is also in $L^{p}(\nu)$. Thus,
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}=\lVert|Q_{t}f|\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert Q_{t}(|f|)\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert|f|\rVert_{L^{p}(\nu)}=\lVert f\rVert_{L^{p}(\nu)}\,,$
which finishes the proof. ∎
For $f:=\mathrm{d}\mu/\mathrm{d}\nu\in L^{p}(\nu)$ (which is equivalent to $\mathcal{R}_{p}(\mu\mathbin{\|}\nu)<\infty$), hypercontractivity of $Q_{t}$ implies that
$\Bigl{\|}\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\Bigr{\|}_{L^{q(t)}(\nu_{t})}=\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f\rVert_{L^{p}(\nu)}=\Bigl{\|}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{\|}_{L^{p}(\nu)}\,.$
Then, $\mathcal{R}_{q(t)}(\mu_{t}\mathbin{\|}\nu_{t})\leq\mathcal{R}_{p}(\mu\mathbin{\|}\nu)$ follows from
$\mathcal{R}_{p}(\mu\mathbin{\|}\nu)=\frac{p}{p-1}\log\,\Bigl{\|}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{\|}_{L^{p}(\nu)}\geq\frac{p}{p-1}\,\log\,\Bigl{\|}\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\Bigr{\|}_{L^{q(t)}(\nu_{t})}\geq\frac{q(t)}{q(t)-1}\,\log\,\Bigl{\|}\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\Bigr{\|}_{L^{q(t)}(\nu_{t})}\,.$
This looks similar to the DPI for Rényi divergence, but we remark that the Rényi order of the RHS is larger that that of the LHS.
Improved convergence rate of PS under LSI.
Just as in the SDPI under ($\mathsf{PI}$), the SDPI under ($\mathsf{LSI}$) was established in [CCSW22, KO25]: for $q>1$ and two probability measures $\mu,\nu$ with $\mu\ll\nu$,
$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{LSI}}(\nu))^{1/q}}\,.$ (3.1)
Combining this contraction with hypercontractivity, we can derive a stronger mixing result of $\mathsf{PS}$: when $\pi_{k}$ denoting the law of the $k$-th sample returned by $\mathsf{PS}$ with initial $\pi_{0}$ and target $\pi$, for any given $\varepsilon>0$ and $p,q>1$, we have $\mathcal{R}_{q}(\pi_{N}\mathbin{\|}\pi)\leq\varepsilon$ after $N\lesssim h^{-1}C_{\mathsf{LSI}}(\pi)\log\frac{q\mathcal{R}_{2\wedge p}(\pi_{0}\mathbin{\|}\pi)}{(p-1)\,\varepsilon}$ iterations.
Proof of Corollary 1.3.
We first iterate $\mathsf{PS}$ $N_{0}\asymp(2\wedge p)\,h^{-1}C_{\mathsf{LSI}}(\pi)\log\frac{\mathcal{R}_{(2\wedge p)}(\pi_{0}\mathbin{\|}\pi)}{\varepsilon}$ times to bring $\pi_{N_{0}}$ within $\varepsilon$-distance to $\pi$ in $\mathcal{R}_{2\wedge p}$. Next, we iterate it $N_{1}\asymp h^{-1}C_{\mathsf{LSI}}(\pi)\log_{2}\frac{q-1}{1\wedge(p-1)}$ more times, where we invoke DPI and hypercontractivity each iteration as follows: for $i\leq N_{1}$ and $\pi_{i}^{X}=\pi_{i}$,
$\mathcal{R}_{q_{i+1}}(\pi_{N_{0}+i+1}\mathbin{\|}\pi)\leq\mathcal{R}_{q_{i+1}}(\pi_{N_{0}+i+1}^{Y}\mathbin{\|}\pi^{Y})\leq\mathcal{R}_{q_{i}}(\pi_{N_{0}+i}\mathbin{\|}\pi)\,,$
where $(q_{i+1}-1)/(q_{i}-1)=1+h/C_{\mathsf{LSI}}(\pi)$. Then,
$\frac{q_{N_{1}}-1}{1\wedge(p-1)}=\frac{q_{N_{1}}-1}{q_{0}-1}=\prod_{i=1}^{N_{1}}\frac{q_{i}-1}{q_{i-1}-1}=\bigl{(}1+\frac{h}{C_{\mathsf{LSI}}(\pi)}\bigr{)}^{N_{1}}\geq\frac{q-1}{1\wedge(p-1)}\,.$
Therefore, $\mathsf{PS}$ suffices to iterate $N=N_{0}+N_{1}\lesssim h^{-1}C_{\mathsf{LSI}}(\pi)\log\frac{q\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}{(1\wedge(p-1))\,\varepsilon}$ times to attain the desired guarantee. ∎
As a simple corollary, this result leads to a mixing time of $\mathsf{PS}$ in the well-conditioned setting, which improves the linear dependence of $\mathsf{PS}$ on $q$ [CCSW22].
Proposition 3.2 (Zeroth-order sampling in well-conditioned setting).
Let $\pi\propto e^{-V}$ with $\alpha$-strongly convex and $\beta$-smooth $V$ over $\mathbb{R}^{d}$, given access to an evaluation and proximal oracle for $V$. Let $\pi_{k}$ denote the law of the $k$-th sample returned by $\mathsf{PS}$. Given $q>1$ and $\varepsilon>0$, we can attain $\mathcal{R}_{q}(\pi_{N}\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(\kappa d\log\frac{q}{\varepsilon})$ queries to both oracles, where $\kappa=\beta/\alpha$.
Proof.
One can find an initial distribution $\pi_{0}$ such that $\mathcal{R}_{\infty}(\pi_{0}\mathbin{\|}\pi)\leq d\log\kappa$ (see [LST21, Lemma 18]). Using $\mathcal{R}_{2}\leq\mathcal{R}_{\infty}$, $C_{\mathsf{LSI}}(\pi)\leq\alpha^{-1}$, and Corollary 1.3 with $q_{0}=2$, we can conclude that $\mathsf{PS}$ suffices to iterate $\widetilde{O}(\kappa d\log\frac{q}{\varepsilon})$ times. It is also known from [LST21] that the backward step can be implemented via rejection sampling for step size $h=\nicefrac{{1}}{{\beta d}}$, using $O(1)$ many evaluation queries (in expectation) and one proximal query per iteration. Combining these together finishes the proof. ∎
4 Order-preserving annealing for simpler warm-start generation
In this section, we streamline prior warm-start generation algorithms and their proofs. Above all, we interweave an overall analysis with hypercontractivity under the heat flow, eventually establishing $\mathcal{R}_{q}$-warmness.
Simpler warm-start generation.
Let $\mathcal{K}$ be a convex body with $B_{1}(0)\subset\mathcal{K}$ (by translation), $\mathbb{E}_{\pi}[|\cdot|^{2}]\leq R^{2}$, and $\Lambda=\lVert\operatorname{cov}\pi\rVert$. We simplify prior annealing approaches in [KV25a], moving along a sequence $\{\mu_{i}\}_{i\in[m]}$ of distributions with $\mu_{i}=\gamma_{\sigma_{i}^{2}}|_{\mathcal{K}}=\pi\gamma_{\sigma_{i}^{2}}$.
This prior algorithm consists of the following several steps: (i) Preprocessing: Truncate $\mathcal{K}$ to the ball of radius $O(R)$ such that $\pi(\bar{\mathcal{K}})\geq 0.99$ where $\bar{\mathcal{K}}:=\mathcal{K}\cap B_{O(R)}(0)$, (ii) Initialization: Sample from $\mu_{1}=\gamma_{d^{-1}}|_{\bar{\mathcal{K}}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$, (iii) Annealing: While $\sigma_{i}^{2}\lesssim R^{2}$, use $\mathsf{PS}$ to sample from $\mu_{i+1}$ with initial distribution $\mu_{i}$, where $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,(1+\frac{\sigma_{i}}{q^{1/2}R})$ for $q=\operatorname{polylog}\nicefrac{{dR}}{{\varepsilon}}$. In particular, there are two modes: (iii-1) normal mode: when $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}$, iterate $\mathsf{PS}$ $\widetilde{O}(d^{2}\sigma_{i}^{2})$ times, (iii-2) faster mode: when $\sigma_{i}^{2}\gtrsim qR\Lambda^{1/2}$, iterate $\mathsf{PS}$ $\widetilde{O}(d^{2}R\Lambda^{1/2})$ times. (iv) Termination: Run $\mathsf{PS}$ to sample from $\pi$ with initial distribution $\gamma_{R^{2}}|_{\bar{\mathcal{K}}}$.
We show that the preprocessing step is redundant and also that the faster mode can be replaced by early-stopping: using $\mathsf{PS}_{\textup{Gauss}}$4.1) below,
  1. 1.
    Sample from $\mu_{1}=\pi\gamma_{d^{-1}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$.
  2. 2.
    While $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{Gauss}}$ such that $\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl{(}1+\frac{\sigma_{i}}{R}\bigr{)}\,.$
We need two ingredients to obtain a final guarantee: (1) the query complexity of $\mathsf{PS}$ for truncated Gaussians and (2) closeness of two annealing distributions. The former was already studied in previous work, but it will be further simplified in §4.1 as done for uniform sampling in an earlier section. The latter was proven for $\chi^{2}$-divergence by a localization argument [CV18] and then was extended to $\mathcal{R}_{q}$ in [KV25a]. We show in §4.2 that this is in fact a simple consequence of ($\mathsf{LSI}$). Combining these together, we establish a query complexity of generating $\mathcal{R}_{q}$-warmness.
4.1 Proximal sampler for truncated Gaussian
We study $\mathsf{PS}$ for $\mu^{X}:=\pi\gamma_{\sigma^{2}}$, Gaussian distributions truncated to a convex body $\mathcal{K}$ containing $B_{1}(0)$. Its prototype was already studied in [KZ25, KV25a]. The streamlined algorithm, denoted by $\mathsf{PS}_{\textup{Gauss}}$, can be readily deduced from the general framework of $\mathsf{PS}$.
Starting with $x_{0}\sim\mu_{0}^{X}$, step size $h>0$, threshold $\tau$, and $s:=\sigma^{2}/(h+\sigma^{2})\in(0,1)$,
  • [Forward] $y_{k+1}\sim\mu^{Y|X=x_{k}}=\gamma_{h}(\cdot-x_{k})$.
[…]
As in the uniform-sampling case, its analysis is split into two parts: a mixing analysis and per-iteration complexity. The per-iteration analysis will be proceeded in a similar fashion, but the mixing analysis can be improved due to the hypercontractivity of $\mathsf{PS}$.
Mixing analysis.
We already established the convergence rate of $\mathsf{PS}$ under ($\mathsf{LSI}$) (see Corollary 1.3). It is classical that $C_{\mathsf{LSI}}(\mu^{X})\leq\sigma^{2}$ (e.g., use the Bakry–Émery criterion, and the convex boundary term does not harm in satisfying this criterion; see [Kla09, BJ24]). Thus, in order to be $\varepsilon$-close to $\mu^{X}$ in $\mathcal{R}_{q}$, for $q_{0}>1$, $\mathsf{PS}_{\textup{Gauss}}$ suffices to iterate
$N\lesssim h^{-1}\sigma^{2}\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{X}\mathbin{\|}\mu^{X})}{(2\wedge q_{0}-1)\,\varepsilon}\qquad\text{iterations}\,.$
Per-step query complexity.
Denoting by $\ell(y)=\mathbb{P}_{\mathcal{N}(sy,shI_{d})}(\cdot\in\mathcal{K})$ the success probability of the rejection sampling, we can readily check that
$\mu^{Y}(y)=(\mu^{X}*\gamma_{h})(y)=\frac{s^{d/2}\ell(y)\exp(-\frac{1}{2s\sigma^{2}}\,|sy|^{2})}{\int_{\mathcal{K}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2})\,\mathrm{d}x}\,,$
and also recall a concentration phenomenon of $\mu^{Y}$ from [KZ25]: for $R=s^{-1}\mathcal{K}_{\delta}$, if $\delta\geq hd$,
$\mu^{Y}(\mathbb{R}^{d}\backslash R)\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+\delta d+hd^{2}\bigr{)}\,.$
Taking $h=(cd^{2})^{-1}$ and $\delta=d^{-1}$ for some $c\geq 1$, we have $\mu^{Y}(\mathbb{R}^{d}\backslash R)\leq e^{-c/2+2}$. Just as we used the bound of $\operatorname{vol}\mathcal{K}_{\delta}/\operatorname{vol}\mathcal{K}\leq e^{\delta d}$, we need the following “expansion” lemma (see [KZ25, Lemma 24]):
$s^{-d/2}\int_{\mathcal{K}_{\epsilon}}\exp\bigl{(}-\frac{1}{2s\sigma^{2}}\,|z|^{2}\bigr{)}\,\mathrm{d}z\leq 2e^{hd^{2}+\epsilon d}\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2\sigma^{2}}\,|z|^{2}\bigr{)}\,\mathrm{d}z\,.$
This leads to the following bound: by change of variables via $y=s^{-1}v$ below,
$\frac{\int_{R}s^{d/2}\exp(-\frac{1}{2s\sigma^{2}}\,|sy|^{2})\,\mathrm{d}y}{\int_{\mathcal{K}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2})\,\mathrm{d}x}=\frac{\int_{\mathcal{K}_{\delta}}s^{-d/2}\exp(-\frac{1}{2s\sigma^{2}}\,|v|^{2})\,\mathrm{d}v}{\int_{\mathcal{K}}\exp(-\frac{1}{2\sigma^{2}}\,|v|^{2})\,\mathrm{d}v}\leq 2e^{hd^{2}+\delta d}\leq e^{3}\,.$ (4.1)
We show that the failure probability of one stage in the $k$-th iteration is at most $1/2$ under suitable choices of $h$ and $\tau$. For a conjugate pair $(p,q)$ with $p,q>1$, we have $\mathbb{E}_{\mu_{k}^{Y}}[(1-\ell)^{\tau}]\leq M_{q}\,\{\mathbb{E}_{\mu^{Y}}[(1-\ell)^{\tau p}]\}^{1/p}$ and
$\int_{\mathbb{R}^{d}}(1-\ell)^{\tau p}\,\mathrm{d}\mu^{Y}=\int_{\mathbb{R}^{d}\backslash R}\cdot+\int_{R\cap[\ell\tau\geq\log(3M_{q})]}\cdot+\int_{R\cap[\ell\tau<\log(3M_{q})]}\cdot=:\mathsf{A}+\mathsf{B}+\mathsf{C}\,.$
Under $c\asymp\log(3M_{q})^{p}$ and $\tau\asymp M_{q}^{p}\log(3M_{q})$, this can be bounded by $(\nicefrac{{1}}{{2M_{q}}})^{p}$ due to
$\mathsf{A}$ $\leq\mu^{Y}(\mathbb{R}^{d}\backslash\mathcal{K}_{\delta})\leq e^{-c/2+2}\,,$
$\mathsf{B}$ $\leq\int_{R\cap[\ell\tau\geq\log(3M_{q})]}e^{-\ell\tau p}\,\mathrm{d}\mu^{Y}\leq\bigl{(}\frac{1}{3M_{q}}\bigr{)}^{p}\,,$
$\mathsf{C}$ $\leq\int_{R\cap[\ell\tau<\log(3M_{q})]}\mathrm{d}\mu^{Y}=\int_{R\cap[\ell\tau<\log(3M_{q})]}\frac{s^{d/2}\ell(y)\exp(-\frac{1}{2s\sigma^{2}}\,|sy|^{2})}{\int_{\mathcal{K}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2})\,\mathrm{d}x}$
[…]
Thus, $\mathbb{E}_{\mu_{k}^{Y}}[(1-\ell)^{\tau}]\leq 1/2$, and the expected number of trials during the $k$-th iteration is bounded by $2\tau$ as in §2.2. Combining all these together, we can prove Theorem 1.4 on the query complexity of $\mathsf{PS}_{\textup{Gauss}}$.
Proof of Theorem 1.4.
For $M:=\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{X}\mathbin{\|}\mu^{X})$, we set $h^{-1}\asymp d^{2}M$ and $\tau\asymp Me^{M}$. As mentioned in the mixing analysis, $\mathsf{PS}_{\textup{Gauss}}$ returns a sample $X_{N}$ such that $\mathcal{R}_{q}(\operatorname{law}X_{N}\mathbin{\|}\mu^{X})\leq\varepsilon$ by iterating
$N\lesssim Md^{2}\sigma^{2}\log\frac{qM}{(2\wedge q_{0}-1)\,\varepsilon}\quad\text{times}\,.$
Thus, the expected number of total queries during $N$ iterations is
$2\tau N\lesssim M^{2}e^{M}d^{2}\sigma^{2}\log\frac{qM}{(2\wedge q_{0}-1)\,\varepsilon}=\widetilde{O}\bigl{(}d^{2}\sigma^{2}\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{X}\mathbin{\|}\mu^{X})}\log\frac{q}{(2\wedge q_{0}-1)\,\varepsilon}\bigr{)}\,,$
which completes the proof. ∎
4.2 Functional inequalities to closeness and early-stopping
We will show that two consecutive Gaussians are $O(1)$-close in $\mathcal{R}_{q}$, finding a connection to ($\mathsf{LSI}$). We then show that annealing can be stopped around $\sigma_{i}^{2}\asymp qR\Lambda^{1/2}\log^{1/2}d$, much earlier than $\sigma_{i}^{2}\asymp R^{2}$, finding a connection to ($\mathsf{PI}$). Moreover, the second result improves a mathematical result established in [KV25a] as an immediate corollary. We then prove Theorem 1.5 on the query complexity of our annealing algorithm.
Mollification.
To avoid any regularity issues, we work with mollifications of log-concave distributions. As detailed in [Eva10, §C.4], the standard mollifier $\eta\in C^{\infty}(\mathbb{R}^{d})$ is defined as
$\eta(x):=\begin{cases}C\exp\bigl{(}\frac{1}{\lVert x\rVert^{2}-1}\bigr{)}&\lVert x\rVert\leq 1\,,\\ 0&\lVert x\rVert>1\,,\end{cases}$ (4.2)
where the constant $C$ is chosen to ensure $\int\eta=1$. For $\varepsilon>0$, one can consider
$\eta_{\varepsilon}(x):=\frac{1}{\varepsilon^{d}}\,\eta\bigl{(}\frac{x}{\varepsilon}\bigr{)}\,,$
whose support is bounded by $B_{\varepsilon}(0)$. Then, for locally integrable $f:\mathbb{R}^{d}\to\mathbb{R}$, its mollification $f_{\varepsilon}:=f*\eta_{\varepsilon}$ satisfies $f_{\varepsilon}\in C^{\infty}(\mathbb{R}^{d})$ and $f_{\varepsilon}\to f$ almost everywhere as $\varepsilon\to 0$.
Streamlined analysis.
Recall Scheffé’s lemma that if a sequence $\pi_{n}$ of probability densities converges to a probability measure $\pi$ almost everywhere, then $\pi_{n}$ converges to $\pi$ in $\mathsf{TV}$-distance (thus weakly). Hence, $\pi_{t}:=\pi*\eta_{t}\to\pi$ weakly as $t\to 0$ due to the property of the mollifier. Then, it is a property of weak convergence that $\int\gamma_{\sigma^{2}}\,\mathrm{d}\pi_{t}\to\int\gamma_{\sigma^{2}}\,\mathrm{d}\pi$ for any Gaussian density $\gamma_{\sigma^{2}}$. Therefore, probability measures $\pi_{t}\gamma_{\sigma^{2}}$ converge to $\pi\gamma_{\sigma^{2}}$ almost surely (and thus weakly).
A Rényi version of LSI was established in [VW23, Lemma 5] that for $q>1$ and smooth $\mu/\nu$,
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{qC_{\mathsf{LSI}}(\nu)}{2}\,\mathsf{RFI}_{q}(\mu\mathbin{\|}\nu)\,,$ ($\mathsf{R\text{-}LSI}$)
where $(\nicefrac{{\mu}}{{\nu}})^{q}\log\nicefrac{{\mu}}{{\nu}}$ is continuous, locally uniform, and integrable.
Lemma 4.1 (Closeness).
Let $\pi$ be a log-concave probability measure over $\mathbb{R}^{d}$ with $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$. For $\alpha,\sigma^{2}>0$ and $q>1$, probability measures $\pi\gamma_{\sigma^{2}}$ and $\pi\gamma_{\sigma^{2}(1+\alpha)}$ satisfy
$\mathcal{R}_{q}(\pi\gamma_{\sigma^{2}}\mathbin{\|}\pi\gamma_{\sigma^{2}(1+\alpha)})\leq\frac{q^{2}\alpha^{2}R^{2}}{2\sigma^{2}}\,.$
In particular, they are $O(1)$-close in $\mathcal{R}_{q}$ if $\alpha\leq\nicefrac{{\sigma}}{{qR}}$.
Proof.
Let $\mu\propto\pi\gamma_{\sigma^{2}}$ and $\nu\propto\pi\gamma_{\sigma^{2}(1+\alpha)}$, which clearly satisfies $\mu\ll\nu$. For $t>0$, consider mollified probability measures $\mu_{t}\propto\pi_{t}\gamma_{\sigma^{2}}$ and $\nu_{t}\propto\pi_{t}\gamma_{\sigma^{2}(1+\alpha)}$. On $\operatorname{supp}\nu_{t}$,
$\nabla\log\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}=\nabla\log\frac{\gamma_{\sigma^{2}}}{\gamma_{\sigma^{2}(1+\alpha)}}=\frac{\alpha x}{(1+\alpha)\,\sigma^{2}}\,,$
which admits a smooth extension (as $\frac{\alpha x}{(1+\alpha)\sigma^{2}}$) to $\mathbb{R}^{d}$. Then, for some $\delta>0$ independent of $t$,
$\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\nu_{t})=q\,\frac{\mathbb{E}_{\nu_{t}}[(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}})^{q}\,|\nabla\log\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}|^{2}]}{\mathbb{E}_{\nu_{t}}[(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}})^{q}]}=\frac{q\alpha^{2}}{\sigma^{4}\,(1+\alpha)^{2}}\,\frac{\mathbb{E}_{\nu_{t}}[(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}})^{q}\,|\cdot|^{2}]}{\mathbb{E}_{\nu_{t}}[(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}})^{q}]}=\frac{q\alpha^{2}}{\sigma^{4}\,(1+\alpha)^{2}}\,\frac{\mathbb{E}_{\pi_{t}}[\gamma_{\delta}\,|\cdot|^{2}]}{\mathbb{E}_{\pi_{t}}\gamma_{\delta}}\,.$ (4.3)
Using ($\mathsf{R\text{-}LSI}$) with $C_{\mathsf{LSI}}(\nu_{t})\leq\sigma^{2}\,(1+\alpha)$,
$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{qC_{\mathsf{LSI}}(\nu_{t})}{2}\,\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{q^{2}\alpha^{2}}{2\sigma^{2}}\,\frac{\mathbb{E}_{\pi_{t}}[\gamma_{\delta}\,|\cdot|^{2}]}{\mathbb{E}_{\pi_{t}}\gamma_{\delta}}\,.$
As noted earlier, $\mu_{t}\to\mu$ and $\nu_{t}\to\nu$ weakly. By [AFP00, Theorem 2.34], the lower semi-continuity of $\mathcal{R}_{q}$ implies that $\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})$. Also, $\lim_{t\downarrow 0}\mathbb{E}_{\pi_{t}}[\gamma_{\delta}\,|\cdot|^{2}]\,/\,\mathbb{E}_{\pi_{t}}\gamma_{\delta}=\mathbb{E}_{\pi}[\gamma_{\delta}\,|\cdot|^{2}]\,/\,\mathbb{E}_{\pi}\gamma_{\delta}$ due to the weak convergence of $\pi_{t}\to\pi$. Thus,
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{q^{2}\alpha^{2}}{2\sigma^{2}}\,\frac{\mathbb{E}_{\pi}[\gamma_{\delta}\,|\cdot|^{2}]}{\mathbb{E}_{\pi}\gamma_{\delta}}\,.$
We now need a general fact that for any integrable function $\rho:\mathbb{R}^{d}\to\mathbb{R}_{+}$,
$\frac{\int|\cdot|^{2}\,\gamma_{\delta}\,\mathrm{d}\rho\,}{\int\gamma_{\delta}\,\mathrm{d}\rho}\leq\frac{\int|\cdot|^{2}\,\mathrm{d}\rho}{\int\mathrm{d}\rho}=\mathbb{E}_{\rho}[|\cdot|^{2}]\,.$
Observe that $\operatorname{cov}_{\rho}(f,g)=\frac{1}{2}\,\mathbb{E}_{(X,Y)\sim\rho^{\otimes 2}}[\{f(X)-f(Y)\}\,\{g(X)-g(Y)\}]$. As $f(\cdot)=|\cdot|^{2}$ increases in ||$\cdot$|| ⋅ | while $g=\gamma_{\delta}$ decreases, $0\geq\operatorname{cov}_{\rho}(f,g)=\mathbb{E}_{\rho}[fg]-\mathbb{E}_{\rho}f\,\mathbb{E}_{\rho}g$, which proves the claim. Using this fact with $\rho=\pi$, we complete the proof of the main claim. ∎
Lemma 4.2 (Early-stopping).
Let $\pi$ be a log-concave probability measure over $\mathbb{R}^{d}$ with $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$. For $q\geq 2$, if $\sigma^{2}\gtrsim qR\,\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{1/2}d$, then $\mathcal{R}_{q}(\pi\gamma_{\sigma^{2}}\mathbin{\|}\pi)=O(1)$.
Proof.
Let $\mu\propto\pi\gamma_{\sigma^{2}}$ be the probability measure, and $\mu_{t}$ and $\pi_{t}$ the mollifications of $\mu$ and $\pi$. Recall from ($\mathsf{R\text{-}PI}$) that
$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\pi_{t})\leq\log\Bigl{(}1-\frac{qC_{\mathsf{PI}}(\pi_{t})}{4}\,\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\pi_{t})\Bigr{)}^{-1}\,.$ (4.4)
By [Mil09, §5] or [CG20, Theorem 3.13], if two log-concave probability measures $\nu_{1}$ and $\nu_{2}$ satisfy $\lVert\nu_{1}-\nu_{2}\rVert_{\mathsf{TV}}\leq\varepsilon$ for $\varepsilon\in(0,1)$, then
$C_{\mathsf{PI}}(\nu_{1})\lesssim\frac{1\vee\log\frac{1}{1-\varepsilon}}{(1-\varepsilon)^{2}}\,C_{\mathsf{PI}}(\nu_{2})\,.$
Since $\pi_{t}\to\pi$ in $\mathsf{TV}$-distance (by Scheffé), and $\pi_{t}$ is log-concave (by Prékopa–Leindler), we have $\limsup_{t\downarrow 0}C_{\mathsf{PI}}(\pi_{t})\lesssim C_{\mathsf{PI}}(\pi)$. Next, emulating the proof of Lemma 4.1 with $\alpha=\infty$, (4.3) yields
$\limsup_{t\downarrow 0}\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\pi_{t})=\frac{q}{\sigma^{4}}\,\frac{\mathbb{E}_{\pi}[\gamma_{\sigma^{2}}\,|\cdot|^{2}]}{\mathbb{E}_{\pi}\gamma_{\sigma^{2}}}\leq\frac{q\,\mathbb{E}_{\pi}[|\cdot|^{2}]}{\sigma^{4}}=\frac{qR^{2}}{\sigma^{4}}\,.$
Putting these together and using $C_{\mathsf{PI}}(\pi)\lesssim\lVert\operatorname{cov}\pi\rVert\log d$ [Kla23],
$\limsup_{t\downarrow 0}\frac{qC_{\mathsf{PI}}(\pi_{t})}{4}\,\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\pi_{t})\lesssim\frac{q^{2}R^{2}\,\lVert\operatorname{cov}\pi\rVert\log d}{\sigma^{4}}\,.$
Thus, if $\sigma^{2}\gtrsim qR\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{1/2}d$, then the RHS can be made smaller than (say) $0.1$. Putting this back to (4.4), it follows that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)=O(1)$. ∎
Covariance of strongly log-concave distributions.
[KV25a] showed that $\lVert\operatorname{cov}\pi\gamma_{\sigma^{2}}\rVert\lesssim\lVert\operatorname{cov}\pi\rVert$ if $h\gtrsim R\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{2}d\log^{2}\nicefrac{{R^{2}}}{{\Lambda}}$ for $R=\mathbb{E}_{\pi}|\cdot|$. This can be improved as an immediate corollary of the early-stopping result. For the barycenter $b_{\pi}$ of $\pi$, any unit vector $v\in\mathbb{R}^{d}$, and $\mu=\pi\gamma_{\sigma^{2}}$ with $\sigma^{2}\gtrsim R\,\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{1/2}d$,
$\Bigl{(}\int\bigl{(}v^{\mathsf{T}}(x-b_{\pi})\bigr{)}^{2}\,\mathrm{d}\mu(x)\Bigr{)}^{2}\leq\Bigl{\|}\frac{\mathrm{d}\mu}{\mathrm{d}\pi}\Bigr{\|}_{L^{2}(\pi)}^{2}\int\bigl{(}v^{\mathsf{T}}(x-b_{\pi})\bigr{)}^{4}\,\mathrm{d}\pi(x)\lesssim\Bigl{\|}\frac{\mathrm{d}\mu}{\mathrm{d}\pi}\Bigr{\|}_{L^{2}(\pi)}^{2}\lVert\operatorname{cov}\pi\rVert^{2}\,,$
where we used the reverse Hölder in the last inequality. Thus, $\lVert\operatorname{cov}\mu\rVert\lesssim\lVert\frac{\mathrm{d}\mu}{\mathrm{d}\pi}\rVert_{L^{2}(\pi)}\lVert\operatorname{cov}\pi\rVert\,.$ Since $\mathcal{R}_{2}(\mu\mathbin{\|}\pi)=O(1)$, we have
$\lVert\operatorname{cov}\mu\rVert\lesssim\lVert\operatorname{cov}\pi\rVert\,.$
Query complexity of Rényi-warmness generation.
Now that all pieces are ready, we show that an $O(1)$-warm start in $\mathcal{R}_{q}$ can be generated using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$ membership queries for $q\geq 2$, where $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$.
Proof of Theorem 1.5.
The initialization by rejection sampling requires $O(1)$ queries in expectation due to
$\frac{\int_{\mathbb{R}^{d}}e^{-d\,|x|^{2}/2}\,\mathrm{d}x}{\int_{\mathcal{K}}e^{-d\,|x|^{2}/2}\,\mathrm{d}x}=\frac{1}{\gamma_{d^{-1}}(\mathcal{K})}\leq\frac{1}{\gamma_{d^{-1}}(B_{1})}\lesssim 1\,.$
In the annealing phase, suppose that we are given $\bar{\mu}_{i}$ such that $\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$. By a weak triangle inequality for Rényi divergence (with $\lambda=2/3$),
$\mathcal{R}_{2}(\bar{\mu}_{i}\mathbin{\|}\mu_{i+1})\leq\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})+\mathcal{R}_{4}(\mu_{i}\mathbin{\|}\mu_{i+1})\leq O(1)+2=O(1)\,,$
where the last inequality follows from Lemma 4.1. By Theorem 1.4, from an $O(1)$-warmness in $\mathcal{R}_{2}$-divergence, $\mathsf{PS}_{\textup{Gauss}}$ needs $O(d^{2}\sigma_{i}^{2})$ queries to sample $X_{i+1}^{*}\sim\bar{\mu}_{i+1}$ such that $\mathcal{R}_{3}(\bar{\mu}_{i+1}\mathbin{\|}\mu_{i+1})=O(1)$. Also, doubling of initial $\sigma_{i}^{2}$ requires $O(R/\sigma_{i})$ phases, so each doubling needs $O(d^{2}\sigma_{i}R)$ queries in total. Since there are at most $O(\log qdR\Lambda^{1/2})$ doublings, and $\sigma_{i}^{2}\leq\sigma_{\text{last}}^{2}:=O(qR\Lambda^{1/2}\log^{1/2}d)$, we can obtain a sample $X^{*}$ such that $\mathcal{R}_{3}(\bar{\mu}_{\text{last}}\mathbin{\|}\mu_{\text{last}})=O(1)$, using
$O\bigl{(}q^{1/2}d^{2}R^{3/2}\Lambda^{1/4}\log^{1/4}d\times\log qdR\Lambda^{1/2}\bigr{)}=\widetilde{O}(q^{1/2}d^{2}R^{3/2}\Lambda^{1/4})\quad\text{queries}\,.$
By hypercontractivity of $\mathsf{PS}$ (Corollary 1.3), iterating $\mathsf{PS}_{\textup{Gauss}}$ started from $X^{*}$ towards $\mu_{\text{last}}$ $O(d^{2}\sigma_{\text{last}}^{2}\log_{2}q)$ times leads to a new sample $X^{\prime}$ such that $\mathcal{R}_{2q}(\operatorname{law}X^{\prime}\mathbin{\|}\mu_{\text{last}})=O(1)$. Hence, the entire annealing has used
$\widetilde{O}\bigl{(}q^{1/2}d^{2}R^{3/2}\Lambda^{1/4}+qd^{2}R\Lambda^{1/2}\bigr{)}\leq\widetilde{O}\bigl{(}qd^{2}R^{3/2}\Lambda^{1/4}\bigr{)}\quad\text{queries (due to }R^{2}\geq\Lambda\text{)}\,.$
Lastly, we conclude the proof by noting the early-stopping result (Lemma 4.2) and the weak triangle inequality, which results in $\mathcal{R}_{q}(\operatorname{law}X^{\prime}\mathbin{\|}\pi)=O(1)$. ∎
5 Log-concave sampling under an evaluation oracle
To extend our approach to log-concave distributions, we generalize what we did for uniform sampling, streamlining algorithms in [KV25a] with their samplers replaced by our new samplers.
We begin with a cleaner analysis for log-concave sampling from a warm start in §5.1 and then move onto an order-preserving annealing algorithm for warm-start generation in §5.2. To this end, we improve the complexity of a sampler for these annealing distributions.
5.1 Zeroth-order log-concave sampling with balanced guarantees
We take the exponential-lifting approach ($\mathsf{exp\text{-}lifting}$) studied in [KV25b], improving their sampler and simplifying analysis similarly to the uniform sampling case.
A mixing analysis.
Assume that $\pi_{0}^{X}$ is $M_{q}$-warm with respect to $\pi^{X}$. Note that one can generate a sample $(X^{\prime},T^{\prime})$ whose joint law is also $M_{q}$-warm with respect to $\pi^{X,T}$ by first drawing $X^{\prime}\sim\pi_{0}^{X}$ and then sampling $T^{\prime}\sim\pi^{T|X=X^{\prime}}\propto e^{-dt}|_{[V(X^{\prime})/d,\infty)}$.
Since the convergence rate of $\mathsf{PS}$ was already analyzed in §2.1, we can just reuse Proposition 2.4. The only adjustment is that we now require a bound on $C_{\mathsf{PI}}(\pi^{X,T})$, rather than $C_{\mathsf{PI}}(\pi^{X})$, as we run $\mathsf{PS}$ to sample from $\pi^{X,T}$. Since $C_{\mathsf{PI}}(\pi^{X,T})\lesssim\lVert\operatorname{cov}\pi^{X,T}\rVert\log d$ due to log-concavity of $\pi^{X,T}$, and in [KV25b, Lemma 2.5], $\lVert\operatorname{cov}\pi^{X,T}\rVert$ is bounded as
$\lVert\operatorname{cov}\pi^{X,T}\rVert\leq 2\,(\lVert\operatorname{cov}\pi^{X}\rVert+160)\lesssim\lVert\operatorname{cov}\pi^{X}\rVert\vee 1\,,$
we have $C_{\mathsf{PI}}(\pi^{X,T})\lesssim(1\vee\lVert\operatorname{cov}\pi^{X}\rVert)\log d$. Hence, by Lemma 2.5, $\mathsf{PS}_{\textup{exp}}$ with step size $h$ can ensure $(\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq)\mathcal{R}_{q}(\pi_{N}^{X,T}\mathbin{\|}\pi^{X,T})\leq\varepsilon$ after
$N\lesssim q\,\bigl{(}1\vee h^{-1}\,(1\vee\lVert\operatorname{cov}\pi^{X}\rVert)\log d\bigr{)}\log\frac{M_{q}}{\varepsilon}\quad\text{iterations}\,.$
Per-step query complexity.
Denoting by $\ell(y)=\mathbb{P}_{\mathcal{N}(y-h\alpha,hI_{d+1})}(\cdot\in\mathcal{K})$ the success probability of rejection sampling, one can check (or see [KV25b, Lemma 2.11]) that for $z=(x,t)\in\mathbb{R}^{d}\times\mathbb{R}$,
$(\pi^{Z}*\gamma_{h})(y)=\pi^{Y}(y)=\frac{\ell(y)\exp(-\alpha^{\mathsf{T}}y+\frac{1}{2}\,h\,|\alpha|^{2})}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}z)\,\mathrm{d}z}\,.$
This has a concentration phenomenon: for $\widetilde{\mathcal{K}}:=\mathcal{K}_{\delta}+h\alpha$, if $\delta\geq 13hd$,
$\pi^{Y}(\widetilde{\mathcal{K}}^{c})\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+13\delta d\bigr{)}\,.$
Hereafter, we take $h=(2\cdot 13^{2}cd^{2})^{-1}$ and $\delta=(13d)^{-1}$ for some $c>0$, under which $\delta\geq 13hd$ is equivalent to $c\geq 1/2$. Recall the following expansion lemma from [KV25b, Lemma 2.13]:
$\frac{\int_{\mathcal{K}_{s}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\leq e^{13sd}\quad\text{for }s>0\,.$
This leads to the following bound: using change of variables via $y=v+h\alpha$ in $(i)$,
$\frac{\int_{\widetilde{\mathcal{K}}}\exp(-\alpha^{\mathsf{T}}y+\frac{1}{2}\,h\,|\alpha|^{2})\,\mathrm{d}y}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\underset{(i)}{=}\frac{\int_{\mathcal{K}_{\delta}}\exp(-\alpha^{\mathsf{T}}v-\frac{1}{2}\,h\,|\alpha|^{2})\,\mathrm{d}v}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\leq\frac{\int_{\mathcal{K}_{\delta}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\leq e^{13\delta d}=e\,.$ (5.1)
As in the analysis for $\mathsf{PS}_{\textup{unif}}$, we show that the failure probability of a stage in the $k$-th iteration is at most $1/2$ under suitable choices of $h$ and $\tau$. Let $(p,q)$ be a pair of conjugates with $p,q\geq 2$. Then, $\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq M_{q}\,\{\mathbb{E}_{\pi^{Y}}[(1-\ell)^{\tau p}]\}^{1/p}$ and
$\int_{\mathbb{R}^{d+1}}(1-\ell)^{\tau p}\,\mathrm{d}\pi^{Y}=\int_{\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}}}\cdot+\int_{\widetilde{\mathcal{K}}\cap[\ell\tau\geq\log(3M_{q})]}\cdot+\int_{\widetilde{\mathcal{K}}\cap[\ell\tau<\log(3M_{q})]}\cdot=:\mathsf{A}+\mathsf{B}+\mathsf{C}\,.$
Under $c\asymp\log(3M_{q})^{p}$ and $\tau\asymp M_{q}^{p}\log(3M_{q})$, the last term can be bounded by $(\nicefrac{{1}}{{2M_{q}}})^{p}$ due to
$\mathsf{A}$ $\leq\pi^{Y}(\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}})\leq e^{-c+1}\,,$
$\mathsf{B}$ $\leq\int_{\widetilde{\mathcal{K}}\cap[\ell\tau\geq\log(3M_{q})]}\exp(-\ell\tau p)\,\mathrm{d}\pi^{Y}\leq\bigl{(}\frac{1}{3M_{q}}\bigr{)}^{p}\,,$
$\mathsf{C}$ $\leq\int_{\widetilde{\mathcal{K}}\cap[\ell\tau<\log(3M_{q})]}\frac{\ell(y)\exp(-\alpha^{\mathsf{T}}y+\frac{1}{2}\,h\,|\alpha|^{2})}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\,\mathrm{d}y\leq\frac{\log(3M_{q})}{\tau}\,\frac{\int_{\widetilde{\mathcal{K}}}\exp(-\alpha^{\mathsf{T}}y+\frac{1}{2}\,h\,|\alpha|^{2})\,\mathrm{d}y}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}$
(5.1)${\leq}\tau^{-1}e\log(3M_{q})\,.$ ) .
Hence, $\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq 1/2$, and the expected number of trials during the $k$-th iteration can be bounded by $2\tau$ just as in §2.2. Putting all these together, we can establish the query complexity of $\mathsf{PS}_{\textup{exp}}$ with balanced guarantees.
Proof of Theorem 1.6.
For $M:=\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$, we set $h^{-1}\asymp d^{2}M$ and $\tau\asymp Me^{M}$. As discussed earlier, $\mathsf{PS}_{\textup{exp}}$ suffices to iterate at most
$N\lesssim qMd^{2}(\Lambda\vee 1)\log d\log\frac{M}{\varepsilon}\quad\text{times}\,.$
Thus, the expected number of total queries during $N$ iterations is
$2\tau N\lesssim qM^{2}e^{M}d^{2}(\Lambda\vee 1)\log d\log\frac{M}{\varepsilon}=\widetilde{O}\bigl{(}qd^{2}\,(\Lambda\vee 1)e^{\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})}\log\frac{1}{\varepsilon}\bigr{)}\,,$
which completes the proof. ∎
5.2 Rényi-warmness generation for log-concave distributions
[…]
Warm-start generation.
For $l=\log 4e$, one may assume that $x_{0}=0$ (by translation) and consider
$\bar{\mathcal{K}}:=\mathcal{K}\cap\{B_{Rl}(0)\times[-21,13l-6]\}\,,$ (5.2)
which satisfies $\pi(\mathbb{R}^{d+1}\backslash\bar{\mathcal{K}})\leq 1/2$ (see [KV25b, Convex truncation and §3.1]). Thus, we can ensure that $\bar{\pi}=\pi|_{\bar{\mathcal{K}}}$ is $O(1)$-close to $\pi$ in $\mathcal{R}_{\infty}$. Below, we let $D:=Rl$ and $\mu_{i}:=\mu_{\sigma_{i}^{2},\rho_{i}}$, and denote by $\bar{\mu}_{i}$ the law of an actual sample such that $\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$.
  1. 1.
    Sample from $\mu_{0}\propto e^{-d\,|x|^{2}/2}|_{\bar{\mathcal{K}}}$ by rejection sampling with proposal $\gamma_{d^{-1}}\otimes\text{Unif}\,([-21,13l-6])$).
  2. 2.
    [Phase I: $\sigma^{2}$-warming] While $d^{-1}\leq\sigma_{i}^{2}\leq 1$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl{(}1+\frac{1}{d^{1/2}}\bigr{)}\,.$
    At the end of Phase I, run $\mathsf{PS}_{\textup{ann}}$ toward $\mu_{i+1}\propto e^{-|x|^{2}/2-t}|_{\bar{\mathcal{K}}}$, obtaining $\bar{\mu}_{i+1}$.
  3. 3.
    [Phase II: $\rho$-annealing] While $1\leq\rho_{i}\leq d$ (and $\sigma_{i}^{2}\approx 1$), move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl{(}1+\frac{1}{d^{1/2}}\bigr{)}^{-1}\quad\&\quad\rho_{i+1}=\min\bigl{(}d,\rho_{i}\,\bigl{(}1+\frac{1}{d^{1/2}}\bigr{)}\bigr{)}\,.$
    This is followed by inner annealing: until $\sigma_{i+1}^{2}\leq 1$, run $\mathsf{PS}_{\textup{ann}}$ toward new $\mu_{i+1}$ defined by
    $\sigma_{i+1}^{2}\leftarrow\sigma_{i+1}^{2}\,\bigl{(}1+\frac{\sigma_{i+1}}{D}\bigr{)}\,,$
  4. 4.
    [Phase III: $\sigma^{2}$-annealing] While $\sigma_{i}^{2}\lesssim qD\,(\Lambda^{1/2}\vee 1)\log^{1/2}d$ (and $\rho_{i}=d$), move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl{(}1+\frac{\sigma_{i}}{D}\bigr{)}\,.$
5.2.1 Proximal sampler for tilted Gaussian
We now improve a sampler for the annealing distributions $\mu_{i}$ proposed in [KV25b]. For $v:=(x,t)\in\mathbb{R}^{d}\times\mathbb{R}$ and $w:=(y,s)\in\mathbb{R}^{d}\times\mathbb{R}$, consider the augmented target given by
$\mu^{V,W}(v,w)\propto\exp\bigl{(}-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t-\frac{1}{2h}\,|w-v|^{2}\bigr{)}\big{|}_{\bar{\mathcal{K}}}\,,$
and $\mathsf{PS}_{\textup{ann}}$ with step size $h$ and threshold $\tau$ alternates the following: for $r:=\frac{\sigma^{2}}{h+\sigma^{2}}<1$, $y_{r}:=ry$, and $h_{r}:=rh$,
  • [Forward] $w_{k+1}\sim\mu^{W|V=v_{k}}=\mathcal{N}(v_{k},hI_{d+1})$.
  • [Backward] $v_{k+1}\sim\mu^{V|W=w_{k+1}}\propto\bigl{[}\mathcal{N}(ry_{k+1},h_{r}I_{d})\otimes\mathcal{N}(s_{k+1}-\rho h,h)\bigr{]}\big{|}_{\bar{\mathcal{K}}}$. To this end, draw $v_{k+1}\sim\mathcal{N}(ry_{k+1},h_{r}I_{d})\otimes\mathcal{N}(s_{k+1}-\rho h,h)$ until $v_{k+1}\in\bar{\mathcal{K}}$. If this rejection loop exceeds $\tau$ trials, discard $w_{k+1}$ and return to the forward step using the same $V=v_{k}$.
We establish its query complexity as follows:
Lemma 5.1.
Consider the annealing distribution $\mu=\mu_{\sigma^{2},\rho}$ above, given access to an evaluation oracle for $V$. For $q_{0}>1$, let $\mu_{k}$ be the law of the $k$-th iterate of $\mathsf{PS}_{\textup{ann}}$, initialized from $\mu_{0}$ toward the target $\mu$, with step size $h\asymp(d^{2}\,\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu))^{-1}$ and threshold $\tau\asymp\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}$. Then, for any $\varepsilon>0$ and $q>1$, we can ensure $\mathcal{R}_{q}(\mu_{N}\mathbin{\|}\mu)\leq\varepsilon$ after $N\asymp d^{2}(\sigma^{2}\vee 1)\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}{(2\wedge q_{0}-1)\,\varepsilon}$ iterations, using
$2\tau N=\widetilde{O}\bigl{(}d^{2}(\sigma^{2}\vee 1)\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}\log\frac{q}{(2\wedge q_{0}-1)\,\varepsilon}\bigr{)}\qquad\text{evaluation queries in expectation}\,.$
Mixing analysis.
Using the Bakry–Émery criterion and bounded perturbation, one can show that $C_{\mathsf{LSI}}(\mu_{i})\lesssim\sigma_{i}^{2}\vee 1$ (see [KV25b, Lemma 3.3]). Thus, when $\mathsf{PS}_{\textup{ann}}$ starts from $\mu_{0}^{V}$, by Corollary 1.3, it needs at most $N\lesssim h^{-1}\,(\sigma^{2}\vee 1)\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{V}\mathbin{\|}\mu^{V})}{(2\wedge q_{0}-1)\,\varepsilon}$ iterations to obtain a sample $X_{N}$ such that $\mathcal{R}_{q}(\operatorname{law}X_{N}\mathbin{\|}\mu^{V})\leq\varepsilon$.
Per-step query complexity.
Denoting by $\ell(w)$ the success probability of each trial at $w=(y,s)$, we recall the density of $\mu^{W}=\mu^{V}*\gamma_{h}$ from [KV25b, Lemma 3.7]:
$\mu^{W}(y,s)=\frac{r^{d/2}\ell(y,s)}{\int_{\bar{\mathcal{K}}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)\,\mathrm{d}x\mathrm{d}t}\,\exp\bigl{(}-\frac{1}{2\tau\sigma^{2}}\,|y_{r}|^{2}\bigr{)}\exp\bigl{(}-\rho s+\frac{1}{2}\,\rho^{2}h\bigr{)}\,.$
Moreover, $\mu^{W}$ has a concentration property (see [KV25b, Lemma 3.8]) that if $\delta\geq 24hd$,
$\mu^{W}(\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}})\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+24\delta d+hd^{2}\bigr{)}\quad\text{for}\quad\widetilde{\mathcal{K}}=\mathopen{}\mathclose{{}\left[\begin{array}[]{cc}r^{-1}I_{d}\\ &1\end{array}}\right]\bar{\mathcal{K}}_{\delta}+\mathopen{}\mathclose{{}\left[\begin{array}[]{c}0\\ \rho h\end{array}}\right]\,.$
Choosing $h=(24^{2}cd^{2})^{-1}$ and $\delta=(1+\nicefrac{{1}}{{24c}})\,d^{-1}$ for some $c\geq 1$, we can ensure that $\delta\geq 24hd$ and $\mu^{W}(\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}})\leq e^{-c/2+1}$. There is also an expansion lemma (see [KV25b, Lemma 3.9]): for $\rho\in(0,d]$,
$r^{-d/2}\int_{\bar{\mathcal{K}}_{s}}e^{-\frac{1}{2r\sigma^{2}}\,|z|^{2}-\rho l}\,\mathrm{d}z\mathrm{d}l\leq 2\exp(hd^{2}+24sd)\int_{\bar{\mathcal{K}}}e^{-\frac{1}{2\sigma^{2}}\,|z|^{2}-\rho l}\,\mathrm{d}z\mathrm{d}l\,.$
Using change of variables and this expansion lemma,
$\frac{\int_{\widetilde{\mathcal{K}}}r^{d/2}\exp(-\frac{1}{2\tau\sigma^{2}}\,|y_{r}|^{2}-\rho s+\frac{1}{2}\,\rho^{2}h)}{\int_{\bar{\mathcal{K}}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)}=\frac{r^{-d/2}\int_{\bar{\mathcal{K}}_{\delta}}\exp(-\frac{1}{2r\sigma^{2}}\,|y|^{2}-\rho s-\frac{1}{2}\,\rho^{2}h)}{\int_{\bar{\mathcal{K}}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)}\leq e^{27}$ (5.3)
Similarly for earlier analyses, the failure probability of one stage in the $k$-th iteration can be bounded by $1/2$ under suitable choices of $h$ and $\tau$. For a conjugate pair $(p,q)$ with $p,q>1$, we have $\mathbb{E}_{\mu_{k}^{W}}[(1-\ell)^{\tau}]\leq M_{q}\,\{\mathbb{E}_{\mu_{k}^{W}}[(1-\ell)^{\tau p}]\}^{1/p}$ and
$\mathbb{E}_{\mu_{k}^{W}}[(1-\ell)^{\tau p}]=\int_{\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}}}\cdot+\int_{\widetilde{\mathcal{K}}\cap[\ell\tau\geq\log(3M_{q})]}\cdot+\int_{\widetilde{\mathcal{K}}\cap[\ell\tau<\log(3M_{q})]}=:\mathsf{A}+\mathsf{B}+\mathsf{C}\,.$
Under $c\asymp\log(3M_{q})^{p}$ and $\tau\asymp M_{q}^{p}\log(3M_{q})$, the RHS can be bounded by $(1/2M_{q})^{p}$ as seen in
$\mathsf{A}$ $\leq\mu^{W}(\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}})\leq e^{-c/2+1}\,,$
$\mathsf{B}$ $\leq\int_{\widetilde{\mathcal{K}}\cap[\ell\tau\geq\log(3M_{q})]}e^{-\ell\tau p}\,\mathrm{d}\mu^{W}\leq\bigl{(}\frac{1}{3M_{q}}\bigr{)}^{p}\,,$
$\mathsf{C}$ $\leq\int_{\widetilde{\mathcal{K}}\cap[\ell\tau<\log(3M_{q})]}\frac{r^{d/2}\ell(y,s)\exp(-\frac{1}{2r\sigma^{2}}\,|y_{r}|^{2})\exp(-\rho s+\frac{1}{2}\,\rho^{2}h)}{\int_{\bar{\mathcal{K}}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)}$
[…]
Thus, $\mathbb{E}_{\mu_{k}^{W}}[(1-\ell)^{\tau}]\leq 1/2$, and per-iteration complexity is bounded by $2\tau$. Combining the mixing analysis and per-iteration analysis proves the query complexity of $\mathsf{PS}_{\textup{ann}}$ (Lemma 5.1).
5.2.2 Query complexity of Rényi-warmness generation
We first recall a result on closeness of two log-concave distributions from [KV25a, Lemma 4.4].
Lemma 5.2.
Let $\mathrm{d}\nu\propto e^{-V}\,\mathrm{d}x$ and $\mathrm{d}\mu\propto e^{-(1+\alpha)\,V}\,\mathrm{d}x$ be log-concave probability measures over $\mathbb{R}^{d}$. For $q>1$ and $\delta>0$ such that $01-q\delta>$ 0 ,
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\begin{cases}\frac{qd\alpha^{2}}{2}&\text{if }\alpha\geq 0\,,\\ \frac{qd\alpha^{2}}{1-q\delta}&\text{if }\alpha\in[-\delta/2,0]\,.\end{cases}$
We now bound the complexity of the overall algorithm as $\widetilde{O}(d^{2.5}+qd^{2}(R^{3/2}\vee 1)(\Lambda^{1/4}\vee 1))$.
Proof of Theorem 1.6.
At initialization, $\mu_{0}$ can be sampled by rejection sampling with $O(1)$ queries in expectation (see [KV25b, Lemma 5.6]). In Phase I, any consecutive distributions are $O(1)$-close in $\mathcal{R}_{2}$ by Lemma 5.2. Since there are $O(d^{1/2}\log d)$ many phases, with each one requiring $\widetilde{O}(d^{2})$ queries, the query complexity of Phase I is $\widetilde{O}(d^{2.5})$. Meanwhile, we relay $\mathcal{R}_{2}$-warmness through the weak triangle inequality (as done in the proof of Theorem 1.5).
In Phase II, closeness during the simultaneous annealing (called the outer annealing hereafter) is justified again by Lemma 5.2, and $\mathsf{PS}_{\textup{ann}}$ requires $\widetilde{O}(d^{2})$ queries for each outer annealing. Each outer annealing is followed by $O(D/d^{1/2})$ many inner phases, where $\sigma^{2}$ goes back to $1$ again. Closeness within the inner annealing follows from Lemma 4.1, and $\mathsf{PS}_{\textup{ann}}$ requires $\widetilde{O}(d^{2})$ queries. Thus, completion of one outer annealing (combined with the inner annealing) requires $\widetilde{O}(d^{2}+d^{1.5}D)$ queries. As there are $O(d^{1/2}\log d)$ outer phases, the total complexity of Phase II is $\widetilde{O}(d^{2.5}+d^{2}D)$.
In Phase III, closeness of annealing distributions follow again from Lemma 4.1. Since doubling of initial $\sigma^{2}$ requires $O(D/\sigma)$ phases, with phase requiring $O(d^{2}\sigma^{2})$ queries from an $O(1)$-warm start, each doubling uses $\widetilde{O}(d^{2}\sigma D)$ queries. As $\sigma^{2}$ is increased until $\sigma^{2}\lesssim qD\,(\Lambda^{1/2}\vee 1)\log^{1/2}d$, it takes Phase III $\widetilde{O}(q^{1/2}d^{2}D^{3/2}(\Lambda^{1/4}\vee 1))$ queries to obtain a sample whose law is $O(1)$-close to $\mu_{\text{last}}$ in $\mathcal{R}_{2}$-divergence. Due to hypercontractivity of $\mathsf{PS}$ again, iterating $\mathsf{PS}_{\textup{ann}}$ toward $\mu_{\text{last}}$ $O(d^{2}\sigma_{\text{last}}^{2}\log_{2}q)$ times returns a sample $Z^{*}$ such that $\mathcal{R}_{2q}(\operatorname{law}Z^{*}\mathbin{\|}\mu_{\text{last}})=O(1)$. Hence, Phase III uses $\widetilde{O}(qd^{2}D^{3/2}\,(\Lambda^{1/4}\vee 1))$ queries in total. Also, by the early-stopping result and weak triangle inequality, we have $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\bar{\pi})=O(1)$. Since $\mathcal{R}_{\infty}(\bar{\pi}\mathbin{\|}\pi)=O(1)$, we obtain that $\mathcal{R}_{q}(\operatorname{law}Z^{*}\mathbin{\|}\pi)=O(1)$.
As $D\asymp R$, adding up complexities of each phase proves the claim. ∎
6 Discussion
We conclude this work by highlighting interesting directions for future research.
Further simplification of general log-concave sampling.
We now have the clean end-to-end routine for warm-start generation and uniform sampling. However, we adopted a more convoluted approach for generating a warm start for general log-concave distributions, since our current approach to log-concave sampling incurs the additive $O(qd^{2}\log\nicefrac{{1}}{{\varepsilon}})$ queries in the complexity. Nonetheless, we believe that general log-concave sampling can be further streamlined, and an intriguing open question is whether this additive factor can be removed. Resolution of this question would lead to a simpler algorithm and analysis.
Irreversible algorithms for uniform sampling.
Unlike the well-conditioned setting with access to a first-order oracle, the complexity of $\mathcal{R}_{2}$-warmness generation (i.e., $d^{2}R^{3/2}\Lambda^{1/4})$ is worse that that of sampling from a warm start (i.e., $d^{2}\Lambda$). A notable key ingredient in [AC24] for achieving the same complexity for warm-start generation as $\mathsf{MALA}$ from a warm start is the use of the Underdamped Langevin Algorithm, an irreversible sampler. In contrast, in our annealing phase, the proximal sampler $\mathsf{PS}_{\textup{Gauss}}$ for truncated Gaussians is a reversible sampler. Developing an irreversible sampler for truncated Gaussians in our general setting could be a promising direction for tightening the complexities of both annealing and sampling.
Acknowledgement.
This work was done while the author was a student researcher at Google Research. The author thanks Sinho Chewi for valuable feedback.