### v2

# Zeroth-order Logconcave Sampling

Abstract
We study the zeroth-order query complexity of sampling from a general logconcave distribution: given access to an evaluation oracle for a convex function $V:\mathbb{R}^{d}\rightarrow\mathbb{R}\cup\{\infty\}$, output a point from a distribution within $\varepsilon$-distance to the density proportional to $e^{-V}$. A long line of work provides efficient algorithms for this problem in TV distance, assuming a pointwise warm start (i.e., in $\infty$-Rényi divergence), and using annealing to generate such a warm start. Here, we address the natural and more general problem of using a $q$-Rényi divergence warm start to generate a sample that is $\varepsilon$-close in $q$-Rényi divergence. Our first main result is an algorithm with this end-to-end guarantee with state-of-the-art complexity for $q=\widetilde{\Omega}(1)$. Our second result shows how to generate a $q$-Rényi divergence warm start directly via annealing, by maintaining $q$-Rényi divergence throughout, thereby obtaining a streamlined analysis and improved complexity. Such results were previously known only under the stronger assumptions of smoothness and access to first-order oracles. We also show a lower bound for Gaussian annealing by disproving a geometric conjecture about quadratic tilts of isotropic logconcave distributions.
Central to our approach, we establish hypercontractivity of the heat adjoint and translate this into improved mixing time guarantees for the Proximal Sampler. The resulting analysis of both sampling and annealing follows a simplified and natural path, directly tying convergence rates to isoperimetric constants of the target distribution.
1 Introduction
We study the zeroth-order query complexity of sampling from an arbitrary logconcave distribution $\pi$ with an arbitrary output guarantee, assuming access to an evaluation oracle for $\log\pi$ up to a fixed unknown additive constant (equivalently, an evaluation oracle for a convex function $V$, with the goal of sampling from the distribution with density proportional to $e^{-V}$). This is a general setting for logconcave sampling—for example, it includes uniform sampling from convex bodies, which will serve as a running example throughout this paper. Results in this general setting provide a theoretical baseline, highlighting the benefits of additional assumptions often made in other sampling work, such as log-smoothness (i.e., $\nabla\log\pi$ is Lipschitz) or access to first-order oracles (i.e., $\nabla\log\pi$). Indeed, much of our work is motivated by gaps between this general case and more favorable settings where first-order oracles are available for unconstrained and log-smooth distributions. In practical terms, zeroth-order samplers are natural default algorithms, since evaluation oracles typically require only function value computations, making them the most general and easiest to implement.
We now make the problem setup more explicit, focusing first on uniform sampling. Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body (i.e., a convex and compact set with nonempty interior) containing the $d$-dimensional unit ball $B_{1}$. Given access to the membership oracle for $\mathcal{K}$ (which answers whether a queried point belongs to $\mathcal{K}$) [GLS93], we ask: how can we sample uniformly from $\mathcal{K}$, and how many oracle queries are required? More formally, given a target accuracy $\varepsilon>0$ and a probability divergence (or distance) $\mathsf{D}$—such as the total variation ($\mathsf{TV}$) distance or $\chi^{2}$-divergence—what is the query complexity of generating a sample whose law is $\varepsilon$-close to the uniform distribution $\pi$ over $\mathcal{K}$ in $\mathsf{D}$?
In this work, we propose a uniform sampler that allows us to establish the first balanced end-to-end complexity guarantee, meaning that the Rényi divergence orders of the initial distribution and the output distribution are the same. This result parallels guarantees found in sampling research in the log-smooth setting. Next, we provide a simple annealing scheme that, for the first time, generates a warm start in $q$-Rényi divergence. To this end, we prove hypercontractivity under simultaneous heat flow, under a logarithmic Sobolev inequality. Finally, we extend these results to arbitrary logconcave distributions under evaluation oracles. Alongside, we present a streamlined analysis leveraging tools from diverse mathematical fields. Combining these contributions, we establish the improved zeroth-order query complexity of sampling from an arbitrary logconcave distribution with $\mathcal{R}_{q}$-guarantees.
1.1 Background and results
In this section, $\pi_{\mathcal{K}}$ denotes the uniform distribution over a convex body $\mathcal{K}$ with diameter $D$, and $\pi$ refers to a logconcave distribution over $\mathbb{R}^{d}$. For clarity, we recall the $q$-Rényi divergence, defined as $\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{q}{q-1}\log\,\lVert\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}\rVert_{L^{q}(\nu)}$ for $q>1$, and its monotonicity in $q$, and $2\,\lVert\cdot\rVert_{\mathsf{TV}}^{2}\leq\mathsf{KL}=\lim_{q\downarrow 1}\mathcal{R}_{q}\leq\mathcal{R}_{2}=\log(1+\chi^{2})\leq\chi^{2}\leq\mathcal{R}_{\infty}=\log\operatorname{ess\,sup}_{\nu}\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}$. See §1.2 for more details. When a sampler is initialized with a distribution $\nu_{0}$ to sample from a target distribution $\nu$, we define the warmness parameter as $M_{q}=\lVert\mathrm{d}\nu_{0}/\mathrm{d}\nu\rVert_{L^{q}(\nu)}$.
1.1.1 Uniform sampling from a warm start
Classical algorithms: Ball walk and Hit-and-run.
The $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$ are well-studied random walks (i.e., Markov chains) for uniform sampling. These algorithms are analyzed through the conductance framework, where a lower bound on conductance yields a mixing-time guarantee. Establishing such a lower bound involves two ingredients: one-step coupling (i.e., the closeness of transition kernels at two nearby points) and a Cheeger isoperimetric inequality for the target distribution. The Cheeger constant of a probability measure $\pi$ over $\mathbb{R}^{d}$ is defined as
$C_{\mathsf{Ch}}(\pi):=\inf_{S:\pi(S)\leq\frac{1}{2}}\frac{\pi(\partial S)}{\pi(S)}\,$,
where $S\subset\mathbb{R}^{d}$ is any open set with smooth boundary, $S_{\eta}=\{x:d(x,S)\leq\eta\}$, and $\pi(\partial S):=\liminf_{\eta\downarrow 0}\frac{\pi(S_{\eta})-\pi(S)}{\eta}$.
The $\mathsf{Ball\ walk}$ was introduced by Lovász [LOV90]; in each step, it samples a random point uniformly from a ball of fixed radius around the current point, accepting it if the point lies within $\mathcal{K}$. Later, Lovász and Simonovits established that the $\mathsf{Ball\ walk}$ satisfies $\mathcal{R}_{\infty}\to\mathsf{TV}$ with complexity $\operatorname{poly}(d,D,M_{\infty}/\varepsilon)$ [LS93, Remark on p. 398], meaning the initial warmness should be given in $\mathcal{R}_{\infty}$-divergence and the final output is $\varepsilon$-close to $\pi_{\mathcal{K}}$ in $\mathsf{TV}$-distance. They also analyzed an indirect approach that uses the $\mathsf{Ball\ walk}$ (with a Metropolis filter) to sample from a logconcave distribution and then post-processes the sample to bring it closer to $\pi_{\mathcal{K}}$. The resulting algorithm, also of type $\mathcal{R}_{\infty}\to\mathsf{TV}$, was shown to have query complexity $\operatorname{poly}(d,D,\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$.
This approach was further improved by Kannan, Lovász, and Simonovits [KLS97], who introduced the $\mathsf{Speedy\ walk}:\mathcal{R}_{\infty}\to\mathsf{TV}$. Similar to [LS93], it begins by sampling from a biased logconcave variant $\pi^{\prime}$ (of $\pi_{\mathcal{K}}$), which is invariant under the following per-step transition: the next point $x_{i+1}$ is sampled uniformly from $\mathcal{K}\cap B_{\delta}(x_{i})$ via rejection sampling, using the uniform distribution over $B_{\delta}(x_{i})$ as the proposal. Then, a post-processing step is used to reduce the bias. They showed that the mixing time of this chain is $O(d^{2}C_{\mathsf{Ch}}^{-2}(\pi^{\prime})\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$ and that each rejection sampling step requires $O(M_{\infty})$ queries, leading to an overall complexity of $\widetilde{O}(M_{\infty}d^{2}D^{2}\log\nicefrac{{1}}{{\varepsilon}})$ (due to $C_{\mathsf{Ch}}(\pi)\gtrsim D^{-1}$ [LS93]). Hereafter, by following convention, the $\mathsf{Ball\ walk}$ refers to the $\mathsf{Speedy\ walk}$ with the post-processing step. Later, progress on the Kannan-Lovász-Simonovits (KLS) conjecture [KLS95] improved this bound to $C_{\mathsf{Ch}}^{2}(\pi)\gtrsim(\lVert\operatorname{cov}\pi\rVert\log d)^{-1}$ [KLA23]. As a result, the complexity of the $\mathsf{Ball\ walk}$ is now $\widetilde{O}(M_{\infty}d^{2}\lVert\operatorname{cov}\pi\rVert\log\nicefrac{{1}}{{\varepsilon}})$.
The $\mathsf{Hit\text{-}and\text{-}Run}$ algorithm [SMI84] was first analyzed by Lovász [LOV99] via $s$-conductance [LS90], with complexity $\widetilde{O}(M_{\infty}d^{2}D^{2}/\varepsilon^{2})$. Lovász and Vempala [LV06b] showed that the conductance of $\mathsf{Hit\text{-}and\text{-}Run}$ is in fact $\Omega(1/dD)$ resulting in a mixing time of $\widetilde{O}(d^{2}D^{2}\log\nicefrac{{M_{2}}}{{\varepsilon}})$. Each step of $\mathsf{Hit\text{-}and\text{-}Run}$ selects a random chord passing the current point and moves to a random point on the chord. Here, uniform sampling from the chord is implemented via rejection sampling with complexity $O(\log D)$, where the proposal distribution is uniform over a slightly longer chord (found by binary search). They established a novel Cheeger isoperimetric inequality in terms of the cross-cut ratio (rather than Euclidean distance) and showed $\mathsf{Hit\text{-}and\text{-}Run}:\mathcal{R}_{2}\to\mathcal{R}_{2}$ with complexity $\widetilde{O}(d^{2}D^{2}\log\nicefrac{{M_{2}}}{{\varepsilon}})$, making it the first random walk with a balanced guarantee (and to date the only walk with logarithmic dependence on both the initial warmness and the target error). From an $O(1)$-warm start in $\mathcal{R}_{\infty}(\geq\mathcal{R}_{2})$, the $\mathsf{Ball\ walk}$ currently has better complexity than $\mathsf{Hit\text{-}and\text{-}Run}$.
Despite these and many other advances over the past four decades, a unified theory for the complexity of this fundamental problem is missing. First, the $\mathsf{Ball\ walk}$ takes an indirect approach to uniform sampling: it samples from a biased distribution to obtain a mixing guarantee, followed by a post-processing step to reduce the bias. Second, conductance-based analyses require an algorithmic modification making the chains lazy (i.e., the chain remains at its current point with some probability). It would be desirable to develop a sampling algorithm that avoids these additional processing steps. Finally, can we replace the “unbalanced” guarantee of the $\mathsf{Ball\ walk}$ (i.e., $\mathcal{R}_{\infty}\to\mathsf{TV}$) to an arguably more natural $\mathcal{R}_{q}\to\mathcal{R}_{q}$ guarantee for any desired $q$? This question becomes even more natural when considering recent developments in sampling under log-smoothness.
Sampling under smoothness.
Research in sampling has made substantial progress by analyzing discretizations of continuous-time processes. A central example is the Langevin diffusion $(X_{t})_{t\geq 0}$ in $\mathbb{R}^{d}$, described by the stochastic differential equation
$\mathrm{d}X_{t}=-\nabla V(X_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}B_{t}\,,\qquad X_{0}\sim\nu_{0}\,$,
where $V:\mathbb{R}^{d}\to\mathbb{R}$ is a potential function satisfying mild regularity conditions, and $B_{t}$ denotes Brownian motion. The law $\nu_{t}$ of $X_{t}$ converges to the stationary distribution $\nu\propto e^{-V}$ as $t\to\infty$. A spectral analysis of the infinitesimal generator for this process yields a connection between the rate of convergence and the Poincaré constant of $\nu$ (see [BGL14, §4]). A probability measure $\pi$ on $\mathbb{R}^{d}$ is said to satisfy a Poincaré inequality with constant $C$ if for any locally Lipschitz function $f\in L^{2}(\pi)$,
$\operatorname{Var}_{\pi}f:=\int\Bigl(f-\int f\,\mathrm{d}\pi\Bigr)^{2}\,\mathrm{d}\pi\leq C\int|\nabla f|^{2}\,\mathrm{d}\pi\,$, ($\mathsf{PI}$)
and the smallest such $C$ is called the Poincaré constant $C_{\mathsf{PI}}(\pi)$. The seminal work of Jordan, Kinderlehrer, and Otto [JKO98] interpreted Langevin diffusion as the $2$-Wasserstein gradient flow of $\mathsf{KL}(\cdot\mathbin{\|}\nu)$, revealing a connection to the Logarithmic Sobolev Inequality (LSI), which is stronger than the Poincaré inequality. A probability measure $\pi$ on $\mathbb{R}^{d}$ is said to satisfy a logarithmic Sobolev inequality with constant $C$ if for any locally Lipschitz function $f\in L^{2}(\pi)$,
$\operatorname{Ent}_{\pi}(f^{2}):=\int f^{2}\log f^{2}\,\mathrm{d}\pi-\int f^{2}\,\mathrm{d}\pi\cdot\log\int f^{2}\,\mathrm{d}\pi\leq 2C\int|\nabla f|^{2}\,\mathrm{d}\pi\,$, ($\mathsf{LSI}$)
and the smallest such $C$ is referred to as the log-Sobolev constant $C_{\mathsf{LSI}}(\pi)$. Under ($\mathsf{PI}$) and ($\mathsf{LSI}$) respectively, the Langevin process converges exponentially:
$\chi^{2}(\nu_{t}\mathbin{\|}\nu)\leq\exp\bigl(-\frac{2t}{C_{\mathsf{PI}}(\nu)}\bigr)\,\chi^{2}(\nu_{0}\mathbin{\|}\nu)\,,\quad\text{and}\quad\mathsf{KL}(\nu_{t}\mathbin{\|}\nu)\leq\exp\bigl(-\frac{2t}{C_{\mathsf{LSI}}(\nu)}\bigr)\,\mathsf{KL}(\nu_{0}\mathbin{\|}\nu)\,.$ (1.1)
These bounds were extended to $\mathcal{R}_{q}$-divergence in [CLL19, VW23].
With the assumption of bounded smoothness (i.e., $\beta:=\lVert\nabla^{2}V\rVert<\infty$), a short-time discretization of the Langevin process yields an algorithm that preserves the balanced guarantees of the continuous process. Specifically, it has been shown to converge in $\mathsf{KL}$ (i.e., $\mathcal{R}_{1}\to\mathcal{R}_{1}$) with complexity $\operatorname{poly}(d/\varepsilon,\beta,C_{\mathsf{LSI}}(\nu),\log\mathsf{KL}(\nu_{0}\mathbin{\|}\nu))$ [VW23]. This result was further extended to $\mathcal{R}_{q}$-divergence [CEL+24] achieving a query complexity of $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{PI}}(\nu),\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$ or $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{LSI}}(\nu),\log\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$. We note that the dependence on $\nicefrac{{1}}{{\varepsilon}}$ can be made polylogarithmic by incorporating a Metropolis filter (e.g., $\mathsf{MALA}$ [CLA+21, WSC22, CG23a]) or by considering $\mathsf{MHMC}$ [CV22, CDW+20, CG23b]). For a comprehensive review, we refer readers to the monograph [CHE25].
The complexity of the $\mathsf{Ball\ walk}$ can be described as $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{PI}}(\nu),\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$. In fact, for logconcave probability measures $\pi$, it holds that
$\frac{1}{4}\leq\frac{C_{\mathsf{Ch}}^{-2}(\pi)}{C_{\mathsf{PI}}(\pi)}\leq 9\,$,
where the first inequality is due to Cheeger [CHE70], and the second follows from the Buser–Ledoux inequality [BUS82, LED04]. Thus, $C_{\mathsf{PI}}$ could be replaced with $C_{\mathsf{Ch}}^{-2}$, although it remains unclear how to interpret the smoothness parameter $\beta$ in the context of convex-body sampling.
The proximal sampler.
Continuing this line of work based on algorithmic diffusion, the $\mathsf{Proximal\ sampler}$ ($\mathsf{PS}$) [LST21] has led to substantial progress on sampling. For a target distribution $\pi^{X}\propto e^{-V}$ over $\mathbb{R}^{d}$, $\mathsf{PS}$ considers the augmented distribution $\pi(x,y)\propto\pi^{X}(x)\,\gamma_{h}(y-x)$ over $\mathbb{R}^{d}\times\mathbb{R}^{d}$, where $\gamma_{h}$ denotes the centered Gaussian with covariance $hI_{d}$. Each iteration consists of two steps: for step size $h>0$,
$\mbox{(i) }y_{k+1}\sim\pi^{Y|X=x_{k}}=\mathcal{N}(x_{k},hI_{d}),\quad\mbox{ and }\quad\mbox{(ii) }x_{k+1}\sim\pi^{X|Y=y_{k+1}}.$
When $\pi^{X}$ is $\beta$-log-smooth, rejection sampling for the backward step $x_{i+1}\sim\pi^{X|Y=y_{i+1}}$ can be performed using $O(1)$ evaluation queries and one proximal query [^1] The proximal query for a function $V:\mathbb{R}^{d}\to\mathbb{R}$ and step size $h>0$ is the function defined as $x\in\mathbb{R}^{d}\mapsto\arg\min\{V(\cdot)+\frac{1}{2h}\,\lVert\cdot-x\rVert^{2}\}$. for any $y_{i+1}$ and $h\asymp(\beta d)^{-1}$.
Further refinements were obtained by [CCS+22], which showed that one iteration of $\mathsf{PS}$ can be interpreted as the composition of two continuous-time stochastic processes—Brownian motion and its time-reversal. Since Brownian motion naturally defines a heat semigroup $(P_{t})_{t\geq 0}$ from $L^{2}(\nu*\gamma_{t})$ to $L^{2}(\nu)$, as $P_{t}f:=f*\gamma_{t}$, and the time-reversal also admits a semigroup $(Q_{t})_{t\geq 0}$ from $L^{2}(\nu)$ to $L^{2}(\nu*\gamma_{t})$, defined as $Q_{t}f:=P_{t}(f\nu)/P_{t}\nu$ [KP23], this allowed [CCS+22] to carry out semigroup calculus, leading to contraction results similar to (1.1). In particular, this framework yields query complexity of $O(qd\beta C_{\mathsf{PI}}(\nu)\,\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu)\log\nicefrac{{1}}{{\varepsilon}})$ for achieving $\varepsilon$-closeness in $\mathcal{R}_{q}$ (i.e., $\mathsf{PS}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ under ($\mathsf{PI}$) and smoothness). See §3.1 for further details on these semigroups.
Proximal sampler for uniform sampling.
Convex body sampling does not satisfy bounded smoothness. Kook, Vempala, and Zhang [KVZ24] adapted these ideas to convex body sampling by introducing the $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ algorithm. They extended the contraction result to constrained distributions and developed a new efficient implementation for sampling from $\pi^{X|Y=y}\propto\gamma_{h}(\cdot-y)|_{\mathcal{K}}$. They introduced a threshold parameter $\tau$: if the number of attempts in the rejection sampling subroutine with proposal $\mathcal{N}(y,hI_{d})$ exceeds $\tau$, then $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ halts and declares failure. To control this failure probability $\eta\in(0,1)$, they carefully set the parameters $h,\eta,\tau$ so that with high probability, the backward step can be completed using $\widetilde{O}(M_{\infty}\operatorname{polylog}\nicefrac{{dD}}{{\eta\varepsilon}})$ queries. Consequently, they established $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ with query complexity $\widetilde{O}(qM_{\infty}d^{2}\Lambda\operatorname{polylog}\nicefrac{{D}}{{\eta\varepsilon}})$ for $\Lambda:=\lVert\operatorname{cov}\pi\rVert$, generalizing $\mathsf{Ball\ walk}$’s output guarantee. Later, Kook and Zhang [KZ25] established a balanced guarantee of $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ using ($\mathsf{LSI}$), with query complexity $\widetilde{O}(M_{\infty}d^{2}D^{2}\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$.
Nevertheless, some algorithmic tricks (e.g., the failure probability) still remain in $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$. Moreover, the required initial warmness remains the stringent $\mathcal{R}_{\infty}$. Kook and Vempala [KV25a] relaxed the warmness to $\mathcal{R}_{c}$ with $c=2+\operatorname{polylog}\nicefrac{{dD}}{{\eta\varepsilon}}=\widetilde{O}(1)$ for achieving $\varepsilon$-closeness in $\mathcal{R}_{2}$ (i.e., $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{c}\to\mathcal{R}_{2}$ under ($\mathsf{PI}$)), with query complexity $\widetilde{O}(M_{c}d^{2}\Lambda\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$ . This result raised the question and gave hope for fully balancing the warmness and output guarantee.
Proximal sampler with restart for balanced guarantees (§2).
We propose the sampler, $\mathsf{PS}_{\textup{unif}}$—essentially $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ with failure replaced by restart: starting with $x_{0}\sim\pi_{0}^{X}$, step size $h$, and threshold $\tau$,
  • [Forward] $y_{k+1}\sim\pi^{Y|X=x_{k}}=\gamma_{h}(\cdot-x_{k})$.
  • [Backward] $x_{k+1}\sim\pi^{X|Y=y_{k+1}}\propto\gamma_{h}(\cdot-y_{k+1})\,\mathds{1}_{\mathcal{K}}(\cdot)$. To this end, draw $x_{k+1}\sim\mathcal{N}(y_{k+1},hI_{d})$ repeatedly until $x_{k+1}\in\mathcal{K}$. If this rejection loop exceeds $\tau$ trials, declare failure and restart scratch from a new $x_{0}\sim\pi_{0}^{X}$.
One iteration refers to the successful completion of both steps. Note that each check $x_{k+1}\in\mathcal{K}$ incurs a single query to the membership oracle.
This algorithm always succeeds and has a complexity guarantee from an $O(1)$-warm start, where the Rényi divergence orders of initial warmness and output distribution guarantee are the same (i.e., $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ under ($\mathsf{PI}$)). This is our first main result. Hereafter, we use $B_{1}$, $B_{2}$, or $B_{2}^{d}$ to denote a $\ell_{2}$-unit ball in $\mathbb{R}^{d}$.
Theorem 1.1 (Uniform sampling from a warm start).
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ given by a membership oracle with $B_{1}\subset\mathcal{K}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$. Given $\varepsilon\in(0,1/100)$, an initial distribution $\pi_{0}$, and $q\geq 2\vee\widetilde{\Omega}(\log(d^{2}\Lambda\log\frac{1}{\varepsilon}))$ such that $M_{q}=\lVert\nicefrac{{\mathrm{d}\pi_{0}}}{{\mathrm{d}\pi}}\rVert_{L^{q}(\pi)}\leq 10$, consider the algorithm that runs $\mathsf{PS}_{\textup{unif}}$ initialized at $\pi_{0}$ and restarts if $\mathsf{PS}_{\textup{unif}}$ fails before completing $N=\widetilde{O}(qd^{2}\Lambda\log^{2}\frac{1}{\varepsilon})$ iterations. There exist choices of $h$ and $\tau$ such that this algorithm restarts with probability at most $1/50$, and its output $Z$ satisfies $\mathcal{R}_{q}(\operatorname{law}Z\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(qd^{2}\Lambda\log^{6}\frac{1}{\varepsilon})$ total membership queries in expectation.
1.1.2 Improved warm-start generation via annealing
$\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ exhibits polynomial dependence on $M_{\widetilde{O}(1)}$, and a “cold” start (e.g., the uniform distribution over a small ball) yields $M_{\widetilde{O}(1)}\lesssim\exp(d,D)$. Hence, efficient warm-start generation is necessary to reduce the complexity overhead incurred by a cold start, so warm-start generation algorithms have been developed alongside sampling algorithms.
Warm-start generation in the well-conditioned setting.
The necessity of warmness is not limited to this general setting; its importance also arises in the well-conditioned setting (i.e., $\log\pi$ is $\alpha$-strongly convex and $\beta$-smooth). In fact, $\mathsf{MALA}$ from an $O(1)$-warm start in $\mathcal{R}_{2}$ can take a step size $h\asymp(\beta d^{1/2})^{-1}$, achieving the first-order complexity of $\widetilde{O}(\kappa d^{1/2}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}}$) in the well-conditioned setting. In contrast, from a feasible start, it can only take a smaller step size $h\asymp(\beta d)^{-1}$, resulting in $\widetilde{O}(\kappa d\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}})$ complexity, where $\kappa:=\beta/\alpha$ denote the condition number of $\log\pi$. Recently, [AC24] showed that combining the Underdamped Langevin Algorithm with $\mathsf{PS}$ can generate an $O(1)$-warm start in $\mathcal{R}_{2}$-divergence using $\widetilde{O}(\kappa d^{1/2})$ first-order queries, thus incurring no overhead for $\mathsf{MALA}$ as a result.
Simulated annealing.
In our general setup, a classical strategy for warm-start generation is an annealing scheme, which involves a sequence of probability distributions $\{\mu_{i}\}_{i\leq m}$ to interpolate between an easy-to-sample initial distribution $\mu_{0}$ and the target distribution $\pi_{\mathcal{K}}$. The key idea is to ensure that each pair of consecutive distributions $\mu_{i}$ and $\mu_{i+1}$ is $O(1)$-close in some appropriate metric (e.g., $\mathcal{R}_{\infty}$-divergence), thereby enabling efficient transitions by a sampler along the sequence.
In the context of convex-body sampling, several annealing schemes have been proposed: (i) uniform distributions over truncated convex bodies intersected with balls of increasing radii [DFK91, LS90, LS93, KLS97], (ii) exponential distributions of the form $e^{-cx}$ with decreasing $c$ over a convex set, derived from the so-called “pencil” construction [LV06c], and (iii) Gaussian distributions with increasing variance, truncated to the body [CV15, CV18, KZ25, KV25a].
Among these, all approaches except [KZ25] provide a warm start $\mu$ satisfying $\lVert\mu-\pi_{\mathcal{K}}\rVert_{\mathsf{TV}}\leq\varepsilon$, while [KZ25] guarantees $\mathcal{R}_{\infty}$-warmness with complexity $\widetilde{O}(d^{2}R^{2})$ for $R^{2}=\mathbb{E}_{\pi_{\mathcal{K}}}[|\cdot|^{2}]$. [KV25a] show that $\mathsf{TV}$-warmness can be achieved with the better complexity of $\widetilde{O}(d^{2}R^{3/2}\Lambda^{1/4}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}})$ . Currently, these are the only known guarantees, and there is no bound for intermediate warmness results, between $\mathsf{TV}$ and $\mathcal{R}_{\infty}$ (i.e., either very weak or strong warmness).
Towards genuine warmness.
Why do we even care about $\mathcal{R}_{q}$-warmness? Given the gap in known complexities between the two extreme settings discussed above, bridging this gap is theoretically intriguing in its own right. In our general setup, it is not clear whether generating $\mathcal{R}_{q}$-warmness without significant overhead is feasible, let alone whether one can tighten the complexities of warm-start generation and sampling from a warm start.
Moreover, $\mathsf{TV}$-warmness arguably serves only as a “pseudo” warm-start for downstream sampling, since subsequent sampling guarantees using this warmness are only in $\mathsf{TV}$-distance due to the necessity of applying triangle inequalities. In practice, this limitation leads to accumulation of sampling errors, resulting in somewhat cumbersome post-processing, such as replacing the original target accuracy $\varepsilon$ by a reduced accuracy $\varepsilon/m$, where $m$ is the number of times the triangle inequality is applied. While this does not significantly affect downstream guarantees for applications (since $\varepsilon^{-1}$ is inside a logarithm), it is a bit unsatisfactory and makes the analysis more complicated.
One might instead generate a $\mathcal{R}_{\infty}$-warm start and then apply the known $\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ guarantee (see Figure 1.1). However, generating $\mathcal{R}_{\infty}$-warmness incurs a higher query complexity than $\mathsf{TV}$-warmness, making it an unnecessarily expensive route to an $\mathcal{R}_{q}$-guarantee. Lastly, when combined with $\mathsf{PS}_{\textup{unif}}$, which achieves balanced guarantees, a $\mathcal{R}_{q}$-warmness result would yield clean end-to-end guarantees for uniform sampling in any desired $\mathcal{R}_{q}$-divergence.
[…]
Figure 1.1: Zeroth-order query complexities for logconcave sampling. Arrows indicate the sampling complexity from an $O(1)$-warm start. Complexities listed next to each left node indicate warmness-generation costs. The dagger sign ($\dagger$) indicates results for uniform sampling over convex bodies (membership oracle). Results without $\dagger$ hold for general logconcave distributions (evaluation oracle). Here, $\bar{R}:=R\vee 1$, $\bar{\lambda}:=\lambda\vee 1$, $\lambda:=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}:=\mathbb{E}_{\pi}[|\cdot|^{2}]$ (so $\lambda\leq R^{2}$). Note that $\mathsf{TV}$-warmness can be treated as $\mathcal{R}_{\infty}$-warmness at the cost of collapsing downstream guarantees to $\mathsf{TV}$-distance.
$\mathsf{TV}$-collapse.
A central difficulty in obtaining $\mathcal{R}_{q}$-warmness through annealing is that we do not yet know how to relay $\mathcal{R}_{q}$-warmness without compromise, leading nearly all previous work to encounter this $\mathsf{TV}$-collapse problem. To illustrate this, consider the annealing distributions $\mu_{0},\mu_{1},\mu_{2}$, where a sampler transitions from $\mu_{0}$ toward $\mu_{1}$. Markov-chain samplers for convex bodies are inherently approximate, producing a measure $\bar{\mu}_{1}$ close to but not exactly $\mu_{1}$, quantified via $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})\leq\varepsilon$ for some $q>1$. In the next phase, the sampler moves toward $\mu_{2}$ starting from the approximate measure $\bar{\mu}_{1}$, not $\mu_{1}$. Ideally, the complexity for this phase would directly depend on $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})$. Unfortunately, the lack of a triangle inequality for $\mathcal{R}_{q}$-divergence prevents us from ensuring $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})=O(1)$, even if $\mathcal{R}_{q}(\mu_{1}\mathbin{\|}\mu_{2})=O(1)$.
To circumvent this issue, prior studies have resorted to using the triangle inequality (or a coupling argument) for $\mathsf{TV}$-distance. Specifically, although $\bar{\mu}_{1}$ differs from $\mu_{1}$, annealing algorithms proceeds as if the initial distribution were $\mu_{1}$, not $\bar{\mu}_{1}$, when running a sampler toward $\mu_{2}$. Denoting the transition kernel of this sampler by $P$ and the output distribution of this sampler by $\bar{\mu}_{2}=\bar{\mu}_{1}P$, we can deduce that
$\lVert\bar{\mu}_{2}-\mu_{2}\rVert_{\mathsf{TV}}\leq\lVert\bar{\mu}_{2}-\mu_{1}P\rVert_{\mathsf{TV}}+\lVert\mu_{1}P-\mu_{2}\rVert_{\mathsf{TV}}\leq\lVert\bar{\mu}_{1}-\mu_{1}\rVert_{\mathsf{TV}}+\lVert\mu_{1}P-\mu_{2}\rVert_{\mathsf{TV}}\,$,
where the second inequality follows from the data-processing inequality (2.1). Since $\mathcal{R}_{q}$-divergence implies $\mathsf{TV}$-distance, and each term on the RHS is bounded by $\varepsilon$, this leads to the accumulation of error over $m$ phases: $\lVert\bar{\mu}_{m}-\mu_{m}\rVert_{\mathsf{TV}}\leq m\varepsilon$. This necessitates accuracy adjustments like replacing $\varepsilon$ with $\varepsilon/m$.
Boosting the Rényi divergence order.
The above discussion highlights the fundamental bottleneck arising from the absence of a triangle inequality for Rényi divergences. One promising direction to overcome this is to leverage a weak triangle inequality (1.3), namely, $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})\lesssim\mathcal{R}_{2q-1}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})+\mathcal{R}_{2q}(\mu_{1}\mathbin{\|}\mu_{2})$. As [KV25a] has already designed annealing schemes that can ensure $\mathcal{R}_{2q}(\mu_{1}\mathbin{\|}\mu_{2})=O(1)$, one might consider starting with a $\mathcal{R}_{2q-1}$-guarantee when sampling from $\mu_{1}$. However, this effectively amounts to requiring a $\mathcal{R}_{\Theta(2^{m}q)}$-guarantee for $\mu_{1}$-sampling when there are $m$ phases. Since $\mathsf{PS}$ has linear dependence on $q$, and prior annealing approaches require $m\gtrsim d^{1/2}$, such a “backtracking” analysis would potentially incur an $\Omega(2^{\sqrt{d}}q)$-multiplicative overhead in the final complexity.
Given the failure of this naïve approach, a pertinent question is whether one can ensure $\mathcal{R}_{2q-1}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})=O(1)$ given $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})<\varepsilon$. Since 2q1>q2q-1>q for $q>1$, this seemingly goes in the wrong direction — opposite the monotonic property of $\mathcal{R}_{q}$-divergence. This suggests an algorithmic question: can we boost the Rényi-divergence order algorithmically (i.e., can we transform $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)=O(1)$ into a stronger bound $\mathcal{R}_{2q-1}(\mu^{\prime}\mathbin{\|}\pi)=O(1)$)? Solving this question would enhance our theoretical toolkit, providing the final missing piece towards a comprehensive approach for convex body sampling and its applications.
Hypercontractivity of semigroups.
Looking back at diffusion for inspiration, perhaps the most relevant mathematical notion to tackle this question is Gross’ hypercontractivity of semigroups [GRO75] (see Proposition 3.1): for a semigroup $(P_{t})_{t\geq 0}$ reversible with respect to an invariant measure $\pi$, it is known that ($\mathsf{LSI}$) for $\pi$ is equivalent to the property that $\|P_{t}f\|_{L^{q(t)}(\pi)}\leq\|f\|_{L^{p}(\pi)}$ for any function $f\in L^{p}(\pi)$, where $q(t)$ is an increasing function with $q(0)=p$. For instance, the continuous-time Langevin diffusion satisfies this property, but it is not immediately clear whether such hypercontractivity is preserved upon discretization. Recent advancements, such as [CEL+24], showed that discretized Langevin processes retain hypercontractivity under LSI when the target $\pi$ is unconstrained and log-smooth.
Can we extend hypercontractivity algorithmically beyond such specialized settings? In particular, sampling with hard constraints—such as sampling from a convex body—or more general smoothness assumptions (e.g., $B_{1}\subseteq\mathcal{K}$) have resisted straightforward solutions. One might consider using reflected Langevin diffusions, but discretizations like the projected Langevin algorithm [BEL18] require access to a projection oracle for $\mathcal{K}$, and existing analytical guarantees similarly collapse to $\mathsf{TV}$-distance (beside incurring much higher complexity).
Hypercontractivity of the heat adjoint (§3).
To address this boosting question, we explore an alternative strategy involving a different Markov semigroup that can be algorithmically realized via $\mathsf{PS}$. Specifically, we aim to establish hypercontractivity for the $L^{2}(\nu)$-adjoint $Q_{t}$ of the heat semigroup $P_{t}$, under ($\mathsf{LSI}$) for the target measure $\nu$. Formally, we state the following hypercontractivity result for $Q_{t}$.
Theorem 1.2 (Hypercontractivity under heat flow).
Let $\nu$ be a probability measure satisfying ($\mathsf{LSI}$) with $C_{\mathsf{LSI}}(\nu)<\infty$. Let $\nu_{t}=\nu*\gamma_{t}$ for $t\geq 0$, and define $q:[0,\infty)\to(1,\infty)$ by
$\frac{q(t)-1}{p-1}=1+\frac{t}{C_{\mathsf{LSI}}(\nu)}\quad\text{for any }p\in(1,\infty)\,.$
Then, the $L^{2}(\nu)$-adjoint $Q_{t}$ of the heat semigroup satisfies
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f\rVert_{L^{p}(\nu)}\quad\text{for all }p\in(1,\infty),t\geq 0,\text{ and }f\in L^{p}(\nu)\,.$
(1) Proof sketch.
Overall, we follow a proof sketch in [BGL14, Theorem 5.2.3] along with technical tweaks. To begin, for any compactly supported, smooth, positive function $f$, we define the probability measure $\mathrm{d}\mu\propto f\,\mathrm{d}\pi$. Denoting the convolved measures $\mu_{t}:=\mu*\gamma_{t}$ and $\pi_{t}:=\pi*\gamma_{t}$, we introduce the functionals $\mathcal{F}(t,q):=\int(\nicefrac{{\mathrm{d}\mu_{t}}}{{\mathrm{d}\pi_{t}}})^{q}\,\mathrm{d}\pi_{t}$ and $\mathcal{G}(t):=\frac{1}{q(t)}\,\log\mathcal{F}(t,q(t))$. Our goal is to construct a suitable function $q(t)$ such that $\partial_{t}\mathcal{G}\leq 0$. Our proof proceeds by differentiating $\mathcal{G}$ in $t$, which involves computing $\partial_{t}\mathcal{F}$ and $\partial_{q}\mathcal{F}$. The time derivative $\partial_{t}\mathcal{F}$ follows from the general de Bruijn identity [KO25], while the exponent derivative $\partial_{q}\mathcal{F}$ gives rise to $\operatorname{Ent}_{\pi_{t}}$, where the log-Sobolev constant $C_{\mathsf{LSI}}(\pi_{t})$ comes into picture.
Combining these ingredients, we derive a differential inequality for $\mathcal{G}(t)$, and $\partial_{t}\mathcal{G}\leq 0$ holds provided that $C_{\mathsf{LSI}}(\pi_{t})\leq(q(t)-1)/(q(t)-1)^{\prime}$. Leveraging $C_{\mathsf{LSI}}(\pi_{t})\leq C_{\mathsf{LSI}}(\pi)+t$ [CHA04], we solve $C_{\mathsf{LSI}}(\pi)+t\leq(q(t)-1)/(q(t)-1)^{\prime}$, which holds when $q(t)-1=(1+\nicefrac{{t}}{{C_{\mathsf{LSI}}(\pi)}})\,(p-1)$. Altogether, these arguments establish the hypercontractivity of the heat adjoint $Q_{t}$. Lastly, we extend this result beyond the space of compactly supported, positive, smooth functions by employing standard approximation techniques in analysis, such as the dominated convergence theorem and Fatou’s lemma, thereby generalizing the statement to $L^{p}(\pi)$-space.
(2) Consequences.
An immediate consequence follows by selecting $f=\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$ for a probability measure $\mu\ll\nu$. Since $Q_{t}f=\frac{P_{t}(f\nu)}{P_{t}\nu}$, this yields $\lVert\mathrm{d}\mu_{t}/\mathrm{d}\nu_{t}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert\mathrm{d}\mu/\mathrm{d}\nu\rVert_{L^{p}(\nu)}$, which implies $\mathcal{R}_{q(t)}(\mu_{t}\mathbin{\|}\nu_{t})\leq\mathcal{R}_{p}(\mu\mathbin{\|}\nu).$ This result has algorithmic implications for $\mathsf{PS}$. Below, the constant $2$ can be replaced by any number larger than $1$.
Corollary 1.3 (Hypercontractivity of proximal sampler).
Let $\pi_{k}^{X}$ be the law of the $k$-th iterate returned by the $\mathsf{Proximal\ sampler}$ with step size $h>0$, initial $\pi_{0}^{X}$, and target $\pi^{X}$. Given $1<p\leq q$, we have $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\mathcal{R}_{p}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$ after $N\lesssim h^{-1}C_{\mathsf{LSI}}(\pi^{X})\log_{2}\frac{q-1}{p-1}$ iterations. Moreover, given $\varepsilon>0$ and $q\geq q_{0}>1$, we have $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
$N\lesssim h^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\pi_{0}^{X}\mathbin{\|}\pi^{X})}{(2\wedge q_{0}-1)\,\varepsilon}\quad\text{iterations\,.}$
This result translates hypercontractivity into an algorithmic outcome, enabling efficient implementation of warm-start generation. Moreover, it leads to $\widetilde{O}(\kappa d\log\frac{q}{\varepsilon})$-mixing of $\mathsf{PS}$ in the well-conditioned setting (Proposition 3.2), improving the linear dependence of $\mathsf{PS}$ on $q$ [CCS+22].
Streamlined warm-start generation (§4).
We simplify the design and analysis of a known Gaussian-annealing scheme for warm-start generation, incorporating the hypercontractivity established for $\mathsf{PS}$. This streamlined approach offers an elegant and clean picture for warm-start generation via annealing schemes.
As with $\mathsf{PS}_{\textup{unif}}$, we propose and analyze its Gaussian variant in §4.1, denoted by $\mathsf{PS}_{\textup{Gauss}}$, specifically tailored for truncated Gaussian distributions $\gamma_{\sigma^{2}}|_{\mathcal{K}}$. Leveraging the hypercontractivity result for $\mathsf{PS}$ and the bound $C_{\mathsf{LSI}}(\gamma_{\sigma^{2}}|_{\mathcal{K}})\leq\sigma^{2}$, we derive an enhanced guarantee for $\mathsf{PS}_{\textup{Gauss}}$. Below, $B_{1}(0)$ denotes the $\ell_{2}$ unit ball centered at the origin.
Theorem 1.4.
Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with $B_{1}(0)\subset\mathcal{K}$ given by a membership oracle, and let $\mu\propto\gamma_{\sigma^{2}}\cdot\mathds{1}_{\mathcal{K}}$ be the Gaussian truncated to $\mathcal{K}$. Given $\varepsilon\in(0,1/100)$, initial distribution $\mu_{0}$, any $q\geq 2\vee\widetilde{\Omega}(\log(d^{2}\sigma^{2}\log\frac{1}{\varepsilon}))$ such that $q\geq q_{0}:=2\vee\widetilde{O}(\log(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\varepsilon}))$ with $M_{q_{0}}=\lVert\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\mu}\rVert_{L^{q_{0}}(\mu)}\leq 10$, consider the algorithm that runs $\mathsf{PS}_{\textup{Gauss}}$ initialized at $\mu_{0}$ and restarts if $\mathsf{PS}_{\textup{Gauss}}$ fails before completing $N=\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon})$ iterations. There exist choices of $h$ and $\tau$ such that this algorithm restarts with probability at most $1/50$, and its output $Z$ satisfies $\mathcal{R}_{q}(\operatorname{law}Z\mathbin{\|}\mu)\leq\varepsilon$, using $\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log^{5}\frac{1}{\varepsilon})$ membership queries in expectation.
Equipped with $\mathsf{PS}_{\textup{Gauss}}$, our warm-start generation process is very simple: (1) sample from $\mu_{1}\propto\pi_{\mathcal{K}}\gamma_{d^{-1}}$ by rejection sampling with the Gaussian proposal $\gamma_{d^{-1}}$, and (2) while $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ toward $\mu_{i+1}\propto\pi_{\mathcal{K}}\gamma_{\sigma_{i+1}^{2}}$ using $\mathsf{PS}_{\textup{Gauss}}$, where $\mathcal{R}_{\widetilde{O}(1)}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$ and $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,(1+\nicefrac{{\sigma_{i}}}{{R}})$.
This simpler algorithm eliminates pre-processing steps, such as truncation of the convex body, and post-processing, such as adjusting the target accuracy. Moreover, a key practical advantage of this streamlined method is the introduction of early-stopping at $\sigma^{2}\approx qR\Lambda^{1/2}\log^{1/2}d$. Consequently, faster sampling phases beyond $\sigma^{2}\gtrsim R\Lambda^{1/2}\log^{1/2}d$ needed in [KV25a] become redundant. On the theoretical side, hypercontractivity enables the propagation of $\mathcal{R}_{q}$-warmness throughout the annealing scheme, leading to the following comprehensive result for warm-start generation:
Theorem 1.5.
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ given by a membership oracle with $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$, . For any $\varepsilon_{\textup{unif}}\in(0,1/100)$ and $q\geq 2\vee\widetilde{\Omega}\bigl(\log\frac{dR}{\varepsilon_{\textup{unif}}}\bigr)$, there exists an algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)\leq 2$, using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$ membership queries in expectation.
Combining this with Theorem 1.1, we obtain the query complexity of convex-body sampling from scratch, with $\mathcal{R}_{q}$-divergence output guarantees for any $q\geq 1$.
Corollary 1.6.
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ given by a membership oracle with $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$. For any $\varepsilon\in(0,1/100)$ and $q\geq 1$, there exists an algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4}\log\frac{1}{\varepsilon}+qd^{2}\Lambda\log^{7}\frac{1}{\varepsilon})$ membership queries in expectation.
Streamlined analysis of Gaussian annealing.
In our analysis of Gaussian annealing (§4.2), we present a principled and streamlined approach to bounding the closeness between successive annealing distributions. Whereas prior works relied on localization arguments (with long calculations) to establish such closeness, we show that this property follows more cleanly from LSI, combined with a mollification argument that allows for a limiting process.
To formalize this idea, we recall a Rényi version of LSI established in [VW23], which states that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\nicefrac{{qC_{\mathsf{LSI}}(\pi)}}{{2}}\,\mathsf{RFI}_{q}(\mu\mathbin{\|}\pi)$ for any $q>1$ and smooth $\mu/\pi$, where $\mathsf{RFI}_{q}(\mu\mathbin{\|}\pi)=q\,\mathbb{E}_{\pi}[(\frac{\mathrm{d}\mu}{\mathrm{d}\pi})^{q}|\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\pi}|^{2}]/\mathbb{E}_{\pi}[(\frac{\mathrm{d}\mu}{\mathrm{d}\pi})^{q}]$. When bounding $\mathcal{R}_{q}(\gamma_{\sigma^{2}}|_{\mathcal{K}}\mathbin{\|}\gamma_{\sigma^{2}(1+\alpha)}|_{\mathcal{K}})$ for a convex body $\mathcal{K}$ and some $\alpha>0$, a naïve calculation—ignoring the truncation to $\mathcal{K}$—suggests that one can take $\alpha\asymp\nicefrac{{\sigma}}{{qR}}$ to make it $O(1)$. This heuristic choice recovers the scaling behavior established in earlier results.
To rigorously account for the effects of truncation to $\mathcal{K}$, we employ a mollification argument using a standard mollifier $\eta$ supported on the unit ball (see (4.3)). For $\varepsilon>0$, we define $\eta_{\varepsilon}(x):=\varepsilon^{-d}\,\eta(x/\varepsilon)$, and convolve both truncated Gaussians with $\eta_{\varepsilon}$. Since convolution with $\eta_{\varepsilon}$ yields a smooth approximation $f_{\varepsilon}:=f*\eta_{\varepsilon}$ that converges to $f$ almost everywhere as $\varepsilon\to 0$ for any locally integrable function $f$, this mollification step allows us to apply the Rényi-LSI to the regularized densities and pass to the limit, thereby recovering the desired closeness bound.
An additional benefit of this approach is that it enables a clean justification for early-stopping in the annealing schedule. Unlike known annealing algorithms, we can terminate annealing when the variance $\sigma^{2}$ reaches approximately $qR\Lambda^{1/2}$, at which point $\mathcal{R}_{q}(\gamma_{\sigma^{2}}|_{\mathcal{K}}\mathbin{\|}\pi_{\mathcal{K}})=O(1)$. This follows directly from a Rényi version of PI (see ($\mathsf{R\text{-}PI}$)), combined with the mollification argument described above. Overall, this approach simplifies the analysis and unifies the reasoning behind closeness and early-stopping in Gaussian annealing.
1.1.3 Logconcave sampling under an evaluation oracle
Our uniform-sampling results extend naturally to general logconcave distributions $\pi\propto e^{-V}$, given access to an evaluation oracle for a convex function $V:\mathbb{R}^{d}\to\mathbb{R}\cup\{\infty\}$. In this setting, the regularity condition $B_{1}\subset\mathcal{K}$ translates to a condition on a level set of $V$, namely, that the ground set $\mathsf{L}_{\pi,g}:=\{x\in\mathbb{R}^{d}:V(x)-\min V\leq 10d\}$ contains a unit ball. This generalization is justified by the fact that $\mathsf{L}_{\pi,g}$ takes up a constant fraction of the $\pi$-measure (see [KV25b, Lemma 2.6]). We emphasize that any logconcave distribution satisfies this condition after rescaling.
The logconcave sampling problem has been studied as an algorithmic problem since Applegate and Kannan [AK91], who used the $\mathsf{Grid}\text{ }\mathsf{walk}:\mathcal{R}_{\infty}\to\mathsf{TV}$, establishing polynomial complexity in the dimension under the assumption of Lipschitzness of the function over the support. Later, Lovász and Vempala [LV07] generalized classical algorithms such as the $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$ to this setting by incorporating a Metropolis filter. They showed that after a suitable pre-processing of $\pi$, $\mathsf{Ball\ walk},\mathsf{Hit\text{-}and\text{-}Run}:\mathcal{R}_{\infty}\to\mathsf{TV}$ with query complexity $\widetilde{O}(\varepsilon^{-4}d^{2}R^{2}M_{\infty}^{4}\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$. Furthermore, [LV06a] showed that the complexity of $\mathsf{Hit\text{-}and\text{-}Run}:\mathcal{R}_{2}\to\mathcal{R}_{2}$ can be improved to $\widetilde{O}(d^{2}R^{2}\operatorname{polylog}\nicefrac{{M_{2}}}{{\varepsilon}})$.
Recent work [KV25b, KV25a] revisited this problem by adapting $\mathsf{PS}$ to the general logconcave setting. Although $\mathsf{PS}$ can in principle be extended to any distribution where $\pi^{X|Y=y}\propto\pi^{X}(\cdot)\,\gamma_{h}(\cdot-y)$ can be sampled (e.g., when given access to a proximal oracle in the log-smooth setting), it is unclear how to implement this step for arbitrary $\pi^{X}$. To address this, [KV25b] introduced the exponential lifting technique: for $\pi^{X}\propto e^{-V}$, consider the lifted distribution
$\mathrm{d}\pi^{X,T}(x,t)\propto e^{-dt}\,\mathds{1}_{\mathcal{K}}(x,t)\,\mathrm{d}x\mathrm{d}t\,,\quad\text{where }\mathcal{K}:=\{(x,t)\in\mathbb{R}^{d}\times\mathbb{R}:V(x)\leq dt\}\,.$ ($\mathsf{exp\text{-}lifting}$)
Here, the $X$-marginal of $\pi^{X,T}$ is exactly $\pi^{X}$, and $\mathcal{K}$ is convex whenever $\pi^{X}$ is logconcave. Thus, logconcave sampling reduces to the simpler problem of sampling from a logconcave distribution with a linear potential over a convex set. We then apply $\mathsf{PS}$ to this exponential distribution, denoted by $\mathsf{PS}_{\textup{exp}}$: for $z=(x,t)\in\mathbb{R}^{d+1}$ and $y\in\mathbb{R}^{d+1}$, it alternates $y\sim\pi^{Y|Z=z}=\mathcal{N}(z,hI_{d+1})$ and $z\sim\pi^{Z|Y=y}\propto\mathcal{N}(y-h\alpha,hI_{d+1})|_{\mathcal{K}}$ for $\alpha=de_{d+1}$, where the second step is implemented via rejection sampling, using the proposal $\mathcal{N}(y-h\alpha,hI_{d+1})$ with threshold $\tau$.
In [KV25b], it was shown that $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ with query complexity $\widetilde{O}(qM_{\infty}d^{2}(\Lambda\vee 1)\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$ and $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ with complexity $\widetilde{O}(M_{\infty}d^{2}(d\vee R^{2})\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$. As in the uniform sampling case, the order of initial warmness was relaxed to $\mathcal{R}_{c}$ with $c=\widetilde{O}(1)$ in [KV25a], where $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{c}\to\mathcal{R}_{2}$ achieves query complexity $\widetilde{O}(M_{c}d^{2}(\Lambda\vee 1)\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$. However, it remained open to balance the divergence order of initial warmness and final output distribution guarantee.
Extension to logconcave sampling (§5).
Our $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ with restart aligns the divergence order of initial warmness with that of the final error guarantee, from an $O(1)$-warm start, at an additive cost of $qd^{2}$ queries. This refines a long line of work on sampling in the classical setting. See §5.1 for further details.
Theorem 1.7 (Zeroth-order logconcave sampling from a warm start).
For a convex function $V:\mathbb{R}^{d}\to\mathbb{R}$, given by an evaluation oracle, let $\pi\propto e^{-V}$ be the logconcave distribution over $\mathbb{R}^{d}$ with $B_{1}\subset\mathsf{L}_{\pi,g}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$. Given $\varepsilon>0$, an initial distribution $\pi_{0}$, and $q\geq 2\vee\widetilde{\Omega}(\log\{d^{2}\,(\Lambda\vee 1)\log\frac{1}{\varepsilon}\})$ such that $M_{q}=\lVert\frac{\mathrm{d}\pi_{0}}{\mathrm{d}\pi}\rVert_{L^{q}(\pi)}\leq 10$, consider the algorithm that runs $\mathsf{PS}_{\textup{exp}}$ initialized at $\pi_{0}$ and restarts if $\mathsf{PS}_{\textup{exp}}$ fails before completing $N=\widetilde{O}(qd^{2}\,(\Lambda\vee 1)\log^{2}\frac{1}{\varepsilon})$ iterations. There exists choices of $h$ and $\tau$ such that this algorithm restarts with probability at most $1/50$, and its output $Z$ satisfies $\mathcal{R}_{q}(\operatorname{law}Z\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(qd^{2}\,(\Lambda\vee 1)\log^{3}\frac{1}{\varepsilon})$ evaluation queries in expectation.
We remark on a connection to the well-conditioned setting. In this case, the ground set contains a ball of radius $\Theta((d/\beta)^{1/2})$ centered at $\arg\min V$. Since $\Lambda\leq\alpha^{-1}$ by the Brascamp–Lieb or Lichnerowicz inequality, $\mathsf{PS}_{\textup{exp}}$ uses $\widetilde{O}(q\,(\kappa d+d^{2})\log\nicefrac{{1}}{{\varepsilon}})$ queries from an $O(1)$-warm start in $\mathcal{R}_{q}$. With the additional $qd^{2}$ queries, this matches the known $\widetilde{O}(q\kappa d\log\nicefrac{{1}}{{\varepsilon}})$-complexity results for zeroth-order samplers in this setting which additionally use a proximal oracle. See §B for further details.
Annealing approaches for uniform sampling have also been extended to arbitrary logconcave distributions [LV06a, KV25b, KV25a]. In the general case, the best-known complexity for warm-start generation is comparable to that of the convex-body case for $\mathsf{TV}$ and $\mathcal{R}_{\infty}$-warmness, with algorithms requiring additional $d^{2}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}}$ queries. In §5.2, we extend our refined warm-start generation to general logconcave distributions.
Theorem 1.8.
For a convex function $V:\mathbb{R}^{d}\to\mathbb{R}$, given by an evaluation oracle, let $\pi\propto e^{-V}$ be the logconcave distribution over $\mathbb{R}^{d}$ with $B_{1}(0)\subset\mathsf{L}_{\pi,g}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$. For any $\varepsilon_{\exp}\in(0,\frac{1}{10})$ and $q\geq 2\vee\widetilde{\Omega}(\log\frac{dD}{\varepsilon_{\exp}})$, there exists an annealing algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)\leq 2$, using $\widetilde{O}(d^{2.5}+qd^{2}\,(R^{3/2}\vee 1)(\Lambda^{1/4}\vee 1))$ evaluation queries in expectation.
Combining the two theorems above, we establish the query complexity of sampling from an arbitrary logconcave distribution in $\mathbb{R}^{d}$ with $\mathcal{R}_{q}$-divergence guarantees, assuming only an evaluation oracle.
Corollary 1.9.
For a convex function $V:\mathbb{R}^{d}\to\mathbb{R}$, given by an evaluation oracle, let $\pi\propto e^{-V}$ be the logconcave distribution over $\mathbb{R}^{d}$ with $B_{1}(0)\subset\mathsf{L}_{\pi,g}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$. For any $\varepsilon\in(0,\frac{1}{10})$ and $q\geq 1$, there exists an algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(d^{2.5}+qd^{2}R^{3/2}\,(\Lambda^{1/4}\vee 1)\log\frac{1}{\varepsilon}+qd^{2}\,(\Lambda\vee 1)\log^{4}\frac{1}{\varepsilon})$ evaluation queries in expectation.
1.1.4 Obstruction to further acceleration
A natural question is whether the annealing complexity in Corollary 1.9 can be improved beyond $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$. For isotropic logconcave distributions in particular, given the $d^{3}$ query complexity of [KV25b] for $\mathcal{R}_{\infty}$-guarantees and our $qd^{2.75}$ complexity for $\mathcal{R}_{q}$-guarantees, it is natural to ask whether the $d^{2.5}$ iteration complexity of the $\mathsf{Speedy\ walk}$ established in [LV24] can be matched as a query complexity, at least for $\mathsf{TV}$-guarantees (note that the speedy walk is an abstraction that, in general, does not have an efficient implementation).
Towards this goal, [KV25a] conjectured that for the uniform distribution over any isotropic convex body $\pi$ in $\mathbb{R}^{d}$, the Gaussian-tilted distribution $\pi\gamma_{t}$ satisfies $\lVert\operatorname{cov}\pi\gamma_{t}\rVert=\widetilde{O}(1)$ for all $t>0$, a property called $\widetilde{O}(1)$-quadratic-tilt stability. Under this conjecture, one can afford more aggressive annealing steps of size $\sigma_{i+1}^{2}=\sigma_{i}^{2}(1+\sigma_{i}^{2}/R)$ (instead of $\sigma_{i+1}^{2}=\sigma_{i}^{2}(1+\sigma_{i}/R)$), leading to an improved warm-start generation complexity of $\widetilde{O}(d^{2.5})$ for isotropic logconcave distributions (see §6.1). We verify in §A that structured bodies—including the hypercube, the simplex, and any isotropic convex body of revolution—satisfy $O(1)$-quadratic-tilt stability, confirming the desired $\widetilde{O}(d^{2.5})$ query complexity of sampling from these structured distributions.
However, in §6.2 we disprove the conjecture in general, establishing a fundamental obstruction to this line of acceleration. Specifically, adapting a construction of Bizeul [BIZ26], we show that there exists an isotropic convex body whose Gaussian-tilted distribution exhibits a peak variance of $d^{1/3}$ (Proposition 6.3), disproving the conjecture and proving a lower bound for the Gaussian annealing approach that has been central to this line of work since [CV14, CV15, CV18].
1.2 Preliminaries
Notation.
For $t>0$, we reserve $\gamma_{t}$ for the centered Gaussian distribution with covariance matrix $tI_{d}$. The indicator function of a set $S\subseteq\mathbb{R}^{d}$ is denoted by $\mathds{1}_{S}(x):=[x\in S]$, and $\mu|_{S}$ refers to a distribution $\mu$ truncated to $S$ (i.e., $\mu|_{S}\propto\mu\cdot\mathds{1}_{S}$). For two probability measures $\mu,\pi$, we use $\mu\pi$ to denote the new distribution with density proportional to $\mu\pi$.
Both $a\lesssim b$ and $a=O(b)$ mean $a\leq cb$ for a universal constant $c>0$. $a=\Omega(b)$ means $a\gtrsim b$, and $a\asymp b$ means $a=O(b)$ and $a=\Omega(b)$. Lastly, $a=\widetilde{O}(b)$ means $a=O(b\operatorname{polylog}b)$. For a positive semi-definite matrix $\Sigma$, $\lVert\Sigma\rVert$ denote the operator norm of $\Sigma$.
Log-concavity.
We call a function $f:\mathbb{R}^{d}\to[0,\infty)$ logconcave if $\log f$ is convex in $\mathbb{R}^{d}$, and a probability measure $\pi$ (or distribution) logconcave if it has a logconcave density function with respect to the Lebesgue measure. We abuse notation by using the same symbol for a distribution and density. We assume that any logconcave distributions considered in this work are non-degenerate; otherwise, we could simply work on an affine subspace on which degenerate distributions are supported. For $t\geq 0$, $\pi$ is called $t$-strongly logconcave if $\log\pi$ is $t$-strongly convex (i.e., $\log\pi-\frac{t}{2}\,|\cdot|^{2}$ is convex). Clearly, log-concavity is preserved under multiplication, and a classical result by Prékopa and Leindler ensures that the convolution also preserves log-concavity. A distribution is called isotropic if its barycenter is at the origin and has the identity covariance matrix. Note that as the logconcave distributions decay exponentially fast at infinity, they have finite moments of all orders.
Probability divergences and distances.
Let $\mu$ and $\nu$ denote probability measures over $\mathbb{R}^{d}$, and $\varphi:\mathbb{R}_{+}\to\mathbb{R}$ be a convex function with $\varphi(1)=0$. The $\varphi$-divergence of $\mu$ toward $\nu$ with $\mu\ll\nu$ is defined as
$D_{\varphi}(\mu\mathbin{\|}\nu):=\int\varphi\bigl(\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr)\,\mathrm{d}\nu\,.$
The total variation ($\mathsf{TV}$) distance, the $\mathsf{KL}$-divergence, and $\chi^{q}$-divergence for $q\in(1,\infty)$ can be recovered through $\varphi(x)=\frac{1}{2}\,|x-1|$, $x\log x$, and $x^{q}-1$, respectively. The $q$-Rényi divergence is defined as
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{1}{q-1}\,\log\bigl(1+\chi^{q}(\mu\mathbin{\|}\nu)\bigr)=\frac{1}{q-1}\,\log\,\Bigl\|\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr\|_{L^{q}(\nu)}^{q}\,.$ (1.2)
The Rényi-infinity divergence is defined as $\mathcal{R}_{\infty}(\mu\mathbin{\|}\nu):=\log\operatorname{ess\,sup}_{\nu}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$. A weak triangle inequality holds for Rényi divergence: for any $q>1,\lambda\in(0,1)$, and probability measures $\mu,\nu,\pi$, it holds that
$\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\frac{q-\lambda}{q-1}\,\mathcal{R}_{\frac{q}{\lambda}}(\mu\mathbin{\|}\nu)+\mathcal{R}_{\frac{q-\lambda}{1-\lambda}}(\nu\mathbin{\|}\pi)\,.$ (1.3)
For $\lambda=q/(2q-1)$, we have $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\frac{2q}{2q-1}\,\mathcal{R}_{2q-1}(\mu\mathbin{\|}\nu)+\mathcal{R}_{2q}(\nu\mathbin{\|}\pi)$. We refer readers to [vH14, MIR17] for more properties of the Rényi divergence.
2 Uniform sampling from convex bodies with balanced Rényi divergence
We begin by describing our approach for the general proximal sampler ($\mathsf{PS}$) and then focus specifically on $\mathsf{PS}_{\textup{unif}}$ for uniform sampling in §2.2, which was proposed and analyzed in [KVZ24]. $\mathsf{PS}$ with step size $h$ alternates between two steps: starting with $k=0$ and $x_{0}\sim\pi_{0}^{X}$, (i) sample $y_{k+1}\sim\pi^{Y|X=x_{k}}$ and (ii) sample $x_{k+1}\sim\pi^{X|Y=y_{k+1}}$. Precisely, let $P_{h}$ and $Q_{h}$ be the transition kernels of the forward and backward steps, respectively, defined as
$P_{h}(x,\mathrm{d}y)=\pi^{Y|X=x}=\mathcal{N}(x,hI_{d})\,,\qquad Q_{h}(y,\mathrm{d}x)=\pi^{X|Y=y}\propto\pi^{X}(\cdot)\,\gamma_{h}(\cdot-y)\,.$
Assume that there exists a suitable proposal distribution $\mu_{y}$ such that $(\mu_{y}\mid\text{accepted})=Q_{h}(y,\cdot)$. For uniform sampling, we take $\mu_{y}=\mathcal{N}(y,hI_{d})$ and accept the proposal if it lies inside $\mathcal{K}$. Given this setup, we implement the backward step via rejection sampling with proposal $\mu_{y_{k+1}}$. If the rejection loop exceeds $\tau$ trials, we declare failure and restart from scratch with a new sample from $\pi_{0}^{X}$.
Kook and Vempala [KV25a] obtained the query complexity of $\mathsf{PS}$ for obtaining a uniform sample whose law is $\varepsilon$-close to the target $\pi^{X}$ in $\mathcal{R}_{2}$-divergence, assuming the warmness is given in $\mathcal{R}_{q}$ with $q\geq\widetilde{O}(1)$. We refine their analysis to obtain the query complexity of $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ given $\lVert\mathrm{d}\pi_{0}^{X}/\mathrm{d}\pi^{X}\rVert_{L^{q}(\pi^{X})}\leq 10$ (i.e., an $O(1)$-warm start in $\mathcal{R}_{q}$). To this end, we decouple the analysis into two components: (1) a mixing analysis and (2) the query complexity of the backward step. The first part determines how many iterations $N$ are required for the sample $X_{N}$ to be $\varepsilon$-close to $\pi^{X}$. The second determines how many queries $\mathsf{PS}$ uses in each iteration. Therefore, the total query complexity of the algorithm is the product of these two components.
2.1 Mixing analysis
The data-processing inequality (DPI) is a fundamental tool in information theory. For any probability measures $\mu,\nu$, Markov kernel $P$, $\varphi$-divergence $D_{\varphi}$, and $q\in(1,\infty]$, it holds that
$D_{\varphi}(\mu P\mathbin{\|}\nu P)\leq D_{\varphi}(\mu\mathbin{\|}\nu)\,,\qquad\quad\text{and}\qquad\quad\mathcal{R}_{q}(\mu P\mathbin{\|}\nu P)\leq\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\,$, (2.1)
where $\mu P:=\int P(x,\cdot)\,\mu(\mathrm{d}x)$ denotes the distribution obtained by applying one step of $P$ starting from $\mu$. This implies that for $D\in\{D_{\varphi},\mathcal{R}_{q}\}$ and any $k\geq 0$,
$D(\pi_{k+1}^{X}\mathbin{\|}\pi^{X})=D(\pi_{k+1}^{Y}Q_{h}\mathbin{\|}\pi^{Y}Q_{h})\leq D(\pi_{k+1}^{Y}\mathbin{\|}\pi^{Y})=D(\pi_{k}^{X}P_{h}\mathbin{\|}\pi^{X}P_{h})\leq D(\pi_{k}^{X}\mathbin{\|}\pi^{X})\,.$
However, this alone is not sufficient to establish a mixing rate. It turns out that adding an independent Gaussian random variable to $X_{k}$ (i.e., convolving $\operatorname{law}X_{k}$ with $\gamma_{h}$) causes $\operatorname{law}X_{k}$ to contract directly toward $\pi^{X}*\gamma_{h}=\pi^{Y}$, with the contraction rate governed by $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$.
This phenomenon, known as the strong data-processing inequality (SDPI) for the Gaussian channel [AG76], has been extensively studied in information theory [PW16, CPW18]. Given a $\varphi$-divergence $D$, the contraction coefficient is defined as
$\eta_{D}(\mu,t):=\sup_{\nu:\,D(\nu\mathbin{\|}\mu)\in\mathbb{R}_{>0}}\frac{D(\nu_{t}\mathbin{\|}\mu_{t})}{D(\nu\mathbin{\|}\mu)}\,$,
where $\nu_{t}:=\nu*\gamma_{t}$ denotes the probability measure obtained by convolving $\nu$ with $\gamma_{t}$. Note that $0\leq\eta_{D}(\mu,t)\leq 1$ by the DPI.
Bounding $\eta_{D}$ relies on the well-known de Bruijn identity, $\partial_{t}\mathsf{KL}(\nu_{t}\mathbin{\|}\mu_{t})=-\frac{1}{2}\,\mathsf{FI}(\nu_{t}\mathbin{\|}\mu_{t})$, where the relative Fisher information is defined as $\mathsf{FI}(\nu\mathbin{\|}\mu):=\mathbb{E}_{\nu}[|\nabla\log\frac{\mathrm{d}\nu}{\mathrm{d}\mu}|^{2}]$; see rigorous proofs by Barron [BAR84] and Klartag and Ordentlich [KO25] (for its generalizations). Other derivations can be found in [GUO09, CCS+22], under the assumption that the probability measures involved are sufficiently smooth to justify differentiation under the integral sign and the vanishing of boundary terms in integration by parts.
Combining this identity with, for example, ($\mathsf{LSI}$) for $\mu$ (i.e., $\mathsf{KL}(\nu_{t}\mathbin{\|}\mu_{t})\leq\nicefrac{{C_{\mathsf{LSI}}(\mu_{t})}}{{2}}\,\mathsf{FI}(\nu_{t}\mathbin{\|}\mu_{t})$) and $C_{\mathsf{LSI}}(\mu_{t})\leq C_{\mathsf{LSI}}(\mu)+t$ [CHA04], one obtains a differential inequality in $\mathsf{KL}$; solving it yields
$\eta_{\mathsf{KL}}(\mu,t)\leq\bigl(1+\frac{t}{C_{\mathsf{LSI}}(\mu)}\bigr)^{-1}\,.$
Under similar regularity assumptions, the SDPI for Rényi divergence under ($\mathsf{PI}$), ($\mathsf{LSI}$), and Latała–Oleszkiewicz inequality was studied in [CCS+22] and further extended to constrained probability measures in [KVZ24]. These results were rigorously unified in [KO25] under a $\varphi$-Sobolev inequality [CHA04] (a generalization of ($\mathsf{PI}$) and ($\mathsf{LSI}$)), with regularity assumptions explicitly identified. We refer the reader to [KO25, Regularity conditions] for a detailed account.
Specifically, [CCS+22] showed that under the regularity assumptions, for any $q\geq 2$ and $\nu\ll\mu$, the following bounds hold:
[…]
and $\chi^{2}(\nu_{t}\mathbin{\|}\mu_{t})\leq(1+t/C_{\mathsf{PI}}(\mu))^{-1}\,\chi^{2}(\nu\mathbin{\|}\mu)$. Applying this result to the forward step, along with the DPI for the backward step, implies that $\mathsf{PS}$ achieves $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ with
$N\lesssim qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{M_{q}}{\varepsilon}\,,\quad\text{where }M_{q}=\Bigl\|\frac{\mathrm{d}\pi_{0}^{X}}{\mathrm{d}\pi^{X}}\Bigr\|_{L^{q}(\pi^{X})}\,.$
We provide a version of this result in terms of $\chi^{q}$-divergence (generalizing the $\chi^{2}$-result of [CCS+22]), as working with $\chi^{q}$-divergence yields cleaner expressions and a direct bound on the contraction coefficient $\eta_{\chi^{q}}$ under ($\mathsf{PI}$). Our proof follows the argument sketched above, combining the general de Bruijn identity in terms of $\chi^{q}$-divergence with ($\mathsf{PI}$). Below, we note that any logconcave probability measure satisfies ($\mathsf{PI}$) with $C_{\mathsf{PI}}<\infty$, as established in [KLS95, BOB99].
Lemma 2.1.
Let $\mu$ be a logconcave probability measure over $\mathbb{R}^{d}$. For any $q\geq 2$, it holds that
$\eta_{\chi^{q}}(\mu,t)\leq\bigl(1+\frac{t}{C_{\mathsf{PI}}(\mu)}\bigr)^{-2/q}\quad\text{for }t>0\,.$
To prove this result, we recall the general version of de Bruijn’s identity from [KO25, Proposition 1.12].
Proposition 2.2 (General de Bruijn’s identity).
Let $\mu$ and $\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\chi^{q}(\nu\mathbin{\|}\mu)<\infty$ and $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$. Then, for any $t>0$ and $q>1$,
$\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})=-\frac{q\,(q-1)}{2}\,\mathbb{E}_{\mu_{t}}\Bigl[\bigl(\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\bigr)^{q}\,\bigl|\nabla\log\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\bigr|^{2}\Bigr]=-\frac{q-1}{2}\,\Bigl\|\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\Bigr\|_{L^{q}(\mu_{t})}^{q}\,\mathsf{RFI}_{q}(\nu_{t}\mathbin{\|}\mu_{t})\,$,
where $\mathsf{RFI}_{q}(\nu\mathbin{\|}\mu)=q\,\mathbb{E}_{\mu}[(\frac{\mathrm{d}\nu}{\mathrm{d}\mu})^{q}\,|\nabla\log\frac{\mathrm{d}\nu}{\mathrm{d}\mu}|^{2}]/\mathbb{E}_{\mu}[(\frac{\mathrm{d}\nu}{\mathrm{d}\mu})^{q}]$ denotes the $q$-Rényi Fisher information.
We also recall a Rényi version of ($\mathsf{PI}$) from [VW23, Lemma 9] that for $q\geq 2$ and probability measure $\nu$ and $\mu$ with $\nu\ll\mu$ and smooth $\nu/\mu$,
$1-\exp\bigl(-\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\bigr)\leq\frac{qC_{\mathsf{PI}}(\mu)}{4}\,\mathsf{RFI}_{q}(\nu\mathbin{\|}\mu)\,.$ ($\mathsf{R\text{-}PI}$)
Proof of Lemma 2.1.
We combine the general de Bruijn identity with the Rényi PI. For the former, we note that $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$ since ($\mathsf{PI}$) ensures exponential integrability [BU83, GM83] (i.e., $\int\exp(s\,|\cdot|)\,\mathrm{d}\mu<\infty$ for $s^{2}<4/C_{\mathsf{PI}}(\mu)$; see [BGL14, Proposition 4.4.2]). For the latter, $\nu_{t}$ and $\mu_{t}$ are smooth by an elementary property of the heat kernel.
Using $\chi^{q}+1=\lVert\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\rVert_{L^{q}(\mu_{t})}^{q}:=L^{q}\geq 1$ and combining the two, we deduce that
$\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq-\frac{2}{qC_{\mathsf{PI}}(\mu_{t})}\,(q-1)\,L^{q}(1-L^{-\frac{q}{q-1}})\,.$
Since $(q-1)\cdot x\,(1-x^{-1/(q-1)})\geq x-1$ for $x\geq 1$, we obtain
$\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq-\frac{2}{qC_{\mathsf{PI}}(\mu_{t})}\,(L^{q}-1)=-\frac{2}{qC_{\mathsf{PI}}(\mu_{t})}\,\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\,.$
Using $C_{\mathsf{PI}}(\mu_{t})\leq C_{\mathsf{PI}}(\mu)+t$ and solving the differential inequality yields
$\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq\bigl(1+\frac{t}{C_{\mathsf{PI}}(\mu)}\bigr)^{-2/q}\,\chi^{q}(\nu\mathbin{\|}\mu)\,$,
which completes the proof. ∎
Remark 2.3 (Rényi version).
Under the same condition $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$, one can rederive (2.2) by following the argument of [CCS+22]. Using the identity $\mathcal{R}_{q}=\frac{1}{q-1}\,\log(1+\chi^{q})$ together with the chain rule and de Bruijn’s identity, we obtain:
$\partial_{t}\mathcal{R}_{q}(\nu_{t}\mathbin{\|}\mu_{t})=\frac{1}{q-1}\,\frac{\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})}{L^{q}}=-\frac{1}{2}\,\mathsf{RFI}_{q}(\nu_{t}\mathbin{\|}\mu_{t})\,.$
When applying the Rényi PI, the inequality $1-e^{-x}\geq 1/2$ for $x\geq 1$ is used when $\mathcal{R}_{q}\geq 1$, and $1-e^{-x}\geq x/2$ for $x\leq 1$ when $\mathcal{R}_{q}\leq 1$. These control the dissipation of the Rényi divergence in each regime.
The mixing time of $\mathsf{PS}$ follows as a direct consequence of the contraction result above. We ignore via the DPI the additional contraction due to the backward step, although the backward step is also known to exhibit similar contraction under suitable regularity assumptions [CCS+22].
Proposition 2.4.
Let $q\geq 2$ and $\mathbb{E}_{\pi^{X}}[|\cdot|^{4}]<\infty$. Then, the law $\pi_{k}^{X}$ of $X_{k}$ returned by the $\mathsf{Proximal\ sampler}$ satisfies $\chi^{q}(\pi_{k+1}^{X}\mathbin{\|}\pi^{X})\leq\eta_{\chi^{q}}(\pi^{X},h)\,\chi^{q}(\pi_{k}^{X}\mathbin{\|}\pi^{X})$. Given $\varepsilon\in(0,1)$, we can ensure that $\chi^{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
$N\lesssim qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{\chi^{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})}{\varepsilon}\quad\text{iterations\,.}$
Moreover, $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
$N\lesssim q\,\bigl(1\vee h^{-1}C_{\mathsf{PI}}(\pi^{X})\bigr)\log\frac{M_{q}}{\varepsilon}\quad\text{iterations\,.}$
We adapt this result to the setting of logconcave probability measures.
Lemma 2.5.
For any $\varepsilon\in(0,1)$, $q\geq 2$, and logconcave probability measure $\pi^{X}$ over $\mathbb{R}^{d}$, we can ensure $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
$N\lesssim q\,\bigl(1\vee h^{-1}\lVert\operatorname{cov}\pi^{X}\rVert\log d\bigr)\log\frac{M_{q}}{\varepsilon}\quad\text{iterations\,.}$
Proof.
Any logconcave probability measures satisfy ($\mathsf{PI}$) (and thus $\mathbb{E}_{\pi^{X}}[|\cdot|^{4}]<\infty$). The claim then follows from the proposition above with $C_{\mathsf{PI}}(\pi^{X})\lesssim\lVert\operatorname{cov}\pi^{X}\rVert\log d$ for logconcave probability measures [KLA23]. ∎
2.2 Per-step analysis
We now focus on the case of $\pi^{X}\propto\mathds{1}_{\mathcal{K}}$, assuming $B_{1}(0)\subset\mathcal{K}$ by translation. Hereafter, we assume that the starting distribution $\pi_{0}^{X}$ is $M_{q}$-warm with respect to $\pi^{X}$. By the DPI, the same holds between $\pi_{1}^{Y}=\pi_{0}^{X}*\gamma_{h}$ and $\pi^{Y}=\pi^{X}*\gamma_{h}$ (i.e., $\lVert\mathrm{d}\pi_{1}^{Y}/\mathrm{d}\pi^{Y}\rVert_{L^{q}(\pi^{Y})}<\infty$), and this property is preserved throughout subsequent iterations (i.e., $\mathcal{R}_{q}(\pi_{k}^{X}\mathbin{\|}\pi^{X})\leq\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$). For brevity, we will use the $L^{q}$-warmness, $\chi^{q}$-warmness, or $\mathcal{R}_{q}$-warmness interchangeably, as justified by (1.2).
2.2.1 High-level idea of analysis
Let us first analyze $\mathsf{PS}_{\textup{unif}}$ as a one-attempt algorithm that may declare failure if the inner rejection sampler exceeds the threshold $\tau$. Fix an iteration horizon $N\in\mathbb{N}$, step size $h>0$, and inner cap $\tau\in\mathbb{N}$.
Let $S_{i}$ denote the survival event that failure has not occurred up to and including step $i$, with $S_{0}=\Omega$ and full success event $S:=S_{N}$. Let $C_{i}\geq 0$ be the number of membership queries used by the backward step at step $i$. The total query cost is then $C:=\sum_{i=1}^{N}\mathds{1}[S_{i-1}]\,C_{i}$. Note that
$\mathbb{E}C=\sum_{i=1}^{N}\mathbb{E}\bigl[\mathds{1}[S_{i-1}]\,C_{i}\bigr]\leq\sum_{i=1}^{N}\mathbb{E}C_{i}\,$,
so it suffices to bound $\mathbb{E}C_{i}$. Similarly, if $F_{i}$ denotes the event of failure at step $i$, and $q_{i}$ denotes the cap-hit probability at step $i$, then
$\mathbb{P}(F_{i})=\mathbb{E}\bigl[\mathds{1}[S_{i-1}]\,q_{i}\bigr]\leq\mathbb{E}q_{i},\qquad\mathbb{P}(S^{c})\leq\sum_{i=1}^{N}\mathbb{E}q_{i}(Y_{i})\,.$
These inequalities allow us to carry out the main per-step analysis under the clean unconditional intermediate laws (e.g., $Y_{i}\sim\pi_{i}^{Y}$), without explicitly tracking the conditional distributions at each step.
(1) Bias from failure.
It may not be immediately clear whether the transition kernel of $\mathsf{PS}$ is preserved under the algorithmic modification of imposing a retry threshold $\tau$. For completeness, we verify that the backward step is indeed unchanged by this modification.
Lemma 2.6.
Upon success of the rejection sampling in the backward step, the accepted sample is distributed according to $\pi^{X|Y=y}$.
Proof.
Consider the proposal $\mu_{y}$ with success probability $p_{y}>0$ such that $X_{i}\sim\mu_{y}$ for each retry $i\in[\tau]$ satisfies $\operatorname{law}(X_{i}\,|\,\text{accepted})=\pi^{X|Y=y}$. Let $X$ be the random sample given by the backward step. For any event $A$ in $\mathbb{R}^{d}$, we have
$\mathbb{P}(X\in A\,|\,\text{accepted})$ $=\frac{\mathbb{P}(X\in A,\,\text{success})}{\mathbb{P}(\text{success})}$
[…]
$=\frac{\sum_{i=1}^{\tau}(1-p_{y})^{i-1}\,\mathbb{P}(X_{i}\in A,\,X_{i}\ \text{accepted})}{\sum_{i=1}^{\tau}(1-p_{y})^{i-1}p_{y}}=\frac{\mathbb{P}(X_{1}\in A,\,X_{1}\ \text{accepted})}{p_{y}}$
$=\mathbb{P}(X_{1}\in A\,|\,\text{accepted})=\pi^{X|Y}(A)\,$,
which completes the proof. ∎
However, the law of output conditioned on success introduces a small bias, which is bounded roughly by the failure probability of the algorithm.
Lemma 2.7 (Bias from success, [KVZ26]).
Let $S$ be the event that the failure is not declared over the required $N$ iterations, and suppose $\mathbb{P}(S)\geq 1-\eta$. Let $\pi_{N}^{X}$ be the law of the uncapped $\mathsf{PS}_{\textup{unif}}$ output after $N$ iterations, and $\hat{\pi}_{N}^{X}$ be the law of $\mathsf{PS}_{\textup{unif}}$ conditioned on $S$ (i.e., $\hat{\pi}_{N}^{X}:=\operatorname{law}(X_{N}\mid S)$). Then,
$\frac{\mathrm{d}\hat{\pi}_{N}^{X}}{\mathrm{d}\pi_{N}^{X}}$ $\leq\frac{1}{1-\eta}\quad\textup{so}\quad\mathcal{R}_{\infty}(\hat{\pi}_{N}^{X}\mathbin{\|}\pi_{N}^{X})\leq\log\frac{1}{1-\eta}\,$,
$\mathcal{R}_{q}(\hat{\pi}_{N}^{X}\mathbin{\|}\pi^{X})$ $\leq\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})+\frac{q}{q-1}\log\frac{1}{1-\eta}\qquad\text{for }q>1\,.$
(2) Restart-until-success.
A practical restart-until-success strategy repeats independent attempts until $S$ occurs. It outputs a sample with (conditional) law $\hat{\pi}_{N}^{X}$ and terminates almost surely whenever $\mathbb{P}(S)>0$. Moreover, if $\mathbb{P}(S)\geq 1-\eta$, then the expected number of attempts is at most $(1-\eta)^{-1}$ and
$\mathbb{E}[C\mid S]\leq\frac{\mathbb{E}C}{\mathbb{P}(S)}\leq\frac{\mathbb{E}C}{1-\eta}\,,\qquad\mathbb{E}C_{\text{restart}}=\frac{\mathbb{E}C}{\mathbb{P}(S)}\leq\frac{\mathbb{E}C}{1-\eta}\,.$
It therefore suffices to bound the one-attempt success probability $\mathbb{P}(S)$ and the unconditional expected cost $\mathbb{E}C$.
2.2.2 Per-step analysis
The analysis follows [KV25a] closely, which establishes the query complexity of $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{\widetilde{O}(1)}\to\mathcal{R}_{2}$. The only difference lies in the required number $N$ of iterations, namely the number needed for contraction in $\mathcal{R}_{q}$. Thus, we recall the relevant results from §2.1 of [KV25a] and adapt them minimally for our setting. With the parameter choices $Z=\frac{16NM_{2}}{\eta}$, $c=\frac{\log\log Z}{2\log Z}$, $\tau=Z^{2}\log^{3}Z$, and $h=\frac{c}{d^{2}}$, the failure probability satisfies $\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq\frac{\eta}{N}$. The per-step complexity is bounded as follows: setting $q=1+\alpha$ with $\alpha\geq 1\vee\log\tau$, it follows that $\mathbb{E}_{\pi_{k}^{Y}}[\frac{1}{\ell}\wedge\tau]\leq 5M_{q}\log^{4}Z$.
We now prove the order-preserving complexity guarantees of $\mathsf{PS}_{\textup{unif}}$ under an $O(1)$-warm start and sufficient condition on $q$.
Proposition 2.8.
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ with $B_{1}\subset\mathcal{K}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$, given access to a membership oracle for $\mathcal{K}$. Given $\varepsilon>0$ , $\eta\in(0,1/2)$, $q\geq 2$, and an initial distribution $\pi_{0}$ with $M_{q}=\lVert\nicefrac{{\mathrm{d}\pi_{0}}}{{\mathrm{d}\pi}}\rVert_{L^{q}(\pi)}$, initialize $\mathsf{PS}_{\textup{unif}}$ with $\pi_{0}$ and iterate it $N=\widetilde{O}(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log\frac{1}{\eta})$ times with $h=(2d^{2}\log Z)^{-1}$ and $\tau=Z^{2}\log^{3}Z$ for $Z:=\frac{16NM_{2}}{\eta}$. If $q\geq 1+\log\tau$ and $M_{q}\leq 10$, then
  • The sufficient condition on $q$ is satisfied when $q\geq 2\vee O(\log\frac{N}{\eta})$.
  • With probability at least $1-\eta$, $\mathsf{PS}_{\textup{unif}}$ iterates $N$ times without failure. Conditioned on this success, the output $Z\sim\nu$ satisfies $\mathcal{R}_{q}(\nu\mathbin{\|}\pi)\leq\varepsilon+2\log\frac{1}{1-\eta}$, with $\widetilde{O}(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log^{5}\frac{1}{\eta})$ membership queries in expectation.
Proof.
The uncapped $\mathsf{PS}_{\textup{unif}}$ can ensure $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after iterating
$N\gtrsim q\,\bigl(1\vee h^{-1}C_{\mathsf{PI}}(\pi)\bigr)\log\frac{M_{q}}{\varepsilon}\,.$
Under the choice of $h=cd^{-2}=\frac{\log\log Z}{2d^{2}\log Z}$ with $Z\asymp\frac{NM_{2}}{\eta}$, the required number $N$ of iterations must satisfy an inequality of the form
$N\gtrsim A\log(BN)\,$,
which can be fulfilled if $N\gtrsim A\log(AB)$. Therefore, the required number of iterations is of order
$N=\widetilde{O}\bigl(qd^{2}C_{\mathsf{PI}}(\pi)\log\frac{M_{q}}{\varepsilon}\log\frac{M_{q}}{\eta}\bigr)=\widetilde{O}\bigl(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log\frac{1}{\eta}\bigr)\,.$
where $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and the last equality follows from the assumption $M_{q}\leq 10$.
We now find a sufficient condition for ensuring $q\geq 1+\log\tau$. Observe that when $q\geq 2$,
$1+\log\tau=1+\frac{q}{q-1}\log(Z\log^{4}Z)\lesssim\log Z\lesssim\log\frac{N}{\eta}\,.$
Hence, when $q\geq 2\vee O(\log\frac{N}{\eta})$, then expected total query complexity of $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ from an $O(1)$-warm start is
$\widetilde{O}\bigl(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log\frac{1}{\eta}\times\log^{4}Z\bigr)=\widetilde{O}\bigl(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log^{5}\frac{1}{\eta}\bigr)\,.$
Lastly, the success probability satisfies $\mathbb{P}(S)\geq\eta$ by construction, and by Lemma 2.7, the output law $\hat{\pi}_{N}^{X}$ conditioned on success satisfies
$\mathcal{R}_{q}(\hat{\pi}_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon+2\log\frac{1}{1-\eta}\,$,
which completes the proof. ∎
Using this proposition, we prove Theorem 1.1.
Proof of Theorem 1.1.
Take $\eta=\varepsilon<0.01$, and restart the algorithm upon failure. Note that the expected number of restarts is at most $0.02$. To invoke the proposition above, $q$ should satisfy $q\geq 2\vee O(\log N)$ for $N=\widetilde{O}(qd^{2}\Lambda\log^{2}\frac{1}{\varepsilon})$. To ensure
$q\geq\log\widetilde{O}\bigl(qd^{2}\Lambda\log^{2}\frac{1}{\varepsilon}\bigr)(\geq\log N)\,$,
let us solve the inequality of the form $q\gtrsim A\log qB$. A sufficient condition for this is simply $q\gtrsim A\log AB=\widetilde{O}(A\log B)$, so it suffices to enforce
$q\geq 2\vee\widetilde{O}\bigl(\log d^{2}\Lambda+\log\log\frac{1}{\varepsilon}\bigr)\,.$ (2.3)
Hence, by Proposition 2.8, the restart version of $\mathsf{PS}_{\textup{unif}}$ uses $\widetilde{O}(qd^{2}\Lambda\log^{6}\frac{1}{\varepsilon})$ total queries in expectation, and its output is distributed as $\hat{\pi}_{N}^{X}$ satisfying $\mathcal{R}_{q}(\hat{\pi}_{N}^{X}\mathbin{\|}\pi^{X})\leq 3\varepsilon$. Retaking $\varepsilon\leftarrow\varepsilon/3$, we complete the proof. ∎
3 Hypercontractivity under the simultaneous heat flow
In this section, we establish hypercontractivity under simultaneous heat flow, which will serve as a stepping stone for relaying a Rényi warmness in annealing in the next section.
3.1 Adjoint of the heat semigroups and the proximal sampler
We first recap the Markov semigroup theory for a gentle introduction to heat semigroup and its adjoint. Interested readers may want to refer to [VAN14, BGL14, CHE25] for further details on the semigroup theory. Meanwhile, we take detour to $\mathsf{PS}$, since it sets up a context in which the introduction of heat semigroups and its adjoint is natural.
Markov semigroup theory.
Given a (time-homogeneous) Markov process $(X_{t})_{t\geq 0}$ on a domain $E$, its associated Markov semigroup $(P_{t})_{t\geq 0}$ is defined as a family of functional operators such that
$(P_{t}f)(x):=\mathbb{E}[f(X_{t})\mid X_{0}=x]\quad\text{for every bounded measurable function }f:E\to\mathbb{R}\,.$
One could readily check that $P_{t}$ is a linear operator with $P_{0}=\mathrm{id}$, $P_{t+s}=P_{t}\circ P_{s}$ for any $t,s\geq 0$ (semigroup property), $P_{t}\mathds{1}=\mathds{1}$ (mass conservation), and $P_{t}f\geq 0$ for any $f\geq 0$ (positivity preserving). Below, we omit $E$ in the integral sign.
A measure $\mu$ is said to be invariant (or stationary) for $P_{t}$ if $\int P_{t}f\,\mathrm{d}\mu=\int f\,\mathrm{d}\mu$ for every $t\geq 0$. In addition, it is said to be reversible for $P_{t}$ if $\int P_{t}f\,g\,\mathrm{d}\mu=\int f\,P_{t}g\,\mathrm{d}\mu$ for any $f,g\in L^{2}(\mu)$ and $t\geq 0$. For any $p\geq 1$, it follows from Jensen’s inequality that $P_{t}$ is a contraction operator on the bounded functions in $L^{p}(\mu)$ (i.e., $\lVert P_{t}f\rVert_{L^{p}(\mu)}\leq\lVert f\rVert_{L^{p}(\mu)}$). This furnishes the extension of $P_{t}$ to $L^{p}(\mu)$ by a density argument.
The proximal sampler.
As examined in [CCS+22] (with further details in [CHE25]), the forward step (i.e., $y_{k+1}\sim\pi^{Y|X=x_{k}}=\mathcal{N}(x_{k},hI_{d})$) can be viewed as simulating Brownian motion for time $h$ started at $x_{k}$; when $\mathrm{d}X_{t}=\mathrm{d}B_{t}$ with $X_{0}=x_{k}$, we have $X_{h}\sim\operatorname{law}y_{k+1}$. The associated Markov semigroup of Brownian motion is often called the heat semigroup, which admits the formula of $P_{t}f=f*\gamma_{t}=:f_{t}$. Based on the classical heat equation $\partial_{t}P_{t}f=\frac{1}{2}\,\Delta P_{t}f$, rich theory for the heat semigroup facilitates one to understand the forward step in a better way (e.g., contraction under simultaneous heat flow).
This prompts a natural question whether one could also introduce a semigroup structure to the backward step. It turns out that the backward step can be interpreted as an SDE $(Z_{t}^{\leftarrow})_{t\geq 0}$, given by
$\mathrm{d}Z_{t}^{\leftarrow}=\nabla\log(\pi^{X}*\gamma_{h-t})\,(Z_{t}^{\leftarrow})\,\mathrm{d}t+\mathrm{d}\bar{B}_{t}\qquad\text{for another Brownian motion }\bar{B}_{t}\,$,
so that if $Z_{0}^{\leftarrow}\sim\pi^{Y}$, then $Z_{h}^{\leftarrow}\sim\pi^{X}$. Though it is non-trivial to justify, one could guess that $Z_{h}^{\leftarrow}\sim\pi^{X|Y=y}$ if $Z_{0}^{\leftarrow}=y$. Hence, one may be able to introduce a semigroup structure to the backward step via, for any suitable test function $f$,
$\overline{Q}_{h}f(y):=\mathbb{E}[f(Z_{h}^{\leftarrow})\mid Z_{0}^{\leftarrow}=y]=\mathbb{E}_{\pi^{X|Y=y}}f\,.$
As seen shortly, Klartag and Putterman [KP23] defined the same object as the $L^{2}(\pi)$-adjoint of the heat semigroups, providing solid ground where similar rigorous theory can be developed for the backward step.
$L^{2}$-adjoint of heat semigroups.
Given a probability measure $\pi$, [KP23] viewed the heat semigroup $P_{t}$ as a contraction operator from $L^{2}(\pi_{t})$ to $L^{2}(\pi)$:
$\lVert P_{t}f\rVert_{L^{2}(\pi)}^{2}=\int(P_{t}f)^{2}\,\pi\leq\int P_{t}(f^{2})\,\pi=\int f^{2}\,P_{t}\pi=\lVert f\rVert_{L^{2}(\pi_{t})}^{2}\,$,
where the equality in the middle follows simply from reversibility of the heat semigroups. This perspective led them to define the adjoint operator $Q_{t}:L^{2}(\pi)\to L^{2}(\pi_{t})$ as $Q_{0}=\mathrm{id}$ and
$Q_{t}f=\frac{P_{t}(f\pi)}{P_{t}\pi}=\frac{P_{t}(f\pi)}{\pi_{t}}\quad\text{for }t>0\,.$
This formula can be naturally derived, though not mentioned in the original manuscript, from the definition of adjoint: for any $f\in L^{2}(\pi)$ and $g\in L^{2}(\pi_{t})$,
$\langle f,P_{t}g\rangle_{L^{2}(\pi)}=\langle Q_{t}f,g\rangle_{L^{2}(\pi_{t})}\,.$
For $\mathrm{d}Z_{t}=\mathrm{d}B_{t}$, one could rewrite the LHS as follows:
$\langle f,P_{t}g\rangle_{L^{2}(\pi)}=\int f\,P_{t}g\,\mathrm{d}\pi=\mathbb{E}\bigl[f(Z_{0})\,\mathbb{E}[g(Z_{t})|Z_{0}]\bigr]=\mathbb{E}[f(Z_{0})\,g(Z_{t})]=\mathbb{E}_{Z_{t}\sim\pi_{t}}\bigl[g(Z_{t})\,\mathbb{E}[f(Z_{0})|Z_{t}]\bigr]\,.$
Equating this with $\langle Q_{t}f,g\rangle_{L^{2}(\pi_{t})}$, we can deduce that
$Q_{t}f(y)=\mathbb{E}[f(Z_{0})|Z_{t}=y]\,.$
Since $\operatorname{law}(Z_{0}|Z_{t}=y)=\pi^{X|Y=y}$ with $\pi^{X}=\pi$,
$Q_{t}f(y)=\int f\,\mathrm{d}\pi^{X|Y=y}=\frac{\int f(x)\,\pi^{X}(x)\,\gamma_{t}(y-x)\,\mathrm{d}x}{\int\pi^{X}(x)\,\gamma_{t}(y-x)\,\mathrm{d}x}=\frac{(f\pi)*\gamma_{t}}{\pi*\gamma_{t}}=\frac{P_{t}(f\pi)}{\pi_{t}}\,.$
We note that this heat adjoint could be viewed as the backward step of $\mathsf{PS}$: $\overline{Q}_{t}f=Q_{t}f$. Also, it follows from the above formula that the adjoint operator satisfies mass preservation ($Q_{t}\mathds{1}=\mathds{1}$) and positivity preserving ($Q_{t}f\geq 0$ if $f\geq 0$). Moreover, it is a contraction operator; for $f\in L^{2}(\pi)\backslash\{0\}$, by Cauchy-Schwarz
$(Q_{t}f)^{2}\,\pi_{t}=\frac{\bigl(P_{t}(f\pi)\bigr)^{2}}{\pi_{t}}\leq\frac{P_{t}(f^{2}\pi)\,P_{t}\pi}{\pi_{t}}=P_{t}(f^{2}\pi)\,.$
and integration of both sides yields $\lVert Q_{t}f\rVert_{L^{2}(\pi_{t})}^{2}\leq\lVert f\rVert_{L^{2}(\pi)}^{2}$. We refer readers to [KV25c, §B.2].
3.2 Hypercontractivity of an adjoint of the heat semigroups
We now show something stronger than mere contraction of $Q_{t}$ from $L^{2}(\pi)$ to $L^{2}(\pi_{t})$, called hypercontractivity. To begin with, we recall hypercontractivity of Markov semigroups under ($\mathsf{LSI}$):
Proposition 3.1 (Gross).
For $C>0$ and $p\in(1,\infty)$, let $q(t)=1+(p-1)\,e^{2t/C}$ for $t\geq 0$, and $P_{t}$ be a reversible Markov semigroup with stationary measure $\nu$. Then, $\nu$ satisfies ($\mathsf{LSI}$) with parameter $C$ if and only if
$\lVert P_{t}f\rVert_{L^{q(t)}(\nu)}\leq\lVert f\rVert_{L^{p}(\nu)}\quad\text{for all }p\in(1,\infty),t\geq 0,\text{ and }f\in L^{p}(\nu)\,.$
It means that when a stationary measure satisfies ($\mathsf{LSI}$), the semigroup $P_{t}$ not only contracts in any $L^{p}$-spaces but also improves integrability (recall $L^{q}\subset L^{p}$ for $p<q$ due to monotonicity of $\lVert\cdot\rVert_{L^{p}}$ in $p$). We now prove Theorem 1.2, hypercontractivity under the simultaneous heat flow with different $q(t)=1+(p-1)\,(1+\nicefrac{{t}}{{C_{\mathsf{LSI}}(\nu)}})$;
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f\rVert_{L^{p}(\nu)}\quad\text{for all }p\in(1,\infty),t\geq 0,\text{ and }f\in L^{p}(\nu)\,.$
Note that the reference measure $\nu_{t}$ also evolves over time $t$.
Proof of Theorem 1.2.
We first prove the claim for $f\in C_{c,+}^{\infty}(\mathbb{R}^{d}):=\{f+c:f\in C_{c}^{\infty}(\mathbb{R}^{d}),f\geq 0,c>0\}$. Let $\bar{f}$ be the normalization of $f$ with respect to $\nu$ (i.e., $\bar{f}=f/\int f\,\mathrm{d}\nu$).
Denoting a new probability measure $\mu$ by $\mathrm{d}\mu=\bar{f}\,\mathrm{d}\nu$, we define $\bar{f}_{t}:=Q_{t}\bar{f}=\mathrm{d}\mu_{t}/\mathrm{d}\nu_{t}$ and
$\mathcal{F}(t,q):=\int\bigl(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr)^{q}\,\mathrm{d}\nu_{t}=\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\,,\qquad\mathcal{G}(t):=\frac{1}{q(t)}\,\log\mathcal{F}\bigl(t,q(t)\bigr)\,.$
It then suffices to show that $\partial_{t}\mathcal{G}\leq 0$ for a suitable choice of $q(t)$. Note that
$\partial_{t}\mathcal{G}=-\frac{q^{\prime}(t)}{q^{2}(t)}\,\log\mathcal{F}+\frac{\partial_{t}\mathcal{F}+q^{\prime}(t)\,\partial_{q}\mathcal{F}}{q(t)\,\mathcal{F}}\,.$
With assumptions of $q(t),q^{\prime}(t)>0$ for a moment, $\partial_{t}\mathcal{G}\leq 0$ is equivalent to
$\frac{1}{q^{\prime}(t)}\,\partial_{t}\mathcal{F}+\partial_{q}\mathcal{F}-\frac{1}{q(t)}\,\mathcal{F}\log\mathcal{F}\leq 0\,.$
We now compute $\partial_{t}\mathcal{F}$ and $\partial_{q}\mathcal{F}$. By general de Bruijn’s identity in Proposition 2.2 (note that $\mathbb{E}_{\nu}[|\cdot|^{4}]<\infty$ since ($\mathsf{LSI}$) implies ($\mathsf{PI}$)),
$\partial_{t}\mathcal{F}=-\frac{q\,(q-1)}{2}\,\mathbb{E}_{\nu_{t}}\bigl[\bigl(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr)^{q-2}\,\bigl|\nabla\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr|^{2}\bigr]=-\frac{q\,(q-1)}{2}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]\,.$
As for $\partial_{q}\mathcal{F}$, we have
$\partial_{q}\mathcal{F}$ $=\partial_{q}\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\underset{(i)}{=}\int\bar{f}_{t}^{q}\log\bar{f}_{t}\,\mathrm{d}\nu_{t}=\frac{1}{q}\,\Bigl(\operatorname{Ent}_{\nu_{t}}(\bar{f}_{t}^{q})+\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\,\log\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\Bigr)$
$\leq\frac{1}{q}\,\Bigl(2C_{\mathsf{LSI}}(\nu_{t})\,\mathbb{E}_{\nu_{t}}[|\nabla(\bar{f}_{t}^{q/2})|^{2}]+\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\,\log\int\bar{f}_{t}^{q}\,\mathrm{d}\nu_{t}\Bigr)$
$=\frac{1}{q}\,\bigl(\frac{q^{2}C_{\mathsf{LSI}}(\nu_{t})}{2}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]+\mathcal{F}\log\mathcal{F}\bigr)\,.$
Differentiation under the integral sign in $(i)$ can be justified as follows: $\partial_{q}\bar{f}_{t}^{q}=(Q_{t}\bar{f})^{q}\log Q_{t}\bar{f}$ is continuous in $q$, and we can take a small $\delta>0$ such that for any $q^{\prime}\in(q-\delta,q+\delta)$ with $q-\delta>1$,
$\bar{f}_{t}^{q^{\prime}}\log\bar{f}_{t}=(\bar{f}_{t}\,\mathds{1}[\bar{f}_{t}\geq 1]+\bar{f}_{t}\,\mathds{1}[\bar{f}_{t}<1])^{q^{\prime}}\log\bar{f}_{t}\leq 2^{q+\delta-1}\,(\bar{f}_{t}^{q+\delta}+\bar{f}_{t}^{q-\delta})\log\bar{f}_{t}\,.$
Hence, it amounts to showing integrability of the RHS. To this end, we first show that $\bar{f}_{t}\in L^{r}(\nu_{t})$ for any $r>1$. By Jensen’s inequality,
$\lVert\bar{f}_{t}\rVert_{L^{r}(\nu_{t})}^{r}=\mathbb{E}_{\nu_{t}}[(Q_{t}\bar{f})^{r}]\leq\mathbb{E}_{\nu_{t}}[Q_{t}(\bar{f}^{r})]=\mathbb{E}_{\nu}[\bar{f}^{r}]=\lVert\bar{f}\rVert_{L^{r}(\nu)}^{r}\,$,
and $\bar{f}\in C_{c,+}^{\infty}(\mathbb{R}^{d})$ ensures that $\lVert\bar{f}\rVert_{L^{r}(\nu)}<\infty$. Thus, integrability of the RHS follows from $L^{s}\subset L^{r}\log L$ for any $r<s$.
Combining the bounds on $\partial_{t}\mathcal{F}$ and $\partial_{q}\mathcal{F}$,
$\frac{1}{q^{\prime}(t)}\,\partial_{t}\mathcal{F}+\partial_{q}\mathcal{F}-\frac{1}{q(t)}\,\mathcal{F}\log\mathcal{F}$
$\leq-\frac{q\,(q-1)}{2\,q^{\prime}(t)}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]+\frac{1}{q}\,\bigl(\frac{q^{2}C_{\mathsf{LSI}}(\nu_{t})}{2}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]+\mathcal{F}\log\mathcal{F}\bigr)-\frac{1}{q(t)}\,\mathcal{F}\log\mathcal{F}$
$=\bigl(C_{\mathsf{LSI}}(\nu_{t})-\frac{q-1}{q^{\prime}(t)}\bigr)\times\frac{q}{2}\,\mathbb{E}_{\nu_{t}}[\bar{f}_{t}^{q-2}\,|\nabla\bar{f}_{t}|^{2}]\,.$
Hence, it suffices to enforce
$\frac{(q-1)^{\prime}}{q-1}\leq\frac{1}{C_{\mathsf{LSI}}(\nu_{t})}\,.$
As $C_{\mathsf{LSI}}(\nu_{t})\leq C_{\mathsf{LSI}}(\nu)+t$, we can guarantee this by solving
$\frac{(q-1)^{\prime}}{q-1}\leq\frac{1}{C_{\mathsf{LSI}}(\nu)+t}\,.$
Integrating both sides over $[0,t]$, and setting $q(0)=p$,
$\log\frac{q(t)-1}{p-1}\leq\log\frac{C_{\mathsf{LSI}}(\nu)+t}{C_{\mathsf{LSI}}(\nu)}\,$,
and thus we can simply take
$\frac{q(t)-1}{p-1}=1+\frac{t}{C_{\mathsf{LSI}}(\nu)}\,$,
which clearly satisfies $q,q^{\prime}\geq 0$. Thus, $\lVert\bar{f}_{t}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert\bar{f}\rVert_{L^{p}(\nu)}$ (so $\lVert f_{t}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f\rVert_{L^{p}(\nu)}$), and hypercontractivity holds on $C_{c,+}^{\infty}(\mathbb{R}^{d})$.
Every $f\in C_{c,+}^{\infty}(\mathbb{R}^{d})$ can be written as $f=f_{\delta}=g+\delta$ for some non-negative $g\in C_{c}^{\infty}(\mathbb{R}^{d})$ and $\delta>0$. We then apply the dominated convergence theorem to $\lVert Q_{t}f_{\delta}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f_{\delta}\rVert_{L^{p}(\nu)}$ as $\delta\to 0$, extending the main claim from $C_{c,+}^{\infty}(\mathbb{R}^{d})$ to $C_{c,\geq 0}^{\infty}(\mathbb{R}^{d}):=C_{c}^{\infty}(\mathbb{R}^{d})\cap\{f\geq 0\}$.
Now, for any non-negative $f\in L^{p}(\nu)$, we can take a sequence $\{f_{n}\}$ of functions in $C_{c,\geq 0}^{\infty}(\mathbb{R}^{d})$ such that $f_{n}\to f$ in $L^{p}$. Moreover, we can take a subsequence $\{n_{k}\}_{k}$ such that $f_{n_{k}}\to f$ almost everywhere. Since $Q_{t}f_{n}\geq 0$ due to positivity-preserving, Fatou’s lemma implies that
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}^{q(t)}=\int\liminf_{k\to\infty}(Q_{t}f_{n_{k}})^{q(t)}\,\mathrm{d}\nu_{t}\leq\liminf_{k\to\infty}\int(Q_{t}f_{n_{k}})^{q(t)}\,\mathrm{d}\nu_{t}\,$,
and thus,
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}\leq\liminf_{k\to\infty}\,\lVert Q_{t}f_{n_{k}}\rVert_{L^{q(t)}(\nu_{t})}\leq\liminf_{k\to\infty}\,\lVert f_{n_{k}}\rVert_{L^{p}(\nu)}=\lVert f\rVert_{L^{p}(\nu)}\,.$
Lastly, for any $f\in L^{p}(\nu)$, note that $|f|=f_{+}+f_{-}$ is also in $L^{p}(\nu)$. Thus,
$\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}=\lVert|Q_{t}f|\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert Q_{t}(|f|)\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert|f|\rVert_{L^{p}(\nu)}=\lVert f\rVert_{L^{p}(\nu)}\,$,
which finishes the proof. ∎
For $f:=\mathrm{d}\mu/\mathrm{d}\nu\in L^{p}(\nu)$ (which is equivalent to $\mathcal{R}_{p}(\mu\mathbin{\|}\nu)<\infty$), hypercontractivity of $Q_{t}$ implies that
$\Bigl\|\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\Bigr\|_{L^{q(t)}(\nu_{t})}=\lVert Q_{t}f\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert f\rVert_{L^{p}(\nu)}=\Bigl\|\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr\|_{L^{p}(\nu)}\,.$
Then, $\mathcal{R}_{q(t)}(\mu_{t}\mathbin{\|}\nu_{t})\leq\mathcal{R}_{p}(\mu\mathbin{\|}\nu)$ follows from
$\mathcal{R}_{p}(\mu\mathbin{\|}\nu)=\frac{p}{p-1}\log\,\Bigl\|\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr\|_{L^{p}(\nu)}\geq\frac{p}{p-1}\,\log\,\Bigl\|\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\Bigr\|_{L^{q(t)}(\nu_{t})}\geq\frac{q(t)}{q(t)-1}\,\log\,\Bigl\|\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\Bigr\|_{L^{q(t)}(\nu_{t})}\,.$
This looks similar to the DPI for Rényi divergence, but we remark that the Rényi order of the RHS is larger that that of the LHS.
Improved convergence rate of PS under LSI.
Just as in the SDPI under ($\mathsf{PI}$), the SDPI under ($\mathsf{LSI}$) was established in [CCS+22, KO25]: for $q>1$ and two probability measures $\mu,\nu$ with $\mu\ll\nu$,
$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{LSI}}(\nu))^{1/q}}\,.$ (3.1)
Combining this contraction with hypercontractivity, we can derive a stronger mixing result of $\mathsf{PS}$: when $\pi_{k}$ denoting the law of the $k$-th sample returned by $\mathsf{PS}$ with initial $\pi_{0}$ and target $\pi$, for any given $\varepsilon>0$ and $p,q>1$, we have $\mathcal{R}_{q}(\pi_{N}\mathbin{\|}\pi)\leq\varepsilon$ after $N\lesssim h^{-1}C_{\mathsf{LSI}}(\pi)\log\frac{q\mathcal{R}_{2\wedge p}(\pi_{0}\mathbin{\|}\pi)}{(p-1)\,\varepsilon}$ iterations.
Proof of Corollary 1.3.
We first iterate $\mathsf{PS}$ $N_{0}\asymp(2\wedge p)\,h^{-1}C_{\mathsf{LSI}}(\pi)\log\frac{\mathcal{R}_{(2\wedge p)}(\pi_{0}\mathbin{\|}\pi)}{\varepsilon}$ times to bring $\pi_{N_{0}}$ within $\varepsilon$-distance to $\pi$ in $\mathcal{R}_{2\wedge p}$. Next, we iterate it $N_{1}\asymp h^{-1}C_{\mathsf{LSI}}(\pi)\log_{2}\frac{q-1}{1\wedge(p-1)}$ more times, where we invoke DPI and hypercontractivity each iteration as follows: for $i\leq N_{1}$ and $\pi_{i}^{X}=\pi_{i}$,
$\mathcal{R}_{q_{i+1}}(\pi_{N_{0}+i+1}\mathbin{\|}\pi)\leq\mathcal{R}_{q_{i+1}}(\pi_{N_{0}+i+1}^{Y}\mathbin{\|}\pi^{Y})\leq\mathcal{R}_{q_{i}}(\pi_{N_{0}+i}\mathbin{\|}\pi)\,$,
where $(q_{i+1}-1)/(q_{i}-1)=1+h/C_{\mathsf{LSI}}(\pi)$. Then,
$\frac{q_{N_{1}}-1}{1\wedge(p-1)}=\frac{q_{N_{1}}-1}{q_{0}-1}=\prod_{i=1}^{N_{1}}\frac{q_{i}-1}{q_{i-1}-1}=\bigl(1+\frac{h}{C_{\mathsf{LSI}}(\pi)}\bigr)^{N_{1}}\geq\frac{q-1}{1\wedge(p-1)}\,.$
Therefore, $\mathsf{PS}$ suffices to iterate $N=N_{0}+N_{1}\lesssim h^{-1}C_{\mathsf{LSI}}(\pi)\log\frac{q\mathcal{R}_{2\wedge p}(\pi_{0}\mathbin{\|}\pi)}{(1\wedge(p-1))\,\varepsilon}$ times to attain the desired guarantee. ∎
As a simple corollary, this result leads to a mixing time of $\mathsf{PS}$ in the well-conditioned setting, which improves the linear dependence of $\mathsf{PS}$ on $q$ [CCS+22].
Proposition 3.2 (Zeroth-order sampling in well-conditioned setting).
Let $\pi\propto e^{-V}$ with $\alpha$-strongly convex and $\beta$-smooth $V$ over $\mathbb{R}^{d}$, given access to an evaluation and proximal oracle for $V$. Let $\pi_{k}$ denote the law of the $k$-th sample returned by $\mathsf{PS}$. Given $q>1$ and $\varepsilon>0$, we can attain $\mathcal{R}_{q}(\pi_{N}\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(\kappa d\log\frac{q}{\varepsilon})$ queries to both oracles, where $\kappa=\beta/\alpha$.
Proof.
One can find an initial distribution $\pi_{0}$ such that $\mathcal{R}_{\infty}(\pi_{0}\mathbin{\|}\pi)\leq d\log\kappa$ (see [LST21, Lemma 18]). Using $\mathcal{R}_{2}\leq\mathcal{R}_{\infty}$, $C_{\mathsf{LSI}}(\pi)\leq\alpha^{-1}$, and Corollary 1.3 with $q_{0}=2$, we can conclude that $\mathsf{PS}$ suffices to iterate $\widetilde{O}(\kappa d\log\frac{q}{\varepsilon})$ times. It is also known from [LST21] that the backward step can be implemented via rejection sampling for step size $h=\nicefrac{{1}}{{\beta d}}$, using $O(1)$ many evaluation queries (in expectation) and one proximal query per iteration. Combining these together finishes the proof. ∎
4 Order-preserving annealing via simpler warm-start generation
In this section, we streamline prior warm-start generation algorithms and their proofs. Above all, we interweave an overall analysis with hypercontractivity under the heat flow, eventually establishing the complexity of generating $\mathcal{R}_{q}$-warmness.
Simpler warm-start generation.
Let $\mathcal{K}$ be a convex body with $B_{1}(0)\subset\mathcal{K}$ (by translation), $\mathbb{E}_{\pi}[|\cdot|^{2}]\leq R^{2}$, and $\Lambda=\lVert\operatorname{cov}\pi\rVert$. We simplify prior annealing approaches in [KV25a], moving along a sequence $\{\mu_{i}\}_{i\in[m]}$ of distributions with $\mu_{i}=\gamma_{\sigma_{i}^{2}}|_{\mathcal{K}}=\pi\gamma_{\sigma_{i}^{2}}$.
This prior algorithm consists of the following several steps: (i) Preprocessing: Truncate $\mathcal{K}$ to the ball of radius $O(R)$ such that $\pi(\bar{\mathcal{K}})\geq 0.99$ where $\bar{\mathcal{K}}:=\mathcal{K}\cap B_{O(R)}(0)$, (ii) Initialization: Sample from $\mu_{1}=\gamma_{d^{-1}}|_{\bar{\mathcal{K}}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$, (iii) Annealing: While $\sigma_{i}^{2}\lesssim R^{2}$, use $\mathsf{PS}$ to sample from $\mu_{i+1}$ with initial distribution $\mu_{i}$, where $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,(1+\frac{\sigma_{i}}{q^{1/2}R})$ for $q=\operatorname{polylog}\nicefrac{{dR}}{{\varepsilon}}$. In particular, there are two modes: (iii-1) normal mode: when $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}$, iterate $\mathsf{PS}$ $\widetilde{O}(d^{2}\sigma_{i}^{2})$ times, (iii-2) faster mode: when $\sigma_{i}^{2}\gtrsim qR\Lambda^{1/2}$, iterate $\mathsf{PS}$ $\widetilde{O}(d^{2}R\Lambda^{1/2})$ times. (iv) Termination: Run $\mathsf{PS}$ to sample from $\pi$ with initial distribution $\gamma_{R^{2}}|_{\bar{\mathcal{K}}}$.
We show that the preprocessing step is redundant and also that the faster mode can be replaced by early-stopping: using $\mathsf{PS}_{\textup{Gauss}}$4.1) below,
  1. 1.
    Sample from $\mu_{1}=\pi\gamma_{d^{-1}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$, and set $\bar{\mu}_{1}:=\mu_{1}$.
  2. 2.
    While $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{Gauss}}$ such that $\mathcal{R}_{\widetilde{O}(1)}(\bar{\mu}_{i+1}\mathbin{\|}\mu_{i+1})=O(1)$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{\sigma_{i}}{R}\bigr)\,.$
We need two ingredients to obtain a final guarantee: (1) the query complexity of $\mathsf{PS}$ for truncated Gaussians ($\mathsf{PS}_{\textup{Gauss}}$) and (2) closeness of two annealing distributions. The former was already studied in previous work, but we will refine the analysis to obtain an order-boosting guarantee via hypercontractivity in §4.1. The latter was proven for $\chi^{2}$-divergence by a localization argument [CV18] and then was extended to $\mathcal{R}_{q}$ in [KV25a]. We show in §4.2 that this is in fact a simple consequence of ($\mathsf{LSI}$). Combining these together, we establish the query complexity of generating $\mathcal{R}_{q}$-warmness.
4.1 Proximal sampler for truncated Gaussian
We study $\mathsf{PS}$ for $\mu^{X}:=\pi\gamma_{\sigma^{2}}$, Gaussian distributions truncated to a convex body $\mathcal{K}$ containing $B_{1}(0)$. Its prototype was already studied in [KZ25, KV25a], and $\mathsf{PS}_{\textup{Gauss}}$ can be readily deduced from the general framework of $\mathsf{PS}$.
Starting with $x_{0}\sim\mu_{0}^{X}$, step size $h>0$, threshold $\tau$, and $s:=\sigma^{2}/(h+\sigma^{2})\in(0,1)$,
  • [Forward] $y_{k+1}\sim\mu^{Y|X=x_{k}}=\gamma_{h}(\cdot-x_{k})$.
  • [Backward] $x_{k+1}\sim\mu^{X|Y=y_{k+1}}\propto\gamma_{sh}(\cdot-sy_{k+1})\,\mathds{1}_{\mathcal{K}}(\cdot)$. To this end, draw $x_{k+1}\sim\mathcal{N}(sy_{k+1},shI_{d})$ repeatedly until $x_{k+1}\in\mathcal{K}$. If this rejection loop exceeds $\tau$ trials, declare failure and restart from scratch with a new sample $x_{0}\sim\mu_{0}^{X}$.
As in the uniform-sampling case, its analysis is split into two parts: a mixing analysis and per-iteration complexity. The failure probability and per-iteration analysis will be proceeded in a similar fashion, but the mixing analysis can be improved due to the hypercontractivity of $\mathsf{PS}$.
Mixing analysis.
We already established the convergence rate of $\mathsf{PS}$ under ($\mathsf{LSI}$) (see Corollary 1.3). It is classical that $C_{\mathsf{LSI}}(\mu^{X})\leq\sigma^{2}$ (e.g., use the Bakry–Émery criterion, and the convex boundary term does not harm in satisfying this criterion; see [KLA09, BJ24]). Thus, in order to be $\varepsilon$-close to $\mu^{X}$ in $\mathcal{R}_{q}$, for $q_{0}>1$, $\mathsf{PS}_{\textup{Gauss}}$ suffices to iterate
$N\lesssim h^{-1}\sigma^{2}\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{X}\mathbin{\|}\mu^{X})}{(2\wedge q_{0}-1)\,\varepsilon}\qquad\text{iterations}\,.$
Per-step analysis.
The analysis is analogous to §2.2, with $\pi_{k}^{Y}$ replaced by $\mu_{k}^{Y}$ and §2.1 of [KV25a] replaced by §2.2. With the parameter choices $Z=\frac{16NM_{2}}{\eta}$, $c=\frac{\log\log Z}{10\log Z}$, $h=\frac{c}{d^{2}}$, $\tau=Z^{2}\log^{3}Z$, the failure probability satisfies $\mathbb{E}_{\mu_{k}^{Y}}[(1-\ell)^{\tau}]\leq\frac{\eta}{N}$, and enforcing $q\geq 1+(1\vee\log\tau)$ yields $\mathbb{E}_{\mu_{k}^{Y}}[\frac{1}{\ell}\wedge\tau]\leq 5M_{q_{0}}\log^{4}Z$.
Proposition 4.1.
Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with $B_{1}(0)\subset\mathcal{K}$, given access to a membership oracle for $\mathcal{K}$, and $\mu\propto\gamma_{\sigma^{2}}\cdot\mathds{1}_{\mathcal{K}}$ be the Gaussian truncated to $\mathcal{K}$. Given $\varepsilon>0$ , $\eta\in(0,1/2)$, $q\geq q_{0}\geq 2$, and an initial distribution $\mu_{0}$ with $M_{q_{0}}=\lVert\nicefrac{{\mathrm{d}\mu_{0}}}{{\mathrm{d}\mu}}\rVert_{L^{q_{0}}(\mu)}$, initialize $\mathsf{PS}_{\textup{Gauss}}$ with $\mu_{0}$ and iterate it $N=\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\eta})$ times with $h=(10d^{2}\log Z)^{-1}$ and $\tau=Z^{2}\log^{3}Z$ for $Z=\frac{16NM_{2}}{\eta}$. If $q_{0}\geq 1+\log\tau$ and $M_{q_{0}}\leq 10$, then
  • The sufficient condition on $q_{0}$ is satisfied when $q_{0}\geq 2\vee O(\log\frac{N}{\eta})$.
  • With probability at least $1-\eta$, $\mathsf{PS}_{\textup{Gauss}}$ iterates $N$ times without failure. Conditioned on this success, the output $Z\sim\nu$ satisfies $\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\leq\varepsilon+2\log\frac{1}{1-\eta}$, with $\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log^{5}\frac{1}{\eta})$ membership queries in expectation.
Proof.
For $q\geq q_{0}\geq 2$ and $(M_{2}\leq)M_{q_{0}}\leq 10$, the uncapped $\mathsf{PS}_{\textup{Gauss}}$ (with the choice of $h=cd^{-2}=\frac{\log\log Z}{10d^{2}\log Z}$ and $Z\asymp NM_{2}/\eta$) returns a sample $X_{N}$ such that $\mathcal{R}_{q}(\operatorname{law}X_{N}\mathbin{\|}\mu^{X})\leq\varepsilon$ after iterating
$h^{-1}\sigma^{2}\log\frac{q\log M_{2}}{\varepsilon}\quad\text{times}\,.$
Repeating the similar argument as in the uniform sampling case, the required number of iterations is of order
$N=\widetilde{O}\bigl(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\eta}\bigr)\,.$ (4.1)
A sufficient condition for ensuring $q_{0}\geq 2\vee\log\tau$ is simply
$1+\log\tau\lesssim\log\frac{N}{\eta}\lesssim\log\widetilde{O}\bigl(\frac{d^{2}\sigma^{2}}{\eta}\log\frac{q}{\varepsilon}\log\frac{1}{\eta}\bigr)\,$, (4.2)
Under this condition, the per-step complexity is $O(\log^{4}Z)$, so the expected number of membership queries for $\mathsf{PS}_{\textup{Gauss}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{q}$ in one attempt is
$\widetilde{O}\bigl(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log^{5}\frac{1}{\eta}\bigr)\,$,
and the success probability satisfies $\mathbb{P}(S)\geq\eta$. As in the analysis of $\mathsf{PS}_{\textup{unif}}$, the conditional law of the final output is distributed as $\hat{\mu}_{N}^{X}$ satisfying $\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\leq\varepsilon+2\log\frac{1}{1-\eta}$. ∎
Combining all these together, we can prove Theorem 1.4 on the query complexity of $\mathsf{PS}_{\textup{Gauss}}$.
Proof of Theorem 1.4.
Take $\eta=\varepsilon<0.01$, and restart the algorithm upon failure. To invoke the proposition above, $q_{0}$ should satisfy $q_{0}\geq 2\vee O(\log N)$ for $N=\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\varepsilon})$. To ensure $q_{0}\geq\log\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\varepsilon})(\geq\log N)$, it suffices to enforce
$q_{0}\geq 2\vee\widetilde{O}\bigl(\log d^{2}\sigma^{2}+\log\log\frac{q}{\varepsilon}\bigr)\,$,
so we will set $q_{0}$ to this as a required base order. We then restrict to $q\geq 2\vee\widetilde{O}(\log d^{2}\sigma^{2}+\log\log\frac{1}{\varepsilon})$ so that $q\geq q_{0}$ holds. Therefore, by Proposition 4.1, the restart version of $\mathsf{PS}_{\textup{Gauss}}$ uses $\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log^{5}\frac{1}{\varepsilon})$ total queries in expectation, and its output is distributed as $\hat{\mu}_{N}^{X}$ satisfying $\mathcal{R}_{q}(\hat{\mu}_{N}^{X}\mathbin{\|}\pi^{X})\leq 3\varepsilon$. Retaking $\varepsilon\leftarrow\varepsilon/3$, we complete the proof. ∎
4.2 Functional inequalities to closeness and early-stopping
We will show that two consecutive Gaussians are $O(1)$-close in $\mathcal{R}_{q}$, finding a connection to ($\mathsf{LSI}$). We then show that annealing can be stopped around $\sigma_{i}^{2}\asymp qR\Lambda^{1/2}\log^{1/2}d$, much earlier than $\sigma_{i}^{2}\asymp R^{2}$, finding a connection to ($\mathsf{PI}$). Moreover, the second result improves a mathematical result established in [KV25a] as an immediate corollary. We then prove Theorem 1.5 on the query complexity of our annealing algorithm.
Mollification.
To avoid any regularity issues, we work with mollifications of logconcave distributions. As detailed in [EVA10, §C.4], the standard mollifier $\eta\in C^{\infty}(\mathbb{R}^{d})$ is defined as
[…]
where the constant $C$ is chosen to ensure $\int\eta=1$. For $\varepsilon>0$, one can consider
$\eta_{\varepsilon}(x):=\frac{1}{\varepsilon^{d}}\,\eta\bigl(\frac{x}{\varepsilon}\bigr)\,$,
whose support is bounded by $B_{\varepsilon}(0)$. Then, for locally integrable $f:\mathbb{R}^{d}\to\mathbb{R}$, its mollification $f_{\varepsilon}:=f*\eta_{\varepsilon}$ satisfies $f_{\varepsilon}\in C^{\infty}(\mathbb{R}^{d})$ and $f_{\varepsilon}\to f$ almost everywhere as $\varepsilon\to 0$.
Streamlined analysis.
Recall Scheffé’s lemma that if a sequence $\pi_{n}$ of probability densities converges to a probability measure $\pi$ almost everywhere, then $\pi_{n}$ converges to $\pi$ in $\mathsf{TV}$-distance (thus weakly). Hence, $\pi_{t}:=\pi*\eta_{t}\to\pi$ weakly as $t\to 0$ due to the property of the mollifier. Then, it is a property of weak convergence that $\int\gamma_{\sigma^{2}}\,\mathrm{d}\pi_{t}\to\int\gamma_{\sigma^{2}}\,\mathrm{d}\pi$ for any Gaussian density $\gamma_{\sigma^{2}}$. Therefore, probability measures $\pi_{t}\gamma_{\sigma^{2}}$ converge to $\pi\gamma_{\sigma^{2}}$ almost surely (and thus weakly).
A Rényi version of LSI was established in [VW23, Lemma 5] that for $q>1$ and smooth $\mu/\nu$,
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{qC_{\mathsf{LSI}}(\nu)}{2}\,\mathsf{RFI}_{q}(\mu\mathbin{\|}\nu)\,$, ($\mathsf{R\text{-}LSI}$)
where $(\nicefrac{{\mu}}{{\nu}})^{q}\log\nicefrac{{\mu}}{{\nu}}$ is continuous, locally uniform, and integrable.
Lemma 4.2 (Closeness).
Let $\pi$ be a logconcave probability measure over $\mathbb{R}^{d}$ with $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$. For $\alpha,\sigma^{2}>0$ and $q>1$, probability measures $\pi\gamma_{\sigma^{2}}$ and $\pi\gamma_{\sigma^{2}(1+\alpha)}$ satisfy
$\mathcal{R}_{q}(\pi\gamma_{\sigma^{2}}\mathbin{\|}\pi\gamma_{\sigma^{2}(1+\alpha)})\leq\frac{q^{2}\alpha^{2}R^{2}}{2\sigma^{2}}\,.$
In particular, they are $O(1)$-close in $\mathcal{R}_{q}$ if $\alpha\leq\nicefrac{{\sigma}}{{qR}}$.
Proof.
Let $\mu\propto\pi\gamma_{\sigma^{2}}$ and $\nu\propto\pi\gamma_{\sigma^{2}(1+\alpha)}$, which clearly satisfies $\mu\ll\nu$. For $t>0$, consider mollified probability measures $\mu_{t}\propto\pi_{t}\gamma_{\sigma^{2}}$ and $\nu_{t}\propto\pi_{t}\gamma_{\sigma^{2}(1+\alpha)}$. On $\operatorname{supp}\nu_{t}$,
$\nabla\log\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}=\nabla\log\frac{\gamma_{\sigma^{2}}}{\gamma_{\sigma^{2}(1+\alpha)}}=\frac{\alpha x}{(1+\alpha)\,\sigma^{2}}\,$,
which admits a smooth extension (as $\frac{\alpha x}{(1+\alpha)\sigma^{2}}$) to $\mathbb{R}^{d}$. Then, for some $\delta>0$ independent of $t$,
$\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\nu_{t})=q\,\frac{\mathbb{E}_{\nu_{t}}[(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}})^{q}\,|\nabla\log\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}|^{2}]}{\mathbb{E}_{\nu_{t}}[(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}})^{q}]}=\frac{q\alpha^{2}}{\sigma^{4}\,(1+\alpha)^{2}}\,\frac{\mathbb{E}_{\nu_{t}}[(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}})^{q}\,|\cdot|^{2}]}{\mathbb{E}_{\nu_{t}}[(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}})^{q}]}=\frac{q\alpha^{2}}{\sigma^{4}\,(1+\alpha)^{2}}\,\frac{\mathbb{E}_{\pi_{t}}[\gamma_{\delta}\,|\cdot|^{2}]}{\mathbb{E}_{\pi_{t}}\gamma_{\delta}}\,.$ (4.4)
Using ($\mathsf{R\text{-}LSI}$) with $C_{\mathsf{LSI}}(\nu_{t})\leq\sigma^{2}\,(1+\alpha)$,
$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{qC_{\mathsf{LSI}}(\nu_{t})}{2}\,\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{q^{2}\alpha^{2}}{2\sigma^{2}}\,\frac{\mathbb{E}_{\pi_{t}}[\gamma_{\delta}\,|\cdot|^{2}]}{\mathbb{E}_{\pi_{t}}\gamma_{\delta}}\,.$
As noted earlier, $\mu_{t}\to\mu$ and $\nu_{t}\to\nu$ weakly. By [AFP00, Theorem 2.34], the lower semi-continuity of $\mathcal{R}_{q}$ implies that $\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})$. Also, $\lim_{t\downarrow 0}\mathbb{E}_{\pi_{t}}[\gamma_{\delta}\,|\cdot|^{2}]\,/\,\mathbb{E}_{\pi_{t}}\gamma_{\delta}=\mathbb{E}_{\pi}[\gamma_{\delta}\,|\cdot|^{2}]\,/\,\mathbb{E}_{\pi}\gamma_{\delta}$ due to the weak convergence of $\pi_{t}\to\pi$. Thus,
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{q^{2}\alpha^{2}}{2\sigma^{2}}\,\frac{\mathbb{E}_{\pi}[\gamma_{\delta}\,|\cdot|^{2}]}{\mathbb{E}_{\pi}\gamma_{\delta}}\,.$
We now need a general fact that for any integrable function $\rho:\mathbb{R}^{d}\to\mathbb{R}_{+}$,
$\frac{\int|\cdot|^{2}\,\gamma_{\delta}\,\mathrm{d}\rho\,}{\int\gamma_{\delta}\,\mathrm{d}\rho}\leq\frac{\int|\cdot|^{2}\,\mathrm{d}\rho}{\int\mathrm{d}\rho}=\mathbb{E}_{\rho}[|\cdot|^{2}]\,.$ (4.5)
Observe that $\operatorname{cov}_{\rho}(f,g)=\frac{1}{2}\,\mathbb{E}_{(X,Y)\sim\rho^{\otimes 2}}[\{f(X)-f(Y)\}\,\{g(X)-g(Y)\}]$. As $f(\cdot)=|\cdot|^{2}$ increases in $|\cdot|$ while $g=\gamma_{\delta}$ decreases, $0\geq\operatorname{cov}_{\rho}(f,g)=\mathbb{E}_{\rho}[fg]-\mathbb{E}_{\rho}f\,\mathbb{E}_{\rho}g$, which proves the claim. Using this fact with $\rho=\pi$, we complete the proof of the main claim. ∎
Lemma 4.3 (Early-stopping).
Let $\pi$ be a logconcave probability measure over $\mathbb{R}^{d}$ with $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$. For $q\geq 2$, if $\sigma^{2}\gtrsim qR\,\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{1/2}d$, then $\mathcal{R}_{q}(\pi\gamma_{\sigma^{2}}\mathbin{\|}\pi)=O(1)$.
Proof.
Let $\mu\propto\pi\gamma_{\sigma^{2}}$ be the probability measure, and $\mu_{t}$ and $\pi_{t}$ the mollifications of $\mu$ and $\pi$. Recall from ($\mathsf{R\text{-}PI}$) that
$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\pi_{t})\leq\log\Bigl(1-\frac{qC_{\mathsf{PI}}(\pi_{t})}{4}\,\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\pi_{t})\Bigr)^{-1}\,.$ (4.6)
By [MIL09, §5] or [CG20, Theorem 3.13], if two logconcave probability measures $\nu_{1}$ and $\nu_{2}$ satisfy $\lVert\nu_{1}-\nu_{2}\rVert_{\mathsf{TV}}\leq\varepsilon$ for $\varepsilon\in(0,1)$, then
$C_{\mathsf{PI}}(\nu_{1})\lesssim\frac{1\vee\log\frac{1}{1-\varepsilon}}{(1-\varepsilon)^{2}}\,C_{\mathsf{PI}}(\nu_{2})\,.$
Since $\pi_{t}\to\pi$ in $\mathsf{TV}$-distance (by Scheffé), and $\pi_{t}$ is logconcave (by Prékopa–Leindler), we have $\limsup_{t\downarrow 0}C_{\mathsf{PI}}(\pi_{t})\lesssim C_{\mathsf{PI}}(\pi)$. Next, emulating the proof of Lemma 4.2 with $\alpha=\infty$, (4.4) yields
$\limsup_{t\downarrow 0}\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\pi_{t})=\frac{q}{\sigma^{4}}\,\frac{\mathbb{E}_{\pi}[\gamma_{\sigma^{2}}\,|\cdot|^{2}]}{\mathbb{E}_{\pi}\gamma_{\sigma^{2}}}\leq\frac{q\,\mathbb{E}_{\pi}[|\cdot|^{2}]}{\sigma^{4}}=\frac{qR^{2}}{\sigma^{4}}\,.$
Putting these together and using $C_{\mathsf{PI}}(\pi)\lesssim\lVert\operatorname{cov}\pi\rVert\log d$ [KLA23],
$\limsup_{t\downarrow 0}\frac{qC_{\mathsf{PI}}(\pi_{t})}{4}\,\mathsf{RFI}_{q}(\mu_{t}\mathbin{\|}\pi_{t})\lesssim\frac{q^{2}R^{2}\,\lVert\operatorname{cov}\pi\rVert\log d}{\sigma^{4}}\,.$
Thus, if $\sigma^{2}\gtrsim qR\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{1/2}d$, then the RHS can be made smaller than (say) $0.1$. Putting this back to (4.6), it follows that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)=O(1)$. ∎
Covariance of strongly logconcave distributions.
[KV25a] showed that $\lVert\operatorname{cov}\pi\gamma_{\sigma^{2}}\rVert\lesssim\lVert\operatorname{cov}\pi\rVert$ if $h\gtrsim R\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{2}d\log^{2}\nicefrac{{R^{2}}}{{\Lambda}}$ for $R=\mathbb{E}_{\pi}|\cdot|$. This can be improved as an immediate corollary of the early-stopping result. For the barycenter $b_{\pi}$ of $\pi$, any unit vector $v\in\mathbb{R}^{d}$, and $\mu=\pi\gamma_{\sigma^{2}}$ with $\sigma^{2}\gtrsim R\,\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{1/2}d$,
$\Bigl(\int\bigl(v^{\mathsf{T}}(x-b_{\pi})\bigr)^{2}\,\mathrm{d}\mu(x)\Bigr)^{2}\leq\Bigl\|\frac{\mathrm{d}\mu}{\mathrm{d}\pi}\Bigr\|_{L^{2}(\pi)}^{2}\int\bigl(v^{\mathsf{T}}(x-b_{\pi})\bigr)^{4}\,\mathrm{d}\pi(x)\lesssim\Bigl\|\frac{\mathrm{d}\mu}{\mathrm{d}\pi}\Bigr\|_{L^{2}(\pi)}^{2}\lVert\operatorname{cov}\pi\rVert^{2}\,$,
where we used the reverse Hölder in the last inequality. Thus, $\lVert\operatorname{cov}\mu\rVert\lesssim\lVert\frac{\mathrm{d}\mu}{\mathrm{d}\pi}\rVert_{L^{2}(\pi)}\lVert\operatorname{cov}\pi\rVert\,.$ Since $\mathcal{R}_{2}(\mu\mathbin{\|}\pi)=O(1)$, we have
$\lVert\operatorname{cov}\mu\rVert\lesssim\lVert\operatorname{cov}\pi\rVert\,.$
4.3 Rényi-warmness generation
Now that all pieces are ready, we show that an $O(1)$-warm start in $\mathcal{R}_{q}$ (equivalently, $L^{q}=O(1)$) can be obtained by using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$ membership queries, where $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$.
Algorithm.
Given $\varepsilon_{\textup{unif}}=\eta_{\textup{unif}}\in(0,1/10)$ and $q\geq 2$, we pick a “base” order $q_{0}(\leq q)$ that will be preserved through annealing. To get a failure-free algorithm, we restart the following algorithm from scratch if $\mathsf{PS}_{\textup{Gauss}}$ declares failure in the middle: for $m$ the number of total phases,
  1. 1.
    Sample from $\mu_{1}=\pi\gamma_{d^{-1}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$, and set $\bar{\mu}_{1}:=\mu_{1}$.
  2. 2.
    While $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{Gauss}}$ such that $\mathcal{R}_{q_{0}}(\bar{\mu}_{i+1}\mathbin{\|}\mu_{i+2})\leq 1$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{\sigma_{i}}{q_{0}R}\bigr)\,.$
    1. (a)
      For $i<m$, run $\mathsf{PS}_{\textup{Gauss}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$ with $N_{i}=\widetilde{O}(d^{2}\sigma_{i}^{2}\log q_{0}\log m)$, where we set $\varepsilon_{i}=1/10$ and $\eta_{i}=(10m)^{-1}$.
    2. (b)
      For $i=m$, run $\mathsf{PS}_{\textup{Gauss}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{q}$ with $N_{i}=\widetilde{O}(qd^{2}R\Lambda^{1/2}\log m)$, where we set $\varepsilon_{i}=1/10$ and $\eta_{i}=(10m)^{-1}$.
Note that $N_{i}$ is chosen to ensure the target accuracies of $\mathsf{PS}_{\textup{Gauss}}$ in each phase (see (4.1)). As seen shortly, given $q\geq\widetilde{O}(1)$, we can take $q_{0}=O(\log q)$ so that the annealing has “linear” dependence on qq. [^2] If we naïvely attempt to preserve $q$-warmness, then the last doubling of $\sigma_{i}^{2}$ incurs (roughly) $qd^{2}\sigma_{m}R\lesssim q^{3/2}d^{2}R^{3/2}\Lambda^{1/4}$ queries (i.e., $q^{3/2}$-dependence).
Choice of parameters.
As $m=\widetilde{O}(q_{0}d^{1/2}R\log q)$, we have $N_{i}\leq\widetilde{O}(qd^{2}R\Lambda^{1/2})=:N_{\max}(q)$. Also, $N_{\textup{unif}}:=\widetilde{O}(qd^{2}\Lambda\log\frac{1}{\eta_{\textup{unif}}}\log\frac{1}{\varepsilon_{\textup{unif}}})$. Recall that $q_{0}$ and $q$ should satisfy sufficient conditions to invoke the established sampling guarantees of $\mathsf{PS}_{\textup{Gauss}}$ and $\mathsf{PS}_{\textup{unif}}$ (Proposition 2.8 and 4.1). Namely, $q\geq 2\vee O(\log\frac{N_{\textup{unif}}}{\eta_{\textup{unif}}})$, $q_{0}\geq 2\vee O(\log mN_{\max})$, and $q\geq q_{0}$. The second condition is implied by $q_{0}\gtrsim\log\widetilde{O}(q_{0}qd^{5/2}R^{2}\Lambda^{1/2}\log q)$, so we will set $q_{0}=\widetilde{O}(\log qdR\Lambda^{1/4})$. To ensure $q\geq q_{0}$, we will restrict $q\geq\widetilde{O}(\log dR\Lambda^{1/4})$. Now, using (2.3) for the first condition, we will assume that
$q\geq 2\vee\widetilde{O}(\log dR\Lambda^{1/4})\vee\widetilde{O}\bigl(\log\frac{d^{2}\Lambda}{\eta_{\textup{unif}}}+\log\log\frac{1}{\varepsilon_{\textup{unif}}}\bigr)\,,\quad q_{0}=\widetilde{O}(\log qdR\Lambda^{1/4})\,.$ (4.7)
By abuse of notation, we sometime write $q\geq\widetilde{O}(1)$ and $q_{0}=\widetilde{O}(\log q)$. We note that $\eta_{i}$ is chosen to bound by $1/10$ the probability that the failure condition is met until the termination of the annealing.
High-level idea of analysis.
We consider an annealing process across a sequence of target distributions $\{\mu_{i}\}_{i=1}^{m}$, where each phase invokes the capped $\mathsf{PS}_{\textup{Gauss}}$ designed to move from (approximately) $\mu_{i}$ to (approximately) $\mu_{i+1}$. Each phase may declare failure due to a capped inner rejection sampler.
Let $S^{(i)}$ be the event that phase $i$ succeeds (no failure of $\mathsf{PS}_{\textup{Gauss}}$ within that phase), and let $S:=\cap_{i=1}^{m-1}S^{(i)}$ be the global success event. Let $C^{(i)}\geq 0$ be the query complexity incurred by phase $i$, and define the one-attempt total cost
$C:=\sum_{i=1}^{m-1}\mathds{1}\bigl\{\bigcap_{j=1}^{i-1}S^{(j)}\bigr\}\,C^{(i)}\,.$
Clearly, $\mathbb{E}C\leq\sum_{i=1}^{m-1}\mathbb{E}C^{(i)}$, so it suffices to upper bound $\mathbb{E}C^{(i)}$. Similarly, letting $F^{(i)}$ denote the event that phase $i$ fails and $q^{(i)}$ denote its cap-hit probability inside phase $i$, we have
$\mathbb{P}(F^{(i)})=\mathbb{E}\bigl[\mathds{1}\bigl\{\cap_{j=1}^{i-1}S^{(j)}\bigr\}\,q^{(i)}\bigr]\leq\mathbb{E}q^{(i)}\,,\qquad\mathbb{P}(S^{c})\leq\sum_{i=1}^{m-1}\mathbb{E}q^{(i)}\,$,
which allows us to bound the overall failure probability via unconditional phase-wise estimates (under the uncapped intermediate laws used in the analysis).
Similar to the analysis of the restarting version of $\mathsf{PS}_{\textup{unif}}$, a restart-until-success annealing repeats independent full attempts (including resampling the initial input from $\mu_{1}$) until $S$ occurs. If $\mathbb{P}(S)\geq 1-\eta$, then the expected number of attempts is at most $(1-\eta)^{-1}$ and
$\mathbb{E}C_{\text{restart}}=\frac{\mathbb{E}C}{\mathbb{P}(S)}\leq\frac{\mathbb{E}C}{1-\eta}\,.$
In summary, we first establish (phase-wise) bounds on unconditional failure probabilities and unconditional expected costs, combine them to control $\mathbb{P}(S)$ and $\mathbb{E}C$ for one attempt, and then translate these into guarantees for the returned-sample distribution $\nu$ and for the restart wrapper.
Proof of Theorem 1.5.
Pick $q$ and $q_{0}$ as in (4.7). The initialization by rejection sampling requires $O(1)$ queries in expectation due to
$\frac{\int_{\mathbb{R}^{d}}e^{-d\,|x|^{2}/2}\,\mathrm{d}x}{\int_{\mathcal{K}}e^{-d\,|x|^{2}/2}\,\mathrm{d}x}=\frac{1}{\gamma_{d^{-1}}(\mathcal{K})}\leq\frac{1}{\gamma_{d^{-1}}(B_{1})}\lesssim 1\,.$
In the annealing phase, suppose that we are given $\bar{\mu}_{i}$ such that $\mathcal{R}_{2q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})\leq\frac{1}{10}$. By a weak triangle inequality for Rényi divergence (with $\lambda=2/3$),
$\mathcal{R}_{q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i+1})\leq 1.1\mathcal{R}_{2q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})+\mathcal{R}_{2q_{0}}(\mu_{i}\mathbin{\|}\mu_{i+1})\leq\frac{1.1}{10}+\frac{1}{2}\leq 1\,$,
where the last inequality follows from Lemma 4.2. By Theorem 1.4, from an $O(1)$-warmness in $\mathcal{R}_{q_{0}}$-divergence (precisely, $\mathcal{R}_{q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i+1})\leq 1$), $\mathsf{PS}_{\textup{Gauss}}$ needs $\widetilde{O}(d^{2}\sigma_{i}^{2}\log^{2}q_{0})$ queries to sample $X_{i+1}^{*}\sim\bar{\mu}_{i+1}$ such that $\mathcal{R}_{2q_{0}}(\bar{\mu}_{i+1}\mathbin{\|}\mu_{i+1})\leq\frac{1}{10}$. Also, doubling of given $\sigma_{i}^{2}$ requires $O(q_{0}R/\sigma_{i})$ phases, so each doubling needs $\widetilde{O}(q_{0}d^{2}\sigma_{i}R)$ queries in total. Since there are at most $O(\log qdR\Lambda^{1/2})$ doublings, and $\sigma_{i}^{2}\leq\sigma_{m}^{2}\asymp qR\Lambda^{1/2}\log^{1/2}d$, we can obtain a sample $X^{*}\sim\bar{\mu}_{m-1}$ such that $\mathcal{R}_{2q_{0}}(\bar{\mu}_{m-1}\mathbin{\|}\mu_{m-1})\leq\frac{1}{10}$, using
$O\bigl(q_{0}q^{1/2}d^{2}R^{3/2}\Lambda^{1/4}\log^{1/4}d\times\log qdR\Lambda^{1/2}\bigr)=\widetilde{O}(q^{1/2}d^{2}R^{3/2}\Lambda^{1/4})\quad\text{queries}\,.$
By hypercontractivity of $\mathsf{PS}$ (Corollary 1.3), iterating $\mathsf{PS}_{\textup{Gauss}}$ started from $X^{*}$ towards $\mu_{m}$ $\widetilde{O}(d^{2}\sigma_{m}^{2}\log q)$ times leads to a new sample $X^{\prime}$ such that $\mathcal{R}_{2q}(\operatorname{law}X^{\prime}\mathbin{\|}\mu_{m})\leq\frac{1}{10}$. Hence, the entire annealing has used
$\widetilde{O}\bigl(q^{1/2}d^{2}R^{3/2}\Lambda^{1/4}+qd^{2}R\Lambda^{1/2}\bigr)\leq\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})\quad\text{queries (due to }R^{2}\geq\Lambda\text{)}\,.$
Note that the early-stopping result (Lemma 4.3) and the weak triangle inequality imply $\mathcal{R}_{q}(\operatorname{law}X^{\prime}\mathbin{\|}\pi)\leq 1$ as desired.
Lastly, restarting the annealing (in case of failure) increases the total query complexity multiplicatively by $10/9$, and the conditional law on success has an extra bias at most $1$ in $\mathcal{R}_{\infty}$. Hence, by $\mathcal{R}_{q}(\operatorname{law}(X^{\prime}\mid\text{success})\mathbin{\|}\pi)\leq 2$. ∎
We now establish the query complexity of convex-body sampling from scratch.
Proof of Corollary 1.6.
We set $\eta_{\textup{unif}}=\varepsilon_{\textup{unif}}=\varepsilon<1/100$ and consider $q\geq\widetilde{O}(1)$. By Theorem 1.1, the expected number of restarts of $\mathsf{PS}_{\textup{unif}}$ is at most $O(1)$, so the warm-start generation will be repeated at most $O(1)$ times. Combining the complexity results in Theorem 1.5 and 1.1, the total complexity of logconcave sampling from scratch is
$\widetilde{O}\bigl(qd^{2}R^{3/2}\Lambda^{1/4}+qd^{2}\Lambda\log^{6}\frac{1}{\varepsilon}\bigr)\,.$
When $1\leq q\leq 2\vee\widetilde{O}(\log\frac{dR}{\varepsilon})=:q_{*}$, we generate the $\mathcal{R}_{q_{*}}$-warmness and use the monotonicity of $\mathcal{R}_{q}$. ∎
5 Logconcave sampling under an evaluation oracle
To extend our approach to logconcave distributions, we generalize what we did for uniform sampling, streamlining algorithms in [KV25a] and obtaining a guarantee for $\mathcal{R}_{q}$-warmness generation.
We begin with analysis for order-preserving logconcave sampling from an $O(1)$-warm start in §5.1 and move to an order-preserving annealing algorithm for warm-start generation in §5.2. Overall, we follow a similar route for the analysis as in the uniform sampling case.
5.1 Zeroth-order logconcave sampling with balanced guarantees
We use the exponential-lifting approach ($\mathsf{exp\text{-}lifting}$) introduced in [KV25b], obtaining an order-preserving guarantee under an $O(1)$-warm start. For $z=(x,t)\in\mathbb{R}^{d+1}$ and $y\in\mathbb{R}^{d+1}$, $\mathsf{PS}_{\textup{exp}}$ alternates $y\sim\pi^{Y|Z=z}=\mathcal{N}(z,hI_{d+1})$ and $z\sim\pi^{Z|Y=y}\propto\mathcal{N}(y-h\alpha,hI_{d+1})|_{\mathcal{K}}$ for $\alpha=de_{d+1}$, where the second step is implemented via rejection sampling, using the proposal $\mathcal{N}(y-h\alpha,hI_{d+1})$ with threshold $\tau$.
A mixing analysis.
Assume that $\pi_{0}^{X}$ is $M_{q}$-warm with respect to $\pi^{X}$. Note that one can generate a sample $(X^{\prime},T^{\prime})$ whose joint law is also $M_{q}$-warm with respect to $\pi^{X,T}$ by first drawing $X^{\prime}\sim\pi_{0}^{X}$ and then sampling $T^{\prime}\sim\pi^{T|X=X^{\prime}}\propto e^{-dt}|_{[V(X^{\prime})/d,\infty)}$.
Since the convergence rate of $\mathsf{PS}$ was already given in §2.1, we can just reuse Proposition 2.4. The only adjustment is that we now require a bound on $C_{\mathsf{PI}}(\pi^{X,T})$, rather than $C_{\mathsf{PI}}(\pi^{X})$, as we run $\mathsf{PS}$ to sample from $\pi^{X,T}$. Since $C_{\mathsf{PI}}(\pi^{X,T})\lesssim\lVert\operatorname{cov}\pi^{X,T}\rVert\log d$ due to log-concavity of $\pi^{X,T}$, and in [KV25b, Lemma 2.5], $\lVert\operatorname{cov}\pi^{X,T}\rVert$ is bounded as
$\lVert\operatorname{cov}\pi^{X,T}\rVert\leq 2\,(\lVert\operatorname{cov}\pi^{X}\rVert+160)\lesssim\lVert\operatorname{cov}\pi^{X}\rVert\vee 1\,$,
we have $C_{\mathsf{PI}}(\pi^{X,T})\lesssim(1\vee\lVert\operatorname{cov}\pi^{X}\rVert)\log d$. Hence, by Lemma 2.5, $\mathsf{PS}_{\textup{exp}}$ with step size $h$ can ensure $(\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq)\mathcal{R}_{q}(\pi_{N}^{X,T}\mathbin{\|}\pi^{X,T})\leq\varepsilon$ after
$N\lesssim q\,\bigl(1\vee h^{-1}\,(1\vee\lVert\operatorname{cov}\pi^{X}\rVert)\log d\bigr)\log\frac{M_{q}}{\varepsilon}\quad\text{iterations}\,.$
Per-step analysis.
It follows from §5.1 in [KV25a] that with the parameter choices $S=\frac{16NM_{2}}{\eta}$, $c=\frac{(\log\log S)^{2}}{13^{2}\log S}$, $\tau=Z^{2}\log^{2}Z$, and $h=c/(13d)^{2}$, the failure probability satisfies $\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq\frac{\eta}{N}$. The per-step complexity is bounded as follows: setting $q=1+\alpha$ with $\alpha\geq 1\vee\log\tau$, it follows that $\mathbb{E}_{\pi_{k}^{Y}}[\frac{1}{\ell}\wedge\tau]\leq 5M_{q}\log S$.
Proposition 5.1.
For a convex function $V:\mathbb{R}^{d}\to\mathbb{R}$, let $\pi\propto e^{-V}$ be the logconcave distribution over $\mathbb{R}^{d}$ with $B_{1}\subset\mathsf{L}_{\pi,g}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$, given access to an evaluation oracle for $V$. Given $\varepsilon>0$, $\eta\in(0,1/2)$, $q\geq 2$, and an initial distribution $\pi_{0}$ with $M_{q}=\lVert\nicefrac{{\mathrm{d}\pi_{0}}}{{\mathrm{d}\pi}}\rVert_{L^{q}(\pi)}$, initialize $\mathsf{PS}_{\textup{exp}}$ with $\pi_{0}$ and iterate it $N=\widetilde{O}(qd^{2}\,(1\vee\Lambda)\log\frac{1}{\varepsilon}\log\frac{1}{\eta})$ times with $h=(13d^{2}\log S)^{-1}$ and $\tau=S^{2}\log^{2}S$ for $S=\frac{16NM_{2}}{\eta}$. If $q\geq 1+\log\tau$ and $M_{q}\leq 10$, then
  • The sufficient condition on $q$ is satisfied when $q\geq 2\vee O(\log\frac{N}{\eta})$.
  • With probability at least $1-\eta$, $\mathsf{PS}_{\textup{exp}}$ iterates $N$ times without failure. Conditioned on this success, the $X$-output $X^{*}\sim\nu$ satisfies $\mathcal{R}_{q}(\nu\mathbin{\|}\pi)\leq\varepsilon+2\log\frac{1}{1-\eta}$, with $\widetilde{O}(qd^{2}\,(1\vee\Lambda)\log\frac{1}{\varepsilon}\log^{2}\frac{1}{\eta})$ evaluation queries in expectation.
Its proof is almost similar to the analysis of $\mathsf{PS}_{\textup{unif}}$ in Proposition 2.8. Putting all these together, we can establish the query complexity of $\mathsf{PS}_{\textup{exp}}$ with balanced guarantees.
Proof of Theorem 1.7.
Take $\varepsilon=\eta<0.01$, and restart $\mathsf{PS}_{\textup{exp}}$ upon failure. Note that the expected number of restarts is at most $0.02$. In order to ensure $q\geq 2\vee O(\log N)$ for $N=\widetilde{O}(qd^{2}\,(1\vee\Lambda)\log^{2}\frac{1}{\varepsilon})$, it suffices to restrict
$q\geq 2\vee\widetilde{O}\bigl(\log(d^{2}\,(1\vee\Lambda))+\log\log\frac{1}{\varepsilon}\bigr)\,.$
By the proposition above, the desired claims follow in a similar way to the proof of Theorem 1.1. ∎
5.2 Rényi-warmness generation for logconcave distributions
We streamline a warm-start algorithm in [KV25a], which generates a warm start for the lifted distribution $\pi\propto e^{-dt}|_{\mathcal{K}}$ by annealing through distributions $\mu_{\sigma^{2},\rho}(x,t)\propto\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)\,\mathds{1}_{\bar{\mathcal{K}}}(x,t)$ for some $\bar{\mathcal{K}}$. In comparison with our warm-start algorithm for uniform distributions, we can also stop the $\sigma^{2}$-annealing earlier than [KV25a], but we could not remove the preprocessing step therein, in order to ensure a finite LSI constant of the annealing distributions. Also, since the complexity of sampling from $\mu_{\sigma^{2},\rho}$ (by $\mathsf{PS}_{\textup{ann}}$ in §5.2.1) is roughly $d^{2}(\sigma^{2}\vee 1)$ (from an $O(1)$-warm start), which has the additive $d^{2}$ queries, we cannot help but following a convoluted approach in [KV25a].
Warm-start generation.
For $l=\log 4e$, one may assume that $x_{0}=0$ (by translation) and consider
$\bar{\mathcal{K}}:=\mathcal{K}\cap\{B_{Rl}(0)\times[-21,13l-6]\}\,$, (5.1)
which satisfies $\pi(\mathbb{R}^{d+1}\backslash\bar{\mathcal{K}})\leq 1/2$ (see [KV25b, Convex truncation and §3.1]). Thus, we can ensure that $\bar{\pi}=\pi|_{\bar{\mathcal{K}}}$ is $\log 2$-close to $\pi$ in $\mathcal{R}_{\infty}$. Below, we let $D:=Rl$ and $\mu_{i}:=\mu_{\sigma_{i}^{2},\rho_{i}}$, and denote by $\bar{\mu}_{i}$ the law of an actual sample such that $\mathcal{R}_{\widetilde{O}(1)}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$.
  1. 1.
    Sample from $\mu_{0}\propto e^{-d\,|x|^{2}/2}|_{\bar{\mathcal{K}}}$ by rejection sampling with proposal $\gamma_{d^{-1}}\otimes\text{Unif}\,([-21,13l-6])$).
  2. 2.
    [Phase I: $\sigma^{2}$-warming] While $d^{-1}\leq\sigma_{i}^{2}\leq 1$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{1}{d^{1/2}}\bigr)\,.$
    At the end of Phase I, run $\mathsf{PS}_{\textup{ann}}$ toward $\mu_{i+1}\propto e^{-|x|^{2}/2-t}|_{\bar{\mathcal{K}}}$, obtaining $\bar{\mu}_{i+1}$.
  3. 3.
    [Phase II: $\rho$-annealing] While $1\leq\rho_{i}\leq d$ (and $\sigma_{i}^{2}\approx 1$), move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{1}{d^{1/2}}\bigr)^{-1}\quad\&\quad\rho_{i+1}=\min\bigl(d,\rho_{i}\,\bigl(1+\frac{1}{d^{1/2}}\bigr)\bigr)\,.$
    This is followed by inner annealing: until $\sigma_{i+1}^{2}\leq 1$, run $\mathsf{PS}_{\textup{ann}}$ toward new $\mu_{i+1}$ defined by
    $\sigma_{i+1}^{2}\leftarrow\sigma_{i+1}^{2}\,\bigl(1+\frac{\sigma_{i+1}}{D}\bigr)\,$,
  4. 4.
    [Phase III: $\sigma^{2}$-annealing] While $\sigma_{i}^{2}\lesssim qD\,(\Lambda^{1/2}\vee 1)\log^{1/2}d$ (and $\rho_{i}=d$), move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{\sigma_{i}}{D}\bigr)\,.$
5.2.1 Proximal sampler for tilted Gaussian
Now consider the sampler for the annealing distributions $\mu_{i}$ proposed in [KV25b]. For $v:=(x,t)\in\mathbb{R}^{d}\times\mathbb{R}$ and $w:=(y,s)\in\mathbb{R}^{d}\times\mathbb{R}$, consider the augmented target given by
$\mu^{V,W}(v,w)\propto\exp\bigl(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t-\frac{1}{2h}\,|w-v|^{2}\bigr)\big|_{\bar{\mathcal{K}}}\,$,
and $\mathsf{PS}_{\textup{ann}}$ with step size $h$ and threshold $\tau$ alternates the following: for $r:=\frac{\sigma^{2}}{h+\sigma^{2}}<1$, $y_{r}:=ry$, and $h_{r}:=rh$,
  • [Forward] $w_{k+1}\sim\mu^{W|V=v_{k}}=\mathcal{N}(v_{k},hI_{d+1})$.
  • [Backward] $v_{k+1}\sim\mu^{V|W=w_{k+1}}\propto\bigl[\mathcal{N}(ry_{k+1},h_{r}I_{d})\otimes\mathcal{N}(s_{k+1}-\rho h,h)\bigr]\big|_{\bar{\mathcal{K}}}$. To this end, draw $v_{k+1}\sim\mathcal{N}(ry_{k+1},h_{r}I_{d})\otimes\mathcal{N}(s_{k+1}-\rho h,h)$ until $v_{k+1}\in\bar{\mathcal{K}}$. If this rejection loop exceeds $\tau$ trials, declare failure.
We establish its query complexity similarly as before.
Proposition 5.2.
Consider the annealing distribution $\mu=\mu_{\sigma^{2},\rho}$ above, given access to an evaluation oracle for $V$. Given $\varepsilon>0$, $\eta\in(0,1/2)$, $q\geq q_{0}\geq 2$, and an initial distribution $\mu_{0}$ with $M_{q_{0}}=\lVert\nicefrac{{\mathrm{d}\mu_{0}}}{{\mathrm{d}\mu}}\rVert_{L^{q_{0}}(\mu)}$, initialize $\mathsf{PS}_{\textup{ann}}$ with $\mu_{0}$ and iterate it $N=\widetilde{O}(d^{2}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}\log\frac{1}{\eta})$ times with $h=(24^{2}d^{2}\log S)^{-1}$ and $\tau=2S^{2}\log^{2}S$ for $S=\frac{16NM_{2}}{\eta}$. If $q_{0}\geq 1+\log\tau$ and $M_{q_{0}}\leq 10$, then
  • The sufficient condition on $q_{0}$ is satisfied when $q_{0}\geq 2\vee O(\log\frac{N}{\eta})$.
  • With probability at least $1-\eta$, $\mathsf{PS}_{\textup{ann}}$ iterates $N$ times without failure. Conditioned on this success, the output $Z\sim\nu$ satisfies $\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\leq\varepsilon+2\log\frac{1}{1-\eta}$, with $\widetilde{O}(d^{2}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}\log^{2}\frac{1}{\eta})$ evaluation queries in expectation.
Proof.
Overall, we follow the analysis of $\mathsf{PS}_{\textup{Gauss}}$ as in Proposition 4.1. Using the Bakry–Émery criterion and bounded perturbation, one can show that $C_{\mathsf{LSI}}(\mu)\lesssim\sigma^{2}\vee 1$ (see [KV25b, Lemma 3.3]). For $q\geq q_{0}>1$, when $\mathsf{PS}_{\textup{ann}}$ starts from $\mu_{0}^{V}$, by Corollary 1.3, in order to be $\varepsilon$-close to $\mu^{V}$ in $\mathcal{R}_{q}$, for $q_{0}>1$, $\mathsf{PS}_{\textup{ann}}$ suffices to iterate
$h^{-1}\,(\sigma^{2}\vee 1)\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{V}\mathbin{\|}\mu^{V})}{(2\wedge q_{0}-1)\,\varepsilon}\qquad\text{iterations}\,.$
Also, [KV25a] showed that under $c=\frac{(\log\log S)^{2}}{4\cdot 24^{2}\log S}$, $h=\frac{c}{24^{2}d^{2}}$, and $\tau=2S^{2}\log^{2}S$ for $S=\frac{16NM_{2}}{\eta}$, the per-step query complexity is bounded by $O(M_{q}\log S)$, and the failure probability is bounded by $\eta/N$.
For $q\geq q_{0}\geq 2$ and $(M_{2}\leq)M_{q_{0}}\leq 10$, the required number $N$ of iterations (of $\mathsf{PS}_{\textup{ann}}$) for $\varepsilon$-closeness in $\mathcal{R}_{q}$ should satisfy $N\gtrsim h^{-1}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}$. Thus, it suffices to take
$N=\widetilde{O}\bigl(d^{2}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}\log\frac{1}{\eta}\bigr)\,.$
Also, a sufficient condition for ensuring $q_{0}\geq 2\vee\log\tau$ is $q_{0}\geq 2\vee O(\log\frac{N}{\eta})$. Under this condition, the per-step complexity is $O(\log S)$, so the expected number of queries for $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{q}$ is
$\widetilde{O}\bigl(d^{2}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}\log^{2}\frac{1}{\eta}\bigr)\,$,
and the success probability is at least $1-\eta$. ∎
5.2.2 Query complexity of Rényi-warmness generation
Overall, we follow the same analysis of the annealing for uniform distributions; see ‘High-level idea of analysis’ in §4.3.
Algorithm.
Given $\varepsilon_{\exp}=\eta_{\exp}\in(0,1/10)$ and $q\geq 2$, we pick a base order $q_{0}(\leq q)$ that will be preserved through annealing, and restart the following algorithm from scratch if $\mathsf{PS}_{\textup{ann}}$ declares failure in the middle.
Recall $D=Rl$ (for a large constant $l>0$ such that $\mathcal{R}_{\infty}(\bar{\pi}\mathbin{\|}\pi)\leq 2$) and $\mu_{i}=\mu_{\sigma_{i}^{2},\rho_{i}}$, and $\bar{\mu}_{i}$ is the law of an actual sample such that $\mathcal{R}_{2q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})\leq\frac{1}{10}$. For $m$ the number of total phases, $\mathsf{PS}_{\textup{ann}}$ is iterated $N_{i}$ times with $\varepsilon_{i}=1/10$ and $\eta_{i}=(10m)^{-1}$ in each phase $i\in[m]$.
  1. 1.
    Sample from $\mu_{0}\propto e^{-d\,|x|^{2}/2}|_{\bar{\mathcal{K}}}$ by rejection sampling with proposal $\gamma_{d^{-1}}\otimes\text{Unif}\,([-21,13l-6])$).
  2. 2.
    [Phase I: $\sigma^{2}$-warming] While $d^{-1}\leq\sigma_{i}^{2}\leq 1$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{1}{(4q_{0}d)^{1/2}}\bigr)\,.$
    At the end of Phase I, run $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$ toward $\mu_{i+1}\propto e^{-|x|^{2}/2-t}|_{\bar{\mathcal{K}}}$, obtaining $\bar{\mu}_{i+1}$.
  3. 3.
    [Phase II: $\rho$-annealing] While $1\leq\rho_{i}\leq d$ (and $\sigma_{i}^{2}\approx 1$), move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{1}{(4q_{0}d)^{1/2}}\bigr)^{-1}\quad\&\quad\rho_{i+1}=\min\bigl(d,\rho_{i}\,\bigl(1+\frac{1}{(4q_{0}d)^{1/2}}\bigr)\bigr)\,.$
    This is followed by inner annealing: until $\sigma_{i+1}^{2}\leq 1$, run $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$ toward new $\mu_{i+1}$ defined by
    $\sigma_{i+1}^{2}\leftarrow\sigma_{i+1}^{2}\,\bigl(1+\frac{\sigma_{i+1}}{q_{0}D}\bigr)\,$,
  4. 4.
    [Phase III: $\sigma^{2}$-annealing] While $\sigma_{i}^{2}\lesssim qD\,(\Lambda^{1/2}\vee 1)\log^{1/2}d$ (and $\rho_{i}=d$), move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{\sigma_{i}}{q_{0}D}\bigr)\,.$
    When i=m1i=m-1 (i.e., the last phase), we move from $\bar{\mu}_{m-1}$ to $\bar{\mu}_{m}$ by $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{q}$.
Choice of parameters.
Note that the number $m$ of inner phases is bounded as
$m\lesssim(q_{0}d)^{1/2}\log d+(q_{0}d)^{1/2}\log d\times\frac{q_{0}^{1/2}D}{d^{1/2}}+q_{0}D\log\bigl(qD\,(\Lambda^{1/2}\vee 1)\bigr)=\widetilde{O}\bigl(q_{0}d^{1/2}D\log q(1\vee\Lambda^{1/2})\bigr)\,.$
Hence, $N_{i}=\widetilde{O}(d^{2}\,(\sigma_{i}^{2}\vee 1)\log^{2}q_{0}m)\leq\widetilde{O}(qd^{2}D\,(\Lambda^{1/2}\vee 1))=:N_{\max}$ and $N_{\exp}:=\widetilde{O}(qd^{2}\,(\Lambda\vee 1)\log\frac{1}{\eta_{\exp}}\log\frac{1}{\varepsilon_{\exp}})$.
We now find sufficient conditions for $q_{0}$ and $q$ to invoke the established sampling guarantees of $\mathsf{PS}_{\textup{ann}}$ and $\mathsf{PS}_{\textup{exp}}$ (see Proposition 5.1 and 5.2); namely, $q\geq 2\vee O(\log\frac{N_{\exp}}{\eta_{\exp}})$, $q_{0}\geq 2\vee O(\log mN_{\max})$, and $q\geq q_{0}$.
Repeating a similar computation as in the uniform-sampling case, we will take $q_{0}=2\vee\widetilde{O}(\log\{qdD\,(\Lambda\vee 1)\})$ and
$q\gtrsim 2\vee\widetilde{O}\bigl(\log\frac{d^{2}\,(\Lambda\vee 1)}{\eta_{\exp}}+\log\log\frac{1}{\varepsilon_{\exp}}\bigr)\vee\widetilde{O}\bigl(\log\{dD\,(\Lambda\vee 1)\}\bigr)\,.$ (5.2)
Also, due to the choice of $\eta_{i}$, the probability that the failure condition is met until the termination of the annealing is bounded by $\frac{1}{10}$.
Complexity analysis.
We first recall a result on closeness of two logconcave distributions from [KV25a, Lemma 4.4].
Lemma 5.3.
Let $\mathrm{d}\nu\propto e^{-V}\,\mathrm{d}x$ and $\mathrm{d}\mu\propto e^{-(1+\alpha)\,V}\,\mathrm{d}x$ be logconcave probability measures over $\mathbb{R}^{d}$. For $q>1$ and $\delta>0$ such that $1-q\delta>0$,
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\begin{cases}\frac{qd\alpha^{2}}{2}&\text{if }\alpha\geq 0\,,\\ \frac{qd\alpha^{2}}{1-q\delta}&\text{if }\alpha\in[-\delta/2,0]\,.\end{cases}$
We now bound the complexity of the overall algorithm as $\widetilde{O}(d^{2.5}+qd^{2}\,(R^{3/2}\vee 1)(\Lambda^{1/4}\vee 1))$.
Proof of Theorem 1.8.
We set $\eta_{\exp}=\varepsilon_{\exp}=\varepsilon<1/100$ and consider $q_{0}$ and $q\geq 2\vee\widetilde{O}(\log\frac{dD}{\varepsilon})$ (to satisfy (5.2)). At initialization, $\mu_{0}$ can be sampled by rejection sampling with $O(1)$ queries in expectation (see [KV25b, Lemma 5.6]). In Phase I, any consecutive distributions satisfy $\mathcal{R}_{2q_{0}}(\mu_{i}\mathbin{\|}\mu_{i+1})\leq 1$ by Lemma 5.3. Since there are $O(q_{0}^{1/2}d^{1/2}\log d)$ many phases, with each one requiring $\widetilde{O}(d^{2}\log^{2}q_{0}m)$ queries from an $O(1)$-warm start, the query complexity of Phase I is $\widetilde{O}(q_{0}^{1/2}d^{2.5}\log^{2}q_{0}m)$. Meanwhile, we relay $\mathcal{R}_{q_{0}}$-warmness through the weak triangle inequality (as done in the proof of Theorem 1.5).
In Phase II, closeness during the simultaneous annealing (called the outer annealing hereafter) is justified again by Lemma 5.3, and $\mathsf{PS}_{\textup{ann}}$ requires $\widetilde{O}(d^{2}\log^{2}q_{0}m)$ queries for each outer annealing. Each outer annealing is followed by $O(q_{0}^{1/2}D/d^{1/2})$ many inner phases, where $\sigma^{2}$ goes back to $1$ again. Closeness within the inner annealing follows from Lemma 4.2, and $\mathsf{PS}_{\textup{ann}}$ requires $\widetilde{O}(d^{2}\log^{2}q_{0}m)$ queries. Thus, completion of one outer annealing (combined with the inner annealing) requires $\widetilde{O}((d^{2}+q_{0}^{1/2}d^{1.5}D)\log^{2}q_{0}m)$ queries. As there are $O(q_{0}^{1/2}d^{1/2}\log d)$ outer phases, the total complexity of Phase II is $\widetilde{O}((q_{0}^{1/2}d^{2.5}+q_{0}d^{2}D)\log^{2}q_{0}m)$.
In Phase III, closeness of annealing distributions follow again from Lemma 4.2. Since doubling of initial $\sigma^{2}$ requires $O(q_{0}D/\sigma)$ phases, with phase requiring $\widetilde{O}(d^{2}\sigma^{2}\log^{2}q_{0}m)$ queries from an $O(1)$-warm start, each doubling uses $\widetilde{O}(q_{0}d^{2}\sigma D\log^{2}q_{0}m)$ queries. As $\sigma^{2}$ is increased while $\sigma^{2}\lesssim qD\,(\Lambda^{1/2}\vee 1)\log^{1/2}d$, it takes Phase III $\widetilde{O}(q^{1/2}q_{0}d^{2}D^{3/2}\,(\Lambda^{1/4}\vee 1)\log^{2}q_{0}m)$ queries to obtain a sample whose law is $O(1)$-close to $\mu_{m}$ in $\mathcal{R}_{q_{0}}$-divergence. Due to hypercontractivity of $\mathsf{PS}$ again, iterating $\mathsf{PS}_{\textup{ann}}$ toward $\mu_{m}$ $O(d^{2}\sigma_{m}^{2}\log^{2}qm)$ times returns a sample $Z^{*}$ such that $\mathcal{R}_{2q}(\operatorname{law}Z^{*}\mathbin{\|}\mu_{m})\leq 1$. Hence, Phase III uses
$\widetilde{O}\bigl(q^{1/2}q_{0}d^{2}D^{3/2}\,(\Lambda^{1/4}\vee 1)\log^{2}q_{0}m+qd^{2}D\,(\Lambda^{1/2}\vee 1)\log^{2}qm\bigr)=\widetilde{O}\bigl(qd^{2}D^{3/2}(\Lambda^{1/4}\vee 1)\log^{2}qm\bigr)$
queries in total. Also, by the early-stopping result and weak triangle inequality, we have $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\bar{\pi})\leq\frac{1.1}{10}+\frac{1}{2}$. Since $\mathcal{R}_{\infty}(\bar{\pi}\mathbin{\|}\pi)\leq\log 2$, we obtain that $\mathcal{R}_{q}(\operatorname{law}Z^{*}\mathbin{\|}\pi)\leq 1$.
As $D\asymp R$, adding up complexities of each phase bound the total query complexity as
$\widetilde{O}\bigl(d^{2.5}+qd^{2}R^{3/2}\,(\Lambda^{1/4}\vee 1)\bigr)\,.$
Lastly, restarting the annealing (in case of failure) increases the total query complexity multiplicatively by $10/9$, and the conditional law on success has an extra bias at most $\log\frac{10}{9}$ in $\mathcal{R}_{\infty}$. Hence, by $\mathcal{R}_{q}(\operatorname{law}(X^{\prime}\mid\text{success})\mathbin{\|}\pi)\leq 2$. ∎
We establish the query complexity of general logconcave sampling from scratch.
Proof of Corollary 1.9.
We set $\eta_{\exp}=\varepsilon_{\exp}=\varepsilon<1/100$ and consider $q_{0}$ and $q\geq 2\vee\widetilde{O}(\log\frac{dD}{\varepsilon})$. By Theorem 1.7, the expected number of restarts of $\mathsf{PS}_{\textup{exp}}$ is at most $O(1)$, so the warm-start generation will be repeated at most $O(1)$ times. Combining the complexity results in Theorem 1.8 and 1.7, the total complexity of logconcave sampling from scratch is
$\widetilde{O}\bigl(d^{2.5}+qd^{2}R^{3/2}\,(\Lambda^{1/4}\vee 1)+qd^{2}\,(\Lambda\vee 1)\log^{3}\frac{1}{\varepsilon}\bigr)\,.$
When $1\leq q\leq q_{*}$, we generate the $\mathcal{R}_{q_{*}}$-warmness and use the monotonicity of $\mathcal{R}_{q}$, which introduces an extra factor of $q_{*}=\widetilde{O}(1)$. ∎
6 Obstruction to further acceleration
In [KV25a], it was conjectured that for the uniform distribution over any isotropic convex body $\pi$ in $\mathbb{R}^{d}$, the largest eigenvalue $\lambda_{t}:=\lVert\operatorname{cov}\pi\gamma_{t}\rVert$ remains as $\widetilde{O}(1)$ for any $t>0$. It was indeed shown that $\lambda_{t}=O(1)$ if $t\leq 1$ or $t\gtrsim R\Lambda^{1/2}\log^{2}d\log^{2}\frac{R^{2}}{\Lambda}$. Note that the latter has been slightly improved to $R\Lambda^{1/2}\log^{1/2}d$ through Lemma 4.3. The rationale behind the conjecture is, as $t$ grows, the Gaussian factor $\gamma_{t}$ push the mass toward the boundary of $\mathcal{K}$, which tends to increase $\lambda_{t}$.
To tackle this question, let us define quadratic-tilt stability.
Definition 6.1 (Quadratic-tilt stability).
Let $\pi$ be a probability measure in $\mathbb{R}^{d}$. For $\lambda>0$, we say that $\pi$ is $\lambda$-quadratic-tilt stable if for any $t>0$,
$\lVert\operatorname{cov}\pi\gamma_{t}\rVert\leq\lambda\,.$
Hence, the conjecture is simply to establish $\widetilde{O}(1)$-quadratic-tilt stability of the uniform distribution over any isotropic convex body.
In §6.1, we show that one can afford more aggressive annealing under $\widetilde{O}(1)$-quadratic-tilt stability, obtaining the $d^{2.5}$-complexity of isotropic logconcave sampling (from scratch). This improves the current complexity of $d^{2.75}$ and matches the $d^{2.5}$ iteration complexity of the $\mathsf{Speedy\ walk}$ in [LV24]. However, in §6.2, we disprove the conjecture, illustrating the limitation of the current annealing approach.
6.1 Acceleration under quadratic-tilt stability
We illustrate how the acceleration is feasible under $\lambda_{t}=\widetilde{O}(1)$ for any $t>0$. Recall the warm-start generation algorithm: (1) sample from $\mu_{1}=\pi\gamma_{d^{-1}}$ by rejection sampling, and (2) while $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{Gauss}}$ such that $\mathcal{R}_{q_{0}}(\bar{\mu}_{i+1}\mathbin{\|}\mu_{i+2})\leq 1$, where $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,(1+\nicefrac{{\sigma_{i}}}{{q_{0}R}})$. If $\lambda_{t}=\widetilde{O}(1)$ for any $t>0$, then the main annealing can be improved to $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,(1+\nicefrac{{\sigma_{i}^{2}}}{{q_{0}R}})$. Note that this allows faster annealing when $\sigma_{i}^{2}\geq 1$.
Proposition 6.2.
Let $\pi$ be any isotropic logconcave distribution over $\mathbb{R}^{d}$ with $\lVert\operatorname{cov}\pi\gamma_{t}\rVert=\widetilde{O}(1)$. For $t>0$, $q\geq 2$, and $\alpha\asymp t/\widetilde{O}(qR)$, the probability measures $\pi\gamma_{t}$ and $\pi\gamma_{t\,(1+\alpha)}$ satisfy
$\mathcal{R}_{q}(\pi_{t}\mathbin{\|}\pi_{t\,(1+\alpha)})\leq\log 2\,.$
Proof.
We previously used the Rényi version of LSI, but we use (4.6) this time:
$\mathcal{R}_{q}(\pi_{t}\mathbin{\|}\pi_{t\,(1+\alpha)})\leq\log\Bigl(1-\frac{qC_{\mathsf{PI}}(\pi_{t\,(1+\alpha)})}{4}\,\mathsf{RFI}_{q}(\pi_{t}\mathbin{\|}\pi_{t\,(1+\alpha)})\Bigr)^{-1}\,.$
Recall that
$\mathsf{RFI}_{q}(\pi_{t}\mathbin{\|}\pi_{t\,(1+\alpha)})\leq\frac{q\alpha^{2}R^{2}}{t^{2}}\,.$
Taking $\alpha\asymp t\,(qR\,\widetilde{O}(1))^{-1}$, we can ensure the term inside the log is smaller than $1/2$. ∎
Hence, as each phase requires roughly $d^{2}\sigma^{2}$ queries from an $O(1)$-warm start, and each doubling of a given $\sigma^{2}$ only needs $q_{0}R/\sigma^{2}$ (instead of $q_{0}R/\sigma$) when $\sigma^{2}\leq q_{0}R\approx d^{1/2}$, the complexity of each doubling would need $d^{2}R$. Since $R^{2}=d$ due to isotropy, the final complexity is improved to $d^{2.5}$.
We mention that isotropic uniform distributions over hypercube, simplex, and any convex bodies of revolution (e.g., a cone, ball) actually satisfy $O(1)$-quadratic-tilt stability. See §A.
6.2 Counterexample
We show that there exists an isotropic convex body which is not $O(1)$-quadratic-tilt stable. Consider the following convex body in $\mathbb{R}^{d+1}$ from [BIZ26]: for a sufficiently large constant $C_{0}>0$,
$\mathcal{K}=(\sqrt{3}B_{\infty}^{d}\times\mathbb{R})\cap\bigl\{(x,\lambda)\in\mathbb{R}^{d}\times\mathbb{R}:|x|\leq\sqrt{d}+C_{0}\,(1-|\lambda|)\bigr\}\,.$
Due to the symmetry (e.g., the sign flip of $x$ and $\lambda$, and the permutation of $x$-coordinates), the barycenter is at the origin, and the covariance of the uniform distribution is of the form
$\Sigma:=\operatorname{cov}\pi=\mathopen{}\mathclose{{\left[\begin{array}[]{cc}aI_{d}\\ &b\end{array}}}\right]\qquad\text{for some }a,b>0\,.$
Bizeul used this instance to prove the sharpness of an estimate on sub-Gaussianity of quadratic-tilted logconcave probability measures. Bizeul [BIZ26] mentions $\lVert\operatorname{cov}\pi\gamma_{t}\rVert\gtrsim\min(t^{-1},1+t^{2}d)$. Here, we verify that the isotropized version of $\mathcal{K}$ has the peak variance of $d^{1/3}$, thus disproving the quadratic-tilt stability conjecture.
6.2.1 Near-isotropy of $\mathcal{K}$
Below, we use $c,c^{\prime},C^{\prime}$ to denote universal constants, which could be different in each line. For fixed $\lambda$, the slice of $\mathcal{K}$ at $\lambda$ is
$S_{\lambda}:=\{x\in[-\sqrt{3},\sqrt{3}]^{d}:|x|\leq r(\lambda)\}\quad\text{for }r(\lambda):=\sqrt{d}+C_{0}\,(1-|\lambda|)\,$,
so the $\lambda$-marginal of $\pi$ satisfies $\pi^{\Lambda}(\lambda)\propto\operatorname{vol}S_{\lambda}$. For $Z\sim\textnormal{Unif}(\sqrt{3}B_{\infty}^{d})$,
$\frac{\operatorname{vol}S_{\lambda}}{\operatorname{vol}(\sqrt{3}B_{\infty}^{d})}=\mathbb{P}\bigl(|Z|\leq r(\lambda)\bigr)\,.$
Recall Hoeffding’s inequality: for independent random variables $X_{1},\dots,X_{d}$ with $X_{i}\in[a_{i},b_{i}]$ a.s., their sum $S_{d}=\sum_{i\in[d]}X_{i}$ satisfies that
$\mathbb{P}(S_{d}-\mathbb{E}S_{d}\geq s)\leq\exp\bigl(-\frac{2s^{2}}{\sum_{i\in[d]}(b_{i}-a_{i})^{2}}\bigr)\quad\text{for any }s>0\,.$
Using the Hoeffding, for $Z=(Z_{1},\dots,Z_{d})\sim\textnormal{Unif}(\sqrt{3}B_{\infty}^{d})$, due to $\mathbb{E}[Z_{i}^{2}]=1$,
$\mathbb{P}(|Z|^{2}-d\leq-s)\leq\exp\bigl(-\frac{2s^{2}}{9d}\bigr)\,.$
Since $|Z|\leq\sqrt{d}-s$ implies $|Z|^{2}\leq d-2s\sqrt{d}+s^{2}\leq d-s\sqrt{d}$ when $s\leq\sqrt{d}$, it follows that for any $s\in[0,\sqrt{d}]$,
$\mathbb{P}(|Z|\leq\sqrt{d}-s)\leq\exp\bigl(-\frac{2s^{2}}{9}\bigr)\quad\text{(and similarly)}\quad\mathbb{P}(|Z|\geq\sqrt{d}+s)\leq\exp\bigl(-\frac{8s^{2}}{9}\bigr)\,.$ (6.1)
Variance in $\Lambda$-direction.
Let us bound $b:=\operatorname{Var}_{\pi}\Lambda$. Since $b=\mathbb{E}[\lambda^{2}]\geq\mathbb{E}\bigl[\lambda^{2}\,\mathds{1}\{|\lambda|\in[\frac{1}{4},\frac{1}{2}]\}\bigr]\geq\frac{1}{16}\,\mathbb{P}(\frac{1}{4}\leq|\lambda|\leq\frac{1}{2})$, it suffices to show that $\mathbb{P}(\frac{1}{4}\leq|\lambda|\leq\frac{1}{2})=\Omega(1)$ to establish $b=\Omega(1)$. Using $\operatorname{vol}\mathcal{K}=\int\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda=2\int_{\geq 0}\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda$,
$\mathbb{P}\bigl(\frac{1}{4}\leq|\lambda|\leq\frac{1}{2}\bigr)\gtrsim\frac{\int_{1/4}^{1/2}\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda}{\operatorname{vol}\mathcal{K}}=\frac{\int_{1/4}^{1/2}\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda}{2\int_{0}^{1+\sqrt{d}/C_{0}}\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda}$
We show that both denominator and numerator are within constant factors of $\operatorname{vol}(\sqrt{3}B_{\infty}^{d})$. Using the Hoeffding (6.1) and taking $C_{0}$ large enough, for $|\lambda|\leq 1/2$ (which implies $r(\lambda)\geq\sqrt{d}+C_{0}/2$),
$\frac{\operatorname{vol}S_{\lambda}}{\operatorname{vol}(\sqrt{3}B_{\infty}^{d})}=\mathbb{P}\bigl(|Z|\leq r(\lambda)\bigr)\geq\mathbb{P}\bigl(|Z|\leq\sqrt{d}+\frac{C_{0}}{2}\bigr)\geq 1-\exp\bigl(-\Theta(C_{0}^{2})\bigr)\geq 0.99\,.$
When $1\leq|\lambda|\lesssim\sqrt{d}$,
$\frac{\operatorname{vol}S_{\lambda}}{\operatorname{vol}(\sqrt{3}B_{\infty}^{d})}=\mathbb{P}\bigl(|Z|\leq\sqrt{d}-C_{0}\,(|\lambda|-1)\bigr)\lesssim\exp(-c\lambda^{2})\,.$
Hence, $b=\Omega(1)$ follows from
$\int_{0}^{1+\sqrt{d}/C_{0}}\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda=\int_{0}^{1}\cdot+\int_{1}^{1+\sqrt{d}/C_{0}}\cdot\leq\operatorname{vol}(\sqrt{3}B_{\infty}^{d})\,\Bigl(1+\int_{1}^{1+\sqrt{d}/C_{0}}e^{-c\lambda^{2}}\,\mathrm{d}\lambda\Bigr)\lesssim\operatorname{vol}(\sqrt{3}B_{\infty}^{d})\,.$
Also, $b=O(1)$ follows from
$\mathbb{E}[\lambda^{2}]$ $=\mathbb{E}[\lambda^{2}\,\mathds{1}\{|\lambda|\leq 1\}]+\mathbb{E}[\lambda^{2}\,\mathds{1}\{|\lambda|\geq 1\}]\leq 1+\frac{\int_{1}^{1+\sqrt{d}/C_{0}}\lambda^{2}\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda}{\int_{0}^{1+\sqrt{d}/C_{0}}\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda}$
$\leq 1+\frac{\int_{1}^{1+\sqrt{d}/C_{0}}\lambda^{2}\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda}{\int_{0}^{1/2}\operatorname{vol}S_{\lambda}\,\mathrm{d}\lambda}\leq 1+\frac{C^{\prime}\int_{1}^{\infty}\lambda^{2}\,\operatorname{vol}(\sqrt{3}B_{\infty}^{d})\,e^{-c\lambda^{2}}\,\mathrm{d}\lambda}{\operatorname{vol}(\sqrt{3}B_{\infty}^{d})}=O(1)\,.$
Variance in $X$-direction.
Let us bound $a:=\operatorname{Var}_{\pi}X_{i}$ for any $i\in[d]$. Since $|x_{i}|\leq\sqrt{3}$, we have $a=\mathbb{E}[X_{i}^{2}]\leq 3$. For the lower bound, observe that
$da=\mathbb{E}_{\pi}[|X|^{2}]$ $\geq\mathbb{E}_{\pi}[|X|^{2}\,\mathds{1}\{|\lambda|\leq 1/2\}]=\mathbb{E}_{\pi^{\Lambda}}\bigl[\mathds{1}\{|\lambda|\leq 1/2\}\,\mathbb{E}_{\pi^{X|\Lambda=\lambda}}[|X|^{2}]\bigr]\,.$
For $|\lambda|\leq 1/2$ (which is equivalent to $r(\lambda)\geq\sqrt{d}+C_{0}/2$),
$\mathbb{E}_{\pi^{X|\Lambda=\lambda}}[|X|^{2}]$ $=\int_{S_{\lambda}}|x|^{2}\,\frac{\mathrm{d}x}{\operatorname{vol}S_{\lambda}}=\frac{\operatorname{vol}(\sqrt{3}B_{\infty}^{d})}{\operatorname{vol}S_{\lambda}}\int_{\sqrt{3}B_{\infty}^{d}}|x|^{2}\,\frac{\mathds{1}[x\in S_{\lambda}]}{\operatorname{vol}(\sqrt{3}B_{\infty}^{d})}\,\mathrm{d}x\geq\int_{\sqrt{3}B_{\infty}^{d}}|x|^{2}\,\frac{\mathds{1}[x\in S_{\lambda}]}{\operatorname{vol}(\sqrt{3}B_{\infty}^{d})}\,\mathrm{d}x$
$=\mathbb{E}_{Z}[|Z|^{2}\,\mathds{1}\{Z\in S_{\lambda}\}]\geq\mathbb{E}_{Z}[|Z|^{2}\,\mathds{1}\{|Z|\leq\sqrt{d}+C_{0}/2\}]$
$=\mathbb{E}[|Z|^{2}]-\mathbb{E}[|Z|^{2}\,\mathds{1}\{|Z|\geq\sqrt{d}+C_{0}/2\}]$
$\geq d-3d\,\mathbb{P}(|Z|\geq\sqrt{d}+C_{0}/2)\geq d-3d\exp(-c^{\prime}C_{0}^{2})\geq d/2\,$,
where we took $C_{0}$ large enough in the last line. Therefore, $a=\Omega(1)$ follows from
$da=\mathbb{E}_{\pi}[|X|^{2}]\gtrsim d\,\mathbb{P}_{\lambda}(|\lambda|\leq 1/2)\gtrsim d\,.$
6.2.2 Lower bound on the variance in $\lambda$-direction
Let us denote Bizeul’s convex body by $\mathcal{K}_{0}$ in place of $\mathcal{K}$. Define the isotropic scaling
$D=\mathrm{diag}(a^{-1/2}I_{d},b^{-1/2})\,,\qquad\mathcal{K}:=D\mathcal{K}_{0}\,.$
Then $\mu=D_{\#}\pi$ is isotropic, and $\mathcal{K}$ can be written as
$\mathcal{K}=\bigl((L\sqrt{3})B_{\infty}^{d}\times\mathbb{R}\bigr)\cap\{(x,\lambda):|x|\leq R(\lambda)\}\,$,
where $L\asymp 1$, $R(\lambda)=Ld^{1/2}+c\,(1-q^{-1}|\lambda|)$, $q\asymp 1$, and $c\asymp 1$. Note that $R(\lambda)=Ld^{1/2}-c\,(|\lambda|/q-1)$ for $|\lambda|\geq q$.
Quadratic tilting.
For $t>0$, define the tilted measure on $\mathcal{K}$ by $\mu_{t}(\mathrm{d}x,\mathrm{d}\lambda)\propto e^{-t\,(\lVert x\rVert^{2}+\lambda^{2})}\,\mathrm{d}x\mathrm{d}\lambda$. When $(X_{t},\Lambda_{t})\sim\mu_{t}$, it holds that $\mathbb{E}\Lambda_{t}=0$ (by symmetry) and $\mathrm{Var}\Lambda_{t}=\mathbb{E}[\Lambda_{t}^{2}]$.
Proposition 6.3.
$\operatorname{Var}\Lambda_{t}\gtrsim d^{1/3}$ for $t\asymp d^{-1/3}$. Precisely,
$\operatorname{Var}\Lambda_{t}\gtrsim\begin{cases}t^{2}d&\text{if }d^{-1/2}\lesssim t\lesssim d^{-1/3}\,,\\ t^{-1}&\text{if }d^{-1/3}\lesssim t\ll 1\,.\end{cases}$
Let $C=[-L,L]^{d}$, and define the slice $S_{\lambda}=\{x\in C:|x|\leq R(\lambda)\}$ for $\lambda\geq 0$. Define the product probability measure on $C$ as
$Q_{t}(\mathrm{d}x)=\frac{\mathds{1}[x\in C]\,e^{-t\,|x|^{2}}}{\int_{C}e^{-t\,|u|^{2}}\,\mathrm{d}u}\,\mathrm{d}x\propto\prod_{i=1}^{d}(e^{-tx_{i}^{2}}\,\mathds{1}[|x_{i}|\leq L])\,$,
and its $\lambda$-marginal density of $\mu_{t}$ satisfies $p_{t}(\lambda)\propto e^{-t\lambda^{2}}\,Q_{t}(S_{\lambda})$.
Concentration of the base measure $Q_{t}$.
Let $Y=|X|^{2}=\sum_{i=1}^{d}X_{i}^{2}$ (w.r.t. $X\sim Q_{t}$), $m_{1}(t)=\mathbb{E}_{Q_{t}}[X_{1}^{2}]$, and $m(t):=\mathbb{E}_{Q_{t}}Y=dm_{1}(t)$. We also define
$I_{0}(t)=\int_{-L}^{L}\exp(-tx_{1}^{2})\,\mathrm{d}x_{1}\,,\qquad Z_{d}(t)=\int_{C}\exp(-t\,|x|^{2})\,\mathrm{d}x\,.$
Note that
$m_{1}(t)=-\partial_{t}\log I_{0}(t)\,,\qquad m_{1}^{\prime}(t)=-\operatorname{Var}_{Q_{t}}(X_{1}^{2})\,.$
When $t=0$, we have $\mathbb{E}[X_{1}^{2}]=L^{2}$ and $\mathbb{E}[X_{1}^{4}]=9L^{4}/5$ , so $\operatorname{Var}_{Q_{0}}(X_{1}^{2})=4L^{4}/5\asymp 1$. Since the $X_{1}$-marginal of $Q_{t}$ has density proportional to $\exp(-tx_{1}^{2})\,\mathds{1}[|x_{1}|\leq L]$, there exists a small constant $t_{0}$ such that for any $t\leq t_{0}$,
$\operatorname{Var}_{Q_{t}}(X_{1}^{2})=\Theta(1)\,.$
As $m_{1}(t)-L^{2}=-\int_{0}^{t}\operatorname{Var}_{Q_{t}}(X_{1}^{2})\,\mathrm{d}t=-\Theta(t)$, we have
$m(t)=dL^{2}-\Theta(dt)\,.$
Since $X_{i}^{2}\in[0,3L^{2}]$ are i.i.d., Hoeffding yields that for any $s\geq 0$,
$Q_{t}\bigl(Y-m(t)\leq-sd^{1/2}\bigr)\leq\exp\bigl(-\frac{2s^{2}}{9L^{4}}\bigr)\,,\qquad Q_{t}\bigl(Y-m(t)\geq sd^{1/2}\bigr)\leq\exp\bigl(-\frac{2s^{2}}{9L^{4}}\bigr)\,.$
Let $M(t)$ be a median of $Y$. Then, the Hoeffding above implies $|M(t)-m(t)|\lesssim d^{1/2}$. Since $m(t)=dL^{2}-\Theta(dt)$, we have
$M(t)=L^{2}d-\Theta(dt)\pm O(d^{1/2})\,.$ (6.2)
Median cutoff at $\lambda^{*}$.
In the paragraph above, we have dealt with the order of the median of $|X|^{2}$ w.r.t. the base measure $Q_{t}$ (not $\mu_{t}$). We now examine how $Q_{t}$ interacts with each slice $S_{\lambda}$.
Hereafter, we assume that $t\gtrsim d^{-1/2}$, for which $M(t)\leq L^{2}d=R(q)^{2}$. By the definition of $R(\lambda)$ and $S_{\lambda}$, as we increase $|\lambda|(\geq q)$, the slice $S_{\lambda}$ is truncated by a ball of smaller radius $R(\lambda)$. Hence, there exists a unique $\lambda^{*}$ such that
$M(t)=R(\lambda^{*})^{2}\,.$
That is, we are attempting to find the point $\lambda^{*}$, where the median of $Q_{t}$ is cut off.
Letting $x=c\,(|\lambda|/q-1)$, we should solve
$x$ $=\sqrt{L^{2}d}-\sqrt{L^{2}d-\Theta(dt)\pm O(d^{1/2})}=\frac{\Theta(dt)\pm O(d^{1/2})}{\sqrt{L^{2}d}+\sqrt{L^{2}d-\Theta(dt)+O(d^{1/2})}}$
$=\frac{\Theta(dt)}{\sqrt{L^{2}d}+\sqrt{L^{2}d-\Theta(dt)+O(d^{1/2})}}\pm O(1)\,.$
When $t\ll 1$ (so that $\Theta(dt)\ll L^{2}d$), we have $x\pm O(1)\asymp\frac{\Theta(dt)}{2\sqrt{L^{2}d}}\asymp t\sqrt{d}$. Therefore, when $t\ll 1$,
$\lambda^{*}\asymp 1+t\sqrt{d}\,.$
Gaussian tail of $p_{t}(\lambda)$ after $\lambda\geq\lambda^{*}$.
The definition of $\lambda^{*}$ implies that for $|\lambda|\leq\lambda^{*}$,
$Q_{t}\bigl(Y\leq R(\lambda)^{2}\bigr)\geq Q_{t}\bigl(Y\leq M(t)\bigr)=\frac{1}{2}\,.$
When $\lambda=\lambda^{*}+s$ for $s>0$, the endpoint $\lambda_{\max}$ (i.e., $R(\lambda_{\max})=0$) satisfies that $\lambda_{\max}=q+qc^{-1}Ld^{1/2}$. Thus,
$0\leq s\leq s_{\max}:=\lambda_{\max}-\lambda^{*}\leq\lambda_{\max}-q\leq qc^{-1}Ld^{1/2}\,.$
Using $s\leq s_{\max}\leq qc^{-1}Ld^{1/2}$ and $R(\lambda^{*}+s)-R(\lambda^{*})=-q^{-1}cs$,
$R(\lambda^{*})^{2}-R(\lambda^{*}+s)^{2}=q^{-1}cs\,\bigl(2R(\lambda^{*})-q^{-1}cs\bigr)\geq q^{-1}cs\,\bigl(2R(\lambda^{*})-Ld^{1/2}\bigr)\,.$ (6.3)
Taking $t\lesssim L^{2}$, we have
[…]
Putting this back to (6.3),
$R(\lambda^{*})^{2}-R(\lambda^{*}+s)^{2}\geq\frac{csLd^{1/2}}{2q}\,.$
Using $|m(t)-M(t)|\lesssim\sqrt{d}$ with $R(\lambda^{*})^{2}=M(t)$,
$R(\lambda^{*}+s)^{2}\leq R(\lambda^{*})^{2}-\frac{csLd^{1/2}s}{2q}=m(t)\pm O(d^{1/2})-\frac{csLd^{1/2}}{2q}\,.$
Hence, for $1\lesssim s\leq s_{\max}$ and some constant $c^{\prime}>0$,
$Q_{t}\bigl(Y\leq R(\lambda^{*}+s)^{2}\bigr)\leq Q_{t}\bigl(Y\leq m(t)-sd^{1/2}\bigr)\leq\exp(-c^{\prime}s^{2})\,.$
First regime: $d^{-1/2}\lesssim t\lesssim d^{-1/3}$.
Let $\bar{p}_{t}(\lambda)=e^{-t\lambda^{2}}\,Q_{t}(S_{\lambda})$ (i.e., an unnormalized density of $p_{t}(\lambda)$). The result above implies that for $\lambda=\lambda^{*}+s$ with $s\gtrsim 1$,
$\bar{p}_{t}(\lambda^{*}+s)\leq e^{-t\,(\lambda^{*}+s)^{2}}e^{-c^{\prime}s^{2}}\leq e^{-c^{\prime}s^{2}}\,.$
Observe that
$\operatorname{Var}\Lambda_{t}=\mathbb{E}[\Lambda_{t}^{2}]=\frac{\int_{\geq 0}\lambda^{2}\bar{p}_{t}(\lambda)\,\mathrm{d}\lambda}{\int_{\geq 0}\bar{p}_{t}(\lambda)\,\mathrm{d}\lambda}\,.$
As for the denominator,
$\int_{\geq 0}\bar{p}_{t}(\lambda)\,\mathrm{d}\lambda=\int_{0}^{\lambda^{*}+\Theta(1)}+\int_{\lambda^{*}+\Theta(1)}^{\infty}\leq\lambda^{*}+\Theta(1)\,.$
As for the numerator, since $t\,(\lambda^{*})^{2}\asymp t^{3}d\lesssim 1$, we have $\exp(-t\lambda^{2})\gtrsim 1$ and $Q_{t}(Y\leq R(\lambda)^{2})\geq 1/2$ on $\lambda\in[0,\lambda^{*}]$. Hence,
$\int_{\geq 0}\lambda^{2}\,\bar{p}_{t}(\lambda)\,\mathrm{d}\lambda\geq\int_{0}^{\lambda^{*}}\lambda^{2}\,\bar{p}_{t}(\lambda)\,\mathrm{d}\lambda\gtrsim\int_{0}^{\lambda^{*}}\lambda^{2}\,\mathrm{d}\lambda\asymp(\lambda^{*})^{3}\,.$
Therefore, in the regime of $t\lesssim d^{-1/3}$,
$\operatorname{Var}\Lambda_{t}\gtrsim\frac{(\lambda^{*})^{3}}{1+\lambda^{*}}\,.$
Since $\lambda^{*}\asymp 1+td^{1/2}\gtrsim 1$ (due to $t\gtrsim d^{-1/2}$), the variance is bounded lower by $(\lambda^{*})^{2}\asymp t^{2}d$.
Second regime: $d^{-1/3}\lesssim t\ll 1$.
We first show that in this regime, $R(\lambda)^{2}\geq M(t)$ when $|\lambda|\lesssim t^{-1/2}$. For $r:=c\,(|\lambda|/q-1)$,
$M(t)$ $=m(t)\pm O(d^{1/2})=L^{2}d-\Theta(dt)\pm O(d^{1/2})\,$,
$R(\lambda)^{2}$ $=L^{2}d-2Lrd^{1/2}+r^{2}\geq L^{2}d-2rLd^{1/2}\,.$
Hence, $R(\lambda)^{2}\geq M(t)$ holds if $\Theta(dt)-O(d^{1/2})\geq 2Lrd^{1/2}$ (equivalently, $r\leq\Theta(d^{1/2}t)-O(1)$ and $|\lambda|\lesssim 1+d^{1/2}t$). When $d^{-1/3}\lesssim t$, we have $t^{-1/2}\lesssim d^{1/2}t$. Then, the claim follows from
$|\lambda|\lesssim\frac{1}{t^{1/2}}\lesssim d^{1/2}t\leq 1+d^{1/2}t\,.$
Let us lower-bound the variance. As for the denominator, due to $Q_{t}(\cdot)\leq 1$,
$\int_{\geq 0}\bar{p}_{t}(\lambda)\,\mathrm{d}\lambda\leq\int_{\geq 0}e^{-t\lambda^{2}}\,\mathrm{d}\lambda\asymp\frac{1}{t^{1/2}}\,.$
As for the numerator, $Q_{t}(Y\leq R(\lambda)^{2})\geq 1/2$ on $|\lambda|\lesssim t^{-1/2}$. Hence, for some constant $c^{\prime}>0$,
$\int\lambda^{2}\,\bar{p}_{t}(\lambda)\,\mathrm{d}\lambda\gtrsim\int_{-c^{\prime}/t^{1/2}}^{c^{\prime}/t^{1/2}}\lambda^{2}\,e^{-t\lambda^{2}}\,\mathrm{d}\lambda\asymp\frac{1}{t^{3/2}}\,.$
Therefore,
$\operatorname{Var}\Lambda_{t}=\frac{\int\lambda^{2}\bar{p}_{t}(\lambda)\,\mathrm{d}\lambda}{2\int_{\geq 0}\bar{p}_{t}(\lambda)\,\mathrm{d}\lambda}\gtrsim\frac{1}{t}\,.$
Acknowledgement.
Part of this work was done while YK was a student researcher at Google Research. The authors thank Sinho Chewi and Andre Wibisono for helpful comments. This work was supported in part by NSF Award CCF-2106444 and a Simons Investigator award.
Appendix A Quadratic-tilt stability of structured distributions
A.1 Hypercube
Consider the isotropic hypercube, $\mathcal{K}=\otimes_{i=1}^{d}[-\sqrt{3},\sqrt{3}]$. Then, any truncated Gaussian $\pi\gamma_{t}$ (here $\pi\propto\mathds{1}_{\mathcal{K}}$) does not violate the conjecture. Since $\pi\gamma_{t}$ can be written as $d$-tensor product of one-dimensional probability measures, $\operatorname{cov}\pi\gamma_{t}$ is a diagonal matrix. Clearly, each diagonal (i.e., variance along any coordinate) is $O(1)$ since the support has diameter of $O(1)$.
A.2 Simplex
Consider the regular $d$-simplex in isotropic position. Let us embed the simplex in the ambient $(d+1)$-dimensional space:
$\Delta_{d}=\Bigl\{x\in\mathbb{R}^{d+1}:x_{i}\geq 0,\sum_{i=1}^{d+1}x_{i}=1\Bigr\}\,$,
centered at $\bar{x}=\frac{1}{d+1}\,(1,\dots,1)$, so that $y:=x-\bar{x}$ lies in the hyperplane $H:=\{y\in\mathbb{R}^{d+1}:\sum_{i}y_{i}=0\}$.
Now, consider the quadratically-tilted measure $\pi\gamma_{t}$ over the isotropized simplex. Using the coordinate symmetry (i.e., permuting coordinates), we can write $C:=\operatorname{cov}\pi\gamma_{t}=\alpha I_{d+1}+\beta\,(J_{d+1}-I_{d+1})$ for some $\alpha,\beta\in\mathbb{R}$, where $J_{d+1}$ is the all-ones matrix. Since $y\in H$ satisfies $y\perp\mathbf{1}$, we have $J_{d+1}y=0$. Thus, the restriction of $C$ to $H$ satisfies
$Cy=(\alpha-\beta)\,y\quad\text{for all }y\in H\,$,
so $\operatorname{cov}(\pi\gamma_{t})|_{H}=(\alpha-\beta)\,I_{d}=:cI_{d}$. Using the similar argument in (4.5),
$\mathbb{E}_{\pi\gamma_{t}}[|Y|^{2}]=\frac{\mathbb{E}_{\pi}[|Y|^{2}\,e^{-|Y|^{2}/(2t)}]}{\mathbb{E}_{\pi}[e^{-|Y|^{2}/(2t)}]}\leq\mathbb{E}_{\pi}[|Y|^{2}]=d\,$,
where the last equality holds since $\pi$ is isotropic. Since $\mathbb{E}_{\pi\gamma_{t}}Y=0$ by symmetry, we have $\operatorname{tr}(\operatorname{cov}\pi\gamma_{t})=\mathbb{E}_{\pi\gamma_{t}}[|Y|^{2}]\leq d$. As $\operatorname{tr}(cI_{d})=cd$, it follows that $c\leq 1$. Hence, $\lVert\operatorname{cov}\pi\gamma_{t}\rVert=c=O(1)$.
A.3 Isotropic convex body of revolution
For an interval $I\subset\mathbb{R}$ and a concave function $r:\mathbb{R}\to\mathbb{R}_{\geq 0}$, let
$\mathcal{K}=\{(s,y)\in\mathbb{R}\times\mathbb{R}^{d-1}:s\in I,\,|y|\leq r(s)\}\,$,
where the uniform distribution over $\mathcal{K}$ is isotropic. Consider a quadratically-tilted measure over $\mathcal{K}$:
$\pi_{t}(\mathrm{d}x)\propto\exp\bigl(-\frac{|x|^{2}}{2t}\bigr)\,\mathds{1}_{\mathcal{K}}(x)\,\mathrm{d}x\qquad\text{for }x=(s,y)\,.$
A.3.1 $y$-direction
Let $\sigma_{\perp}^{2}$ be the variance of $\pi_{t}$ in the $y$-direction. Using the rotational symmetry and the similar argument in (4.5),
$\sigma_{\perp}^{2}\leq\frac{1}{d-1}\,\mathbb{E}_{\pi_{t}}[|Y|^{2}]\leq\frac{1}{d-1}\,\mathbb{E}_{\pi}[|Y|^{2}]\leq\frac{d}{d-1}=O(1)\,.$
A.3.2 $s$-direction
Conditionally on $S=s$, the radial coordinate $U=|Y|$ has density, for $R=r(s)$
$\nu_{t,R}(\mathrm{d}u)\propto u^{d-2}\exp\bigl(-\frac{u^{2}}{2t}\bigr)\,\mathds{1}_{[0,R]}(u)\,\mathrm{d}u\,.$
Hence, integrating $y\in\mathbb{R}^{d-1}$ gives the following $s$-marginal:
$p_{t}(s)\propto\exp\bigl(-\frac{s^{2}}{2t}\bigr)\,H_{t}\bigl(r(s)\bigr)\qquad\text{for }H_{t}(R):=\int_{0}^{R}u^{d-2}e^{-\frac{u^{2}}{2t}}\,\mathrm{d}u\,.$
Lemma A.1.
Let $\eta_{R}$ be the probability measure on $[0,R]$ with density proportional to $u^{d-2}$. If $f$ is increasing and $g$ is decreasing on $[0,R]$, then $\mathbb{E}_{\eta_{R}}[fg]\leq\mathbb{E}_{\eta_{R}}f\,\mathbb{E}_{\eta_{R}}g$. In particular, for $f(u)=u^{2}$ and $g(u)=e^{-u^{2}/(2t)}$,
$\mathbb{E}_{\nu_{t,R}}[U^{2}]\leq\mathbb{E}_{\eta_{R}}[U^{2}]=\frac{d-1}{d+1}\,R^{2}\,.$
Proof.
The first claim can be deduced from the definition of $\operatorname{cov}_{\eta_{R}}(f,g)$ (see (4.5)). For $f(u)=u^{2}$ and $g(u)=e^{-u^{2}/(2t)}$, we have
$\mathbb{E}_{\eta_{R}}[u^{2}e^{-u^{2}/(2t)}]\leq\mathbb{E}_{\eta_{R}}[u^{2}]\,\mathbb{E}_{\eta_{R}}e^{-u^{2}/(2t)}\,.$
Equivalently,
$\int_{0}^{R}u^{d-2}\,\mathrm{d}u\int_{0}^{R}u^{2}u^{d-2}e^{-u^{2}/(2t)}\,\mathrm{d}u\leq\int_{0}^{R}u^{2}u^{d-2}\,\mathrm{d}u\int_{0}^{R}u^{d-2}e^{-u^{2}/(2t)}\,\mathrm{d}u\,$,
which implies that
$\mathbb{E}_{\nu_{t,R}}[U^{2}]\leq\mathbb{E}_{\eta_{R}}[U^{2}]=\frac{\int_{0}^{R}u^{d}\,\mathrm{d}u}{\int_{0}^{R}u^{d-2}\,\mathrm{d}u}=\frac{R^{d+1}/(d+1)}{R^{d-1}/(d-1)}=\frac{d-1}{d+1}\,R^{2}\,.$
This completes the proof. ∎
Maximum of $r(s)$.
From above, $\mathbb{E}[|Y|^{2}\mid S=s]=\frac{d-1}{d+1}\,r(s)^{2}$. Define the axial density $p_{0}(s)\propto r(s)^{d-1}\,\mathds{1}_{I}(s)$ and recall that $p_{0}$ is isotropic logconcave on $\mathbb{R}$. Taking expectation with respect to $s\sim p_{0}$,
$d-1=\mathbb{E}_{\pi}[|Y|^{2}]=\frac{d-1}{d+1}\,\mathbb{E}_{p_{0}}[r(s)^{2}]\,$,
so
$\frac{\int_{I}r(s)^{d+1}\,\mathrm{d}s}{\int_{I}r(s)^{d-1}\,\mathrm{d}s}=\mathbb{E}_{p_{0}}[r(s)^{2}]=d+1\,.$
Using this and concavity of $r(s)$, we show the following:
Lemma A.2.
In the setting above,
$r_{\max}^{2}=\max_{s\in I}r(s)^{2}\leq 16d\,.$
Proof.
Let $R=\max_{I}r(s)$, achieved at $s^{*}$. Then, it follows from the concavity of $r(s)$ that the set $J:=\{s:r(s)\geq R/2\}$ has length at least $|I|/2$. To see this, consider the chord between the left endpoint $(a,r(a))$ and $(s^{*},R)$. Concavity of $r(s)$ implies that $r$ is above $R/2$ (as $r(a)\geq 0$). Repeating the same argument for the right endpoint, the claim follows.
Let us show that
$\frac{\int_{I}r(s)^{d+1}\,\mathrm{d}s}{\int_{I}r(s)^{d-1}\,\mathrm{d}s}\gtrsim R^{2}\,.$
To this end, using $|J|\geq|J^{c}|$,
$\int_{I}r(s)^{d-1}\,\mathrm{d}s=\int_{J}+\int_{J^{c}}\leq\int_{J}+\int_{J}=2\int_{J}r(s)^{d-1}\,\mathrm{d}s\,.$
Also,
$\int_{I}r(s)^{d+1}\,\mathrm{d}s\geq\int_{J}r(s)^{d+1}\,\mathrm{d}s\geq\frac{R^{2}}{4}\int_{J}r(s)^{d-1}\,\mathrm{d}s\,.$
Hence,
$(d+1=\mathbb{E}_{p_{0}}[r(s)^{2}]=)\frac{\int_{I}r(s)^{d+1}\,\mathrm{d}s}{\int_{I}r(s)^{d-1}\,\mathrm{d}s}\geq\frac{R^{2}}{8}\,$,
which completes the proof. ∎
Main proof.
We need some identity:
Lemma A.3.
For $H_{t}(R)=\int_{0}^{R}u^{d-2}e^{-u^{2}/(2t)}\,\mathrm{d}u$, it holds that
$\frac{d-1}{R}\,\bigl(1-\frac{R^{2}}{t\,(d+1)}\bigr)\leq\frac{H_{t}^{\prime}(R)}{H_{t}(R)}=\frac{d-1}{R}-\frac{1}{tR}\,\mathbb{E}_{\nu_{t,R}}[U^{2}]\leq\frac{d-1}{R}\,.$
In particular, if $R^{2}\leq t\,(d+1)/2$, then $\frac{H_{t}^{\prime}(R)}{H_{t}(R)}\geq\frac{d-1}{2R}$.
Proof.
Using IBP and $H_{t}^{\prime}(R)=R^{d-2}e^{-R^{2}/2t}$,
$H_{t}(R)=e^{-\frac{R^{2}}{2t}}\,\frac{R^{d-1}}{d-1}+\frac{1}{d-1}\int_{0}^{R}\frac{u^{d}}{t}\,e^{-\frac{u^{2}}{2t}}\,\mathrm{d}u=\frac{R}{d-1}\,H_{t}^{\prime}(R)+\frac{1}{d-1}\int_{0}^{R}\frac{u^{d}}{t}\,e^{-\frac{u^{2}}{2t}}\,\mathrm{d}u\,.$
Dividing both sides by $RH_{t}(R)/(d-1)$,
$\frac{H_{t}^{\prime}(R)}{H_{t}(R)}=\frac{d-1}{R}-\frac{1}{tR}\frac{\int_{0}^{R}u^{d}e^{-\frac{u^{2}}{2t}}\,\mathrm{d}u}{\int_{0}^{R}u^{d-2}e^{-\frac{u^{2}}{2t}}\,\mathrm{d}u}=\frac{d-1}{R}-\frac{1}{tR}\,\mathbb{E}_{\nu_{t,R}}[U^{2}]\,.$
The claim follows from Lemma A.1. ∎
Since $p_{0}$ is isotropic, we can take a mode within [10,10][-10,10] (see Claim A.4). By Claim A.5, there exist universal $A,c>0$ such that $(\log p_{0})^{\prime}(s)\leq-c$ for all $s\geq A$. Differentiating $\log p_{t}(s)=-\frac{s^{2}}{2t}+\log H_{t}(r(s))+\mathrm{const}$ yields
$(\log p_{t})^{\prime}(s)=-\frac{s}{t}+r^{\prime}(s)\,\frac{H_{t}^{\prime}\bigl(r(s)\bigr)}{H_{t}\bigl(r(s)\bigr)}\,.$
Using Lemma A.2, we take $t\geq 32$ to satisfy $\frac{H_{t}^{\prime}(r(s))}{H_{t}(r(s))}\geq\frac{d-1}{2r(s)}$. On the right tail (where $r^{\prime}(s)\leq 0$ by concavity), we have
$(\log p_{t})^{\prime}(s)\leq-\frac{s}{t}+\frac{d-1}{2}\,\frac{r^{\prime}(s)}{r(s)}=-\frac{s}{t}+\frac{1}{2}\,(\log p_{0})^{\prime}(s)\leq-\frac{c}{2}\qquad\text{for }s\geq A\,.$
Thus, $p_{t}$ has a universal exponential right tail, which implies that second moment contributed by the right tail (after $A=\Omega(1)$) is $O(1)$. We can handle the left tail in a similar manner. Also, the second moment coming from [10,10][-10,10] is $O(1)$. Hence, $\mathbb{E}_{p_{t}}[S^{2}]=O(1)$ and $\mathrm{Var}_{p_{t}}S=O(1)$ for $t\in[32,d^{1/2}]$. We can handle the case of $t\in[1,32]$ by the Brascamp–Lieb inequality inequality.
Missing details.
Claim A.4.
The mode of an isotropic logconcave distribution in $\mathbb{R}$ is within [10,10][-10,10].
Proof.
Let $g$ be the pdf of the isotropic logconcave distribution, and $m$ a mode of $g$. If $m>10$, then log-concavity implies that $g$ is non-decreasing on $(-\infty,m]$. Thus, on $[0,10]$, we have $g(x)\geq g(0)\geq 1/8$ from [LV07],
$1=\int g(x)\,\mathrm{d}x\geq\int_{0}^{10}g(x)\,\mathrm{d}x\geq\frac{10}{8}\,$,
which is contradiction. Repeating the same argument for m<10m<-10, we conclude that $|m|\leq 10$. ∎
Let $A\geq 11$. Then, $g$ is non-increasing on $A-1,\infty)$ since $m\leq 10$. Hence,
$g(A)\leq\int_{A-1}^{\infty}g(u)\,\mathrm{d}u=\mathbb{P}(X\geq A-1)\leq\mathbb{P}(|X|\geq A-1)\leq e^{-(A-1)+1}=e^{-A+2}\,.$
Hence,
$\frac{g(A)}{g(0)}\leq 8e^{-A+2}\,.$
Claim A.5.
There exist universal $A,c>0$ such that $(\log p_{0})^{\prime}(s)\leq-c$ for all $s\geq A$.
Proof.
Take the concave function $\phi(s)=\log g(s)$, where $g(s)$ is the pdf of the isotropic logconcave $p_{0}$. Then, for $A\geq 11$,
$\phi^{\prime}(A)\leq\frac{\phi(A)-\phi(0)}{A}=\frac{1}{A}\,\log\frac{g(A)}{g(0)}\leq\frac{1}{A}\,(\log 8-A+2)\,.$
For $A=\Omega(1)$, $\phi^{\prime}(A)\leq-1/2$. Since $\phi^{\prime}$ is non-increasing due to concavity, for all $x\geq A$,
$\phi^{\prime}(x)=(\log g)^{\prime}(x)\leq\phi^{\prime}(A)\leq-1/2\,$,
which completes the proof. ∎
Appendix B Related work
We survey the literature on applications of zeroth-order logconcave sampling.
B.1 More on uniform sampling
In early work, researchers used many pioneering ideas to obtain polynomial-time algorithms. In [DFK91], the input body $\mathcal{K}$ is first mollified and then sampled using the $\mathsf{Grid}\text{ }\mathsf{walk}$, a random walk on a suitably fine lattice over the body. This algorithm is then analyzed through the conductance of the resulting Markov chain, the well-known framework developed by Jerrum and Sinclair [JS88]. To use this, they applied an isoperimetric inequality of Milman and Schechtman [MS86], showing that the $\mathsf{Grid}\text{ }\mathsf{walk}:\mathcal{R}_{\infty}\to\mathsf{TV}$ requires $\operatorname{poly}(d,D,\nicefrac{{M_{\infty}}}{{\varepsilon}})$ many queries to achieve $\varepsilon$-$\mathsf{TV}$ distance.
Lovász and Simonovits [LS90] generalized the conductance framework to continuous domains, introducing $s$-conductance that lower bounds the conductance of the chain for subsets of measure at least $s>0$. To prove lower bounds on $s$-conductance, they developed the well-known localization lemma, proving a Cheeger isoperimetric inequality for $\pi_{\mathcal{K}}$, with $C_{\mathsf{Ch}}(\pi_{\mathcal{K}})\gtrsim D^{-1}$ (which was originally established by Payne and Weinberger [PW60]). These new tools also improved the complexity of the $\mathsf{Grid}\text{ }\mathsf{walk}$, although it still has dependence of $\operatorname{poly}(d,D,\nicefrac{{M_{\infty}}}{{\varepsilon}})$.
As mentioned in the main body of the paper, Lovász and Simonovits [LS93] used the $\mathsf{Ball\ walk}$ to develop an indirect approach with improved complexity $\operatorname{poly}(d,D,\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$. It was achieved by refining their earlier conductance framework and localization machinery; they proved that any logconcave distribution $\pi$ supported on a convex body of diameter $D$ also satisfies $C_{\mathsf{Ch}}(\pi)\gtrsim D^{-1}$. [KLS97] provided a refined version of these results with improved isoperimetry, an efficient rounding algorithm, the speedy walk and other extensions. The study of logconcave sampling was initiated by Applegate and Kannan [AK91]. Frieze and Kannan [FK99] used a log-Sobolev inequality together with a grid walk to give an improved complexity for logconcave sampling.
Recently, $\mathsf{Coordinate\ Hit\text{-}and\text{-}Run}$ ($\mathsf{CHR}$), a special variant of $\mathsf{Hit\text{-}and\text{-}Run}$, has been shown to mix in polynomial time with respect to the dimension and ($\ell_{2}$ or $\ell_{\infty}$) diameter [NS22, LV23, FER23], and with a logarithmic dependence on the initial warmness [NRS25]. This thus results in $\mathsf{CHR}:\mathcal{R}_{2}\to\mathcal{R}_{2}$, although the mixing time is a larger polynomial than that of hit-and-run.
B.2 KLS conjecture
The Kannan-Lovász-Simonovits (KLS) conjecture [KLS95] posits that the Cheeger constant of any logconcave probability measure is bounded below by $\Omega(\lVert\operatorname{cov}\pi\rVert^{-1/2})$. Starting from $C_{\mathsf{Ch}}^{2}\gtrsim\operatorname{tr}^{-1}(\operatorname{cov})$, subsequent work [ELD13, LV24, CHE21, KL22, KLA23]—notably through Eldan’s stochastic localization—has improved it to $C_{\mathsf{Ch}}^{2}\gtrsim(\lVert\operatorname{cov}\pi\rVert\log d)^{-1}$ [KLA23].
B.3 Rounding
A notable application is the rounding of convex bodies, or more generally, logconcave functions. It aims to make the geometry of the input object more regular—typically by applying an affine transformation that brings it into a nearly-isotropic position. This step is often a necessary preprocessing phase for achieving a fully polynomial-time algorithm, since worst-case shapes (e.g., highly skewed bodies) can cause sampling algorithms to suffer from poor mixing.
Earlier approaches to rounding relied on deterministic methods based on the ellipsoid algorithm [DFK91, LS90], which guarantees $B_{1}(0)\subset\mathcal{K}\subset B_{d^{3/2}}(0)$ after a suitable affine transformation. To improve the sandwiching ratio, randomized rounding methods were pioneered in [LS93] and further developed in [KLS97, LV06c, JLL+21, JLL+26]. These methods use sampling to estimate the covariance matrix and then apply affine transformations to round the body. The best known complexity for isotropic rounding of a convex body is $\widetilde{O}(d^{3.5}\operatorname{polylog}R)$ [JLL+26], with the analysis streamlined in [KZ24]. These rounding techniques have been extended to logconcave distributions [LV06a, KV25b], while preserving the query complexity.
B.4 Volume computation and integration
One of the most well-known applications of uniform sampling is volume computation, a problem with a long and rich history in theoretical computer science and convex geometry; given $\varepsilon>0$, the goal is to estimate the volume $V$ of a convex body within a relative $\varepsilon$-error, $(1-\varepsilon)\operatorname{vol}\mathcal{K}\leq V\leq(1+\varepsilon)\operatorname{vol}\mathcal{K}$.
Computing the volume of a high-dimensional convex body is known to be intractable via deterministic algorithms [ELE86, BF87]. The breakthrough result by Dyer, Frieze, and Kannan [DFK91] showed that this task admits a randomized polynomial-time algorithm using random walks. The standard technique introduced therein, referred to as the multi-phase Monte Carlo method, constructs a sequence of convex bodies such that each successive pair is close in some sense. Rewriting the volume of $\mathcal{K}$ as a telescoping product of volume ratios between successive bodies, one can efficiently estimate each ratio using samples obtained via uniform sampling over convex bodies.
This framework initiated a long line of research focused on improving the mixing time, oracle complexity, and practical implementability of such algorithms [LOV90, AK91, LS90, LS93, KLS95, KLS97, LV06b, CV15, CV18, JLL+21]. The current fastest algorithm for well-rounded convex bodies (a weaker condition than near-isotropic) is due to [CV18] with complexity $\widetilde{O}(d^{3}/\varepsilon^{2})$. A natural generalization of volume computation is the integration of a logconcave function. This general problem was studied in [LV06a, KV25b], establishing the same complexity $\widetilde{O}(d^{3}/\varepsilon^{2})$ when the corresponding distribution is well-rounded; this can be applied after rounding for an arbitrary logconcave function. See [KV25b, §5.2] for a streamlined analysis.